### v3

# Aspects of the commuting graph [^1] To the memory of Richard Parker

Abstract
The commuting graph of a group $G$ is the graph whose vertices are the elements of $G$, two distinct vertices joined if they commute.
Our purpose in this paper is twofold.
  1. (a)
    We discuss the computational problem of deciding whether a given graph is the commuting graph of a finite group; we give a quasipolynomial algorithm, and a polynomial algorithm for the case when the group is an extraspecial $p$-group for $p$ an odd prime.
  2. (b)
    We give new results on the question of whether the commuting graph of a given group is a cograph or a chordal graph, two classes of graphs defined by forbidden subgraphs.
The problems are not unrelated, since there are a number of cases where hard computational problems on graphs are easier when restricted to special classes of graphs; we conjecture that the recognition problem is polynomial for cographs and chordal graphs.
keywords:
finite group , commuting graph , forbidden subgraph , cograph , chordal graph , graph isomorphism , quasipolynomial algorithm , extraspecial group
MSC:
[2020] 05C25 , 05C85 , 20D60
1]arvind@imsc.res.in 2]pjc20@st-andrews.ac.uk 3]xuanlma@xsyu.edu.cn 4]butterson@mail.ru
1 Introduction
Graphs associated with some algebraic structures have been actively investigated, such as, the power graph of a group (cf. [1, 20, 22, 24]), and of course the famous Cayley graphs which have a long history. Moreover, graphs from algebraic structures have valuable applications (cf. [45]): for example, Cayley graphs as classifiers for data mining [46].
Let $G$ be a group and $S$ be a subset of $G$. The commuting graph of $G$ on $S$ is a simple graph with vertex set $S$, and two distinct elements are adjacent in this graph if they commute in $G$. In particular, if $S=G$, then we denote this commuting graph on $G$ by $\Gamma(G)$. In 1955, Brauer and Fowler [12] first introduced the commuting graph of a group and showed that there are only finitely many simple groups with even order such that they have a prescribed centralizer. In fact, in the seminal paper [12], they considered the group commuting graph on the set of all non-trivial elements. Since then, the commuting graphs have received considerable attention and were well studied in the literature. For example, Segev et al. [64, 65, 66] established some longstanding conjectures on division algebras by using some combinatorial parameters of commuting graphs. Britnell and Gill [15] classified all finite quasi-simple groups whose commuting graph is perfect, where the vertex set of the commuting graph is the set consisting of all non-central elements. Giudici and Parker [38] showed that for every positive integer $d$, there is a finite $2$-group $G$ such that the commuting graph of $G$ on the set of all non-central elements has diameter greater than $d$; further results are found in [30]. Morgan and Parker [58] proved that, for a finite group $G$ with trivial center, every connected component of the commuting graph of $G$ on the set of all non-trivial elements has diameter at most $10$. Recently, Lewis et al. [9] extended some results by Morgan and Parker [58] about commuting graphs. Moreover, some further applications of commuting graphs in group theory can be found in [21].
The first part of this paper is devoted to the algorithmic question: Given a graph $\Gamma$, is it isomorphic to the commuting graph of a group? We give a quasi-polynomial algorithm in general, and a polynomial algorithm in the case of extraspecial $p$-groups with $p$ odd.
A number of important graph classes, including perfect graphs, cographs, chordal graphs, split graphs, and threshold graphs, can be defined either structurally or in terms of forbidden induced subgraphs. Forbidden subgraphs of power graphs were studied in [54]. In [21], Cameron described various graphs whose vertex set is the universe of a group and whose edges reflect group structure in some way, and proposed the following question.
Question 1
([21, Question 14]) For which finite groups is the commuting graph a perfect graph, or a cograph, or a chordal graph, or a split graph, or a threshold graph?
Britnell and Gill [15] studied the finite groups whose commuting graph is perfect. Ma and Cameron [51] completely classified the finite groups whose commuting graph is a split graph or a threshold graph.
In the second part of the paper, we address the question of when the commuting graph is a cograph or a chordal graph. We prove that in some sense, the class of such groups generalizes the class of finite solvable groups (see Theorem 11.7 in Subsection 11.3). Also we determine all finite simple groups whose commuting graph is a cograph (see Problem 1 in Subsection 11.3 and Theorem 12.15 at the end of Section 12). The results partly answer Question 1.
These issues are relevant to the recognition question. There are extra hypotheses which could make the problem of recognition of commuting graphs easier. One is to restrict the groups to a known class, such as nilpotent groups. Another is to restrict the graphs to a class in which computation is easier. For example, the result of the second and third authors describes precisely the split or threshold graphs which are commuting graphs; these are the complete graphs, together with the graphs consisting of an $n$-vertex complete graph (where $n$ is odd), with $n$ pendant edges added at one of its vertices. Clearly these are easily recognised. We have not pursued this issue further.
We note that some works, especially on connectedness and diameter, define the commuting graph to have vertex set $G\setminus Z(G)$ (that is, omitting dominating vertices). This has little effect on our results. For the recognition problem, if we are given the commuting graph on $G\setminus Z(G)$, then the commuting graph on $G$ is found by adding $|Z(G)|$ dominating vertices to the graph, and this number is a divisor of $|G\setminus Z(G)|$ so there are not many possibilities. For the second part of the paper, see Proposition 10.2.
We conclude this section with a suggestion for further research, which would link the two parts of the paper together more closely.
Question 2
Suppose that $\Gamma$ is a perfect graph, a cograph or a chordal graph. Is there a polynomial algorithm to decide whether $\Gamma$ is the commuting graph of a group?
2 Preliminaries
In this section, we will introduce some notation and recall some definitions on groups and graphs.
Every group considered is finite. Throughout, $G$ always denotes a finite group and $e_{G}$ or $1_{G}$ denotes the identity element of $G$, and we denote $e_{G}$ simply by $e$ or $1$ if the considered situation is unambiguous. Let $g\in G$. The order $o(g)$ of $g$ in $G$ is the size of the cyclic subgroup $\langle g\rangle$ generated by $g$. The element $g$ is called an involution if $o(g)=2$. The center of $G$, denoted by $Z(G)$, is defined as
$\{x\in G:gx=xg\hbox{ for each $g\in G$}\}.$
Also, the centralizer $C_{G}(g)$ of $g$ in $G$ is the set of all elements that commute with $g$, that is,
$C_{G}(g)=\{x\in G:gx=xg\}.$
Note that $C_{G}(g)$ is a subgroup of $G$. The commutator of two elements $x,y$ is denoted as $[x,y]$, that is, $[x,y]=xyx^{-1}y^{-1}$. Note that $x$ and $y$ commute if and only if $[x,y]=e$. Denote by $\mathbb{Z}_{n}$ the cyclic group with order $n$. For a prime $p$, an elementary abelian $p$-group is a direct product of cyclic groups of order $p$. A Sylow $p$-subgroup of $G$ is a $p$-subgroup whose index is coprime to $p$.
If $G$ is a finite group, the following conditions are equivalent:
  • (a)
    $G$ is nilpotent;
  • (b)
    every Sylow subgroup of $G$ is normal in $G$;
  • (c)
    $G$ is the direct product of its Sylow subgroups.
By $F(G)$ we will denote the Fitting subgroup of $G$, i.e. the largest normal nilpotent subgroup of $G$.
A central product $H\circ K$ of two groups $H$ and $K$ with respect to an isomorphism $\theta$ from a central subgroup $Z$ of $H$ to a central subgroup of $K$ is defined to be
$(H\times K)/\{(z,z\theta):z\in Z\}.$
Given a positive integer $n$, we use $D_{2n}$ to denote the dihedral group with order $2n$, a presentation of $D_{2n}$ is given by
$D_{2n}=\langle a,b:a^{n}=b^{2}=e,bab=a^{-1}\rangle.$ (1)
Note that $D_{2n}$ is an abelian group if and only if $n\in\{1,2\}$, and for all $1\leq i\leq n$,
$o(a^{i}b)=2,~{}~{}D_{2n}=\langle a\rangle\cup\{b,ab,a^{2}b,\ldots,a^{n-1}b\}.$ (2)
Suppose that $A$ is an abelian group. The generalized dihedral group corresponding to $A$, denoted by $D(A)$, is the external semidirect product of $A$ with the cyclic group of order two, with the non-identity element acting as the inverse map on $A$. Viewing this external semidirect product as an internal semidirect product, $A$ is an abelian normal subgroup of index two. A presentation for $D(A)$ is as follows:
$D(A)=\langle A,x:o(x)=2,xgx=g^{-1}\hbox{ for each $g\in A$}\rangle.$ (3)
Observation 1
Let $D(A)$ be the generalized dihedral group as presented in (3). Then we have the following.
  • (i)
    Every element of $D(A)\setminus A$ is an involution;
  • (ii)
    $D(A)$ is abelian if and only if $A$ is an elementary abelian $2$-group, if and only if $D(A)$ is an elementary abelian $2$-group;
  • (iii)
    If $A$ is cyclic of order $n$, then $D(A)\cong D(\mathbb{Z}_{n})\cong D_{2n}$.
Observation 1(ii) means that $D(A)$ is non-abelian if and only if $A$ has an element of order at least $3$. Moreover, Observation 1(iii) implies that the dihedral groups are special cases of generalized dihedral groups. As an example, one can check that $D(\mathbb{Z}_{2}\times\mathbb{Z}_{6})\cong\mathbb{Z}_{2}\times D_{12}$.
Given a positive integer $m$ at least $2$, Johnson [44, pp. 44–45] introduced the generalized quaternion group $Q_{4m}$ of order $4m$, that is,
$Q_{4m}=\langle x,y:x^{m}=y^{2},x^{2m}=e,y^{-1}xy=x^{-1}\rangle.$ (4)
Note that for any $1\leq i\leq 2m$, $o(x^{i}y)=4$ and $Z(Q_{4m})=\{e,x^{m}\}$.
A group $G$ is called a Frobenius group if there is a proper subgroup $H$ of $G$ such that $H\cap H^{g}=1$ whenever $g\in G\setminus H$. Let
$K=\{1_{G}\}\cup(G\setminus(\cup_{g\in G}H^{g}))$
be the Frobenius kernel of $G$. It is well-known (see, for example, [3, 35.24 and 35.25]) that $K\trianglelefteq G$, $G=K\rtimes H$, $C_{G}(h)\leq H$ for each $h\in H$, and $C_{G}(k)\leq K$ for each $k\in K$. Moreover, by the Thompson theorem on finite groups with fixed-point-free automorphisms of prime order [74, Theorem 1], $K$ is nilpotent. Frobenius groups are also characterised as transitive non-regular permutation groups in which non-identity elements fix at most one point.
Every graph considered is simple: that is, it is undirected, with no loops and no multiple edges. Given a graph $\Gamma$, the vertex set and edge set of $\Gamma$ are denoted by $V(\Gamma)$ and $E(\Gamma)$, respectively. Let $\mathcal{F}$ be a collection of finite graphs. If no induced subgraph of $\Gamma$ is isomorphic to a graph in $\mathcal{F}$, then $\Gamma$ is called $\mathcal{F}$-free; and if =$\mathcal{F}=\{\Delta\}$, we say that $\Gamma$ is $\Delta$-free. Sometimes we say instead that $\Gamma$ forbids $\mathcal{F}$ or $\Delta$. Denote by $P_{n}$ and $C_{n}$ the path of length $n-1$ and the cycle of length $n$, respectively. For distinct $x,y\in V(\Gamma)$, if $x$ is adjacent to $y$ in $\Gamma$, then we use $x\sim_{\Gamma}y$ to denote this edge $\{x,y\}$, shortly by $x\sim y$. In particular, denote by
$x_{1}\sim x_{2}\sim\cdots\sim x_{n}$
the induced path of length $n-1$, which is a subgraph of $\Gamma$ induced by the set $\{x_{1},x_{2},\ldots,x_{n}\}$, and is isomorphic to $P_{n}$ with $\{x_{i},x_{i+1}\}\in E(\Gamma)$ for all $11\leq i\leq n-$ 1. Similarly, in $\Gamma$, the induced cycle of length $n$ by $\{x_{1},x_{2},\ldots,x_{n}\}$ is an induced subgraph, and is denoted by
$x_{1}\sim x_{2}\sim\cdots\sim x_{n}\sim x_{1},$
where $\{x_{1},x_{n}\}\in E(\Gamma)$ and $\{x_{i},x_{i+1}\}\in E(\Gamma)$ for all $11\leq i\leq n-$ 1.
The neighbourhood of a vertex $v$ is $N(v)=\{w\in V(\Gamma):w\sim v\}$, and the closed neighbourhood of $v$ is $\bar{N}(v)=\{v\}\cup N(v)$. A vertex $v$ of a graph $\Gamma$ is said to be dominating if $v\sim w$ for all $w\in V(\Gamma)\setminus\{v\}$.
Part I Recognising the commuting graph
3 The Recognition Problem
The main focus of this part is the recognition problem of commuting graphs. It is basically an algorithmic question. Given an undirected graph $X=(V,E)$ as input, the aim is to design an algorithm that efficiently checks if there is a group $G$ with $|V|$ elements such that the input graph $X$ is isomorphic to $\Gamma(G)$, and if so to efficiently determine such an isomorphism.
We summarize our results.
  • (a)
    We obtain a deterministic polynomial-time algorithm for the case of extraspecial $p$-groups for odd primes $p$. Extraspecial $p$-groups are a special case of $p$-groups of nilpotence class $2$. A formal definition is in Section 7.
  • (b)
    For the general case, we obtain a quasipolynomial-time algorithm. The algorithm is essentially based on a simple $2^{O(\log^{3}n)}$ time algorithm that enumerates the multiplication tables of all distinct groups of order $n$ (where by distinct we mean mutually non-isomorphic groups of order $n$). This enumeration algorithm combined with Babai’s quasipolynomial time algorithm for graph isomorphism [4] yields a simple $2^{O(\log^{3}n)}$ time algorithm to check if a given $n$ vertex graph $X$ is the commuting graph of an $n$-element group: for each $n$-element group $G$ that is output by the enumeration algorithm we check if $\Gamma(G)$ is isomorphic to $X$ by using Babai’s isomorphism test [4].
  • (c)
    A natural question in connection with our algorithm for recognizing the commuting graphs of extraspecial groups is whether groups with the same commuting graphs are isoclinic. This holds for extraspecial groups which is exploited by the algorithm, and it is natural to conjecture that this property holds for all groups of nilpotence class 2 (of which extraspecial groups are a subclass). We present counter-examples for class-3 nilpotent groups and conjecture that the property holds for class-2 nilpotent groups.
  • (d)
    Additionally, we also show an efficient reduction of the commuting graph recognition problem to the recognition problem of commuting graphs of indecomposable groups. We also note recognizing the commuting graph of generalized dihedral groups along with an extension to Frobenius groups.
Some related work.  There is a result by Giudici and Kuzma [37] that shows the following: every $n$-vertex graph $X$ with at least two vertices of degree $n-1$ is realizable as the commuting graph of a semigroup. It is easy to see that their construction actually gives a polynomial-time algorithm for finding a semigroup with $n$ elements and a bijection from it to the vertex set of $X$ such that the edges of $X$ realize the commuting relation of the semigroup. It is a nearly complete answer to the question in the semigroups setting.
Solomon and Woldar [69] have shown that the commuting graph $\Gamma(G)$ of a finite simple group $G$ uniquely determines the group. More precisely, they have shown that for any finite group $H$ and any finite simple group $G$, their commuting graphs $\Gamma(H)$ and $\Gamma(G)$ are isomorphic if and only if the groups $G$ and $H$ are isomorphic. An interesting question is whether or not the proof of this result [69] yields a polynomial-time algorithm for checking if a given graph $X$ is the commuting graph of a simple group.
4 Basic properties
We being with some preliminary observations that are well-known in the literature (see, e.g., the survey [21]).
Let $G$ be a finite group. What are the cliques of $\Gamma(G)$? If a vertex subset $S$ induces a clique in $\Gamma(G)$ then $S$ is a commuting subset of elements of $G$. Conversely, every commuting subset of elements of $G$ forms a clique in $\Gamma(G)$. Which cliques of $\Gamma(G)$ correspond to subgroups of $G$? Although we cannot directly infer the group multiplication from $\Gamma(G)$, we can observe the following.
Lemma 4.1
A vertex subset $S$ is a maximal clique of $X$ iff $S$ is a maximal abelian subgroup of $G$.
Proof 1
Suppose $H$ is a maximal abelian subgroup of $G$. Clearly, $H$ is a clique in $X$. If $x\notin H$ is adjacent to all of $H$ then $x$ commutes with all of $H$ implying that $\langle H\cup\{x\}\rangle$ is an abelian subgroup of $G$ strictly larger than $H$. Hence the clique induced by $H$ is maximal. Conversely, by a similar argument, if $S$ is a maximal clique in $X$ then $S$ is a maximal abelian subgroup of $G$$\qed$
4.1 The vertex degrees of $\Gamma(G)$ and conjugacy classes of $G$
Let $X=(G,E)$ be the commuting graph of a finite group $G$. Let $x^{G}$ denote the conjugacy class of $x\in G$:
$x^{G}=\{g^{-1}xg\mid g\in G\},$
which is the orbit of $x$ under the conjugation action of $G$ on itself. Let $\deg(v)$ denote the degree of a node $v$ in the graph $X$. The orbit-stabilizer lemma [19] directly implies the following.
Proposition 4.2
For each $x\in G$ its centralizer $C_{G}(x)$ is the closed neighborhood $\bar{N}(x)$ of $x$ in $\Gamma(G)$, and
$|C_{G}(x)|=1+\deg(x)={\frac{|G|}{|x^{G}|}},\textrm{ for all }x\in G.$
Let $n=|G|$, let $m$ denote the number of edges in the commuting graph $\Gamma(G)$, and let $k$ denote the number of conjugacy classes. As $\sum_{x\in G}\deg(x)=2m$, we have:
$2m+n=\sum_{x\in G}{\frac{|G|}{|x^{G}|}}=|G|\cdot\sum_{x\in G}{\frac{1}{|x^{G}|}}=n\cdot k.$ (5)
Thus the number of conjugacy classes of $G$ is $k=(2m+n)/n$, which, by the above equation, can be inferred from the commuting graph. Thus, for example, the only regular commuting graphs are the complete graphs, which are the commuting graphs of abelian groups.
Let $f(n)$ denote the minimum number of conjugacy classes in a group of order $n$. The function $f(n)$ is quite well studied. Landau [50] showed that $f(n)\to\infty$ as $n\to\infty$. The first lower bound was obtained by Brauer [11] and Erdős and Turán [33], who showed that $f(n)\geq\log\log n$ (logarithms to base $2$). This was improved to $\epsilon\log n/(\log\log n)^{8}$ by Pyber [61]. The exponent $8$ in the previous expression was reduced to $7$ by Keller [47], and then to $3+\epsilon$ by Baumeister, Maróti and Tong-Viet [8], for any $\epsilon>0$. It is conjectured that a bound of the form $f(n)\geq C\log n$ holds for some constant $C$.
This is relevant to the recognition problem because of the following observation, which follows immediately from equation (5).
Proposition 4.3
An $n$-vertex graph $X$ with fewer than $n(f(n)-1/2)$ edges cannot be the commuting graph of a group, where $f(n)$ is the minimum number of conjugacy classes that an $n$ element group can have.
At the other extreme, we can rule out very dense incomplete graphs by the $5/8$-theorem [41] for finite groups. More precisely, for any finite nonabelian group $G$ with $n$ elements:
$|C|=|\{(x,y)\mid x,y\in G\textrm{ and }xy=yx\}|\leq\frac{5}{8}\cdot n^{2},$
where $C\subset G\times G$ denotes the set of ordered pairs $(x,y)$ of commuting elements. Then, in the commuting graph $\Gamma(G)=(G,E)$ the number of edges
$|E|=\frac{|C|-n}{2}\leq\frac{5}{16}\cdot n^{2}-\frac{n}{2}\leq\frac{5}{8}\cdot{n\choose 2}.$
Putting it together, we have:
Proposition 4.4
An incomplete graph $X=(V,E)$ cannot be the commuting graph of a group if $|E|>\frac{5}{8}\cdot{|V|\choose 2}$.
4.2 The commuting graph and maximal abelian subgroups
For a finite group $G$, let $\mathcal{M}\subseteq 2^{G}$ denote the set of all maximal abelian subgroups of $G$. Associated with $G$ is the natural hypergraph $(G,\mathcal{M})$, where the hyperedges are precisely the maximal abelian subgroups of $G$.
Proposition 4.5
The commuting graph of a finite group determines the hypergraph of its maximal abelian subgroups, and, conversely the maximal abelian subgroups hypergraph of the group determines its commuting graph.
Proof 2
Let $G$ be a finite group. The maximal abelian subgroups are the maximal cliques in $\Gamma(G)$; and the edges of $\Gamma(G)$ are the $2$-element subsets of maximal abelian subgroups.$\qed$
Remark 1
Since the commuting graph $\Gamma(G)$ of a finite simple group $G$ uniquely determines the group [69], by Proposition 4.5 it follows that the set of maximal abelian groups of a finite simple group $G$ uniquely determines $G$.
For a finite group $G$, the number of maximal abelian subgroups is bounded by (${|G|}\choose{\log|G|}$ because every subgroup of $G$ has a generating set of size bounded by $\log|G|$. Thus, the hypergraph of maximal abelian subgroups has size at most $n^{\log n}$ for $n$ element groups.
This simple bound is tight apart from a constant in the exponent. To see this, consider the extraspecial group $G$ of order $p^{2n+1}$ and exponent $p$, where $p$ is an odd prime. The centre has order $p$, and $G/Z(G)$ is isomorphic to a $2n$-dimensional vector space $V$ over the field $F$ of $p$ elements, with the bilinear form from $V\times V$ to $F$ corresponding to the commutation map from $G/Z(G)\times G/Z(G)$ to $Z(G)$. Maximal abelian subgroups contain the centre, and correspond to maximal totally isotropic subspaces of $V$. It is known that the number of such subspaces is $\prod_{i=1}^{n}(p^{i}+1)$ (see [73]). As
$\prod_{i=1}^{n}(p^{i}+1)>p^{n(n+1)/2}>p^{n(2n+1)/4}=|G|^{n/4},$
it follows that the number of maximal abelian subgroups of the extraspecial $p$-group $G$ is $|G|^{\Omega(\log|G|)}$ for any odd prime $p$.
5 Commuting graphs of product groups
A natural approach to the problem of recognizing commuting graphs is to examine the connection between product decompositions of the commuting graph $\Gamma(G)$ and the direct product decomposition of the group $G$.
Let $G$ and $H$ be finite groups. Consider the commuting graph $\Gamma(G\times H)$ of their direct product. We recall the definition of strong products of graphs.
Definition 1
[42, Chapter 4, Page 35] Let $X=(V,E)$ and $X^{\prime}=(V^{\prime},E^{\prime})$ be simple undirected graphs. Their strong product, denoted $X\boxtimes X^{\prime}$, is a simple undirected graph with the cartesian product $V\times V^{\prime}$ as its vertex set and edges defined as follows: distinct pairs $(u,u^{\prime})$ and $(v,v^{\prime})$ are adjacent if and only if one of the following holds:
  • (a)
    $u=v$ and $(u^{\prime},v^{\prime})\in E^{\prime}$,
  • (b)
    $u^{\prime}=v^{\prime}$ and $(u,v)\in E$,
  • (c)
    $(u,v)\in E$ and $(u^{\prime},v^{\prime})\in E^{\prime}$.
The following proposition is immediate from the definition.
Proposition 5.1
For finite groups $G$ and $H$
$\Gamma(G\times H)=\Gamma(G)\boxtimes\Gamma(H).$
The strong product is both commutative and associative [42, Chapter 4], in the sense that $X\boxtimes Y\cong Y\boxtimes X$ and $(X\boxtimes Y)\boxtimes Z)\cong X\boxtimes(Y\boxtimes Z)$.
By a nontrivial graph we mean a graph with more than one vertex. Given a graph $X$, a factorization $X\cong Y\boxtimes Z$ is nontrivial if both $Y$ and $Z$ are nontrivial. A graph $X$ is said to be a prime graph if it does not have a nontrivial factorization. (We warn the reader that this term is also used for graphs defined by arithmetic parameters of a finite group; see for example [53].)
It turns out that any connected graph $X$ has a unique factorization (up to isomorphism) into a strong product of prime graphs [42, Theorem 7.14, Pg 85]. Moreover, this prime factorization can be computed in polynomial time [34]. As a consequence, we obtain the following polynomial-time reduction.
Recall the a group $G$ is said to be indecomposable if it cannot be expressed as the direct product of two non-trivial groups.
Theorem 5.2
The problem of recognizing the commuting graphs of groups is polynomial-time reducible to the problem of recognizing the commuting graphs of indecomposable groups.
Proof 3
Suppose we have an algorithm $\mathcal{A}$ for recognizing the commuting graphs of indecomposable groups. Using $\mathcal{A}$ as subroutine, we present a polynomial-time algorithm for recognizing the commuting graphs of all finite groups.
Let $X=(V,E)$ be an undirected graph on $n$ vertices which is a purported commuting graph.
First, using the polynomial-time algorithm of Feigenbaum and Schäffer [34] we can factorize $X$ as
$X=X_{1}\boxtimes X_{2}\boxtimes\cdots\boxtimes X_{k},$
where each $X_{i}$ is a prime graph on at least two vertices. It follows that $k\leq\log n$. Now, for each subset $S\subseteq[k]$ of the prime graphs we combine them by taking the strong direct product to define the graph
$X_{S}=\boxtimes_{i\in S}X_{i}.$
Notice that any order in which the strong product of these graphs $X_{i}$ is computed yields the same graph, up to isomorphism.
Thus, in time which is polynomial in $n$, we have computed graphs $X_{S}$ for each subset $S$ of $[k]$. Now, we invoke the subroutine $\mathcal{A}$ that checks if $X_{S}$ is the commuting graph of an indecomposable group $G_{S}$, and if so, finds a labeling of the vertices of $X_{S}$ with elements of $G_{S}$ consistent with the commuting relations.
We can now check if the input graph $X$ is the commuting graph of a group with a straightforward dynamic programming strategy based on the following claim which directly follows from Proposition 5.1.
Claim 5.3
For any two disjoint subsets $S,S^{\prime}$ of $[k]$ such that $X_{S}$ and $X_{S^{\prime}}$ are the commuting graphs of groups $G_{S}$ and $G_{S^{\prime}}$ the graph $X_{S}\boxtimes X_{S^{\prime}}$ is the commuting graph of the direct product group $G_{S}\times G_{S^{\prime}}$.
Now, the algorithm works in stages, computing subsets $S$ of $[k]$ along with the graph $X_{S}$ and group $G_{S}$ such that $X_{S}=\Gamma(G_{S})$.
  1. (a)
    for stages 00 to $k$ do
  2. (b)
    Stage 00   we have subsets $S$ such that $G_{S}$ is an indecomposable group. We mark all such subsets $S$ as true. We mark the remaining subsets as false.
  3. (c)
    Stage $i+1$   For each pair of disjoint subsets $S$ and $S^{\prime}$ marked true in Stages $1,2,\ldots,i$, such that $S\cup S^{\prime}$ is marked false, we mark $S\cup S^{\prime}$ true and compute $X_{S\cup S^{\prime}}=X_{S}\boxtimes X_{S^{\prime}}$ and $G_{S\cup S^{\prime}}=G_{S}\times G_{S^{\prime}}$.
  4. (d)
    end-for
  5. (e)
    If $[k]$ is marked true then the input $X$ is the commuting graph of the group $G_{[k]}$ computed above.
The above description checks if $X$ is the commuting graph of a group with at most $2^{k}\leq n$ calls to the subroutine $\mathcal{A}$ and the running time of the remaining computation is clearly polynomially bounded in $n$$\qed$
From Theorem 5.2 we can easily deduce from the Solomon–Woldar theorem [69] that direct products of simple groups too have uniquely determined commuting graphs.
Corollary 5.4
If $G$ is a direct product of simple groups, then $G$ $\Gamma(G)$ is uniquely determined by its commuting graph.
We do not know if the commuting graph of an indecomposable group $G$ is necessarily a prime graph. However, if $G$ is simple then it turns out, as we show below, that $\Gamma(G)$ is a prime graph.
Lemma 5.5
The commuting graph of a finite simple group is a prime graph under the strong product.
Proof 4
We use the fact that, if $G$ is a non-abelian simple group and $g\in G\setminus\{1\}$, then there exists $h\in G$ such that $\langle g,h\rangle=G$ [40]. Now if $\langle g,h\rangle=G$, then
  • (a)
    $g$ and $h$ are nonadjacent in the commuting graph (since $G$ is non-abelian);
  • (b)
    $g$ and $h$ have no non-identity common neighbour in the commuting graph (since a common neighbour would belong to $Z(G)$, but $Z(G)=\{1\}$.
Now suppose for a contradiction that $\Gamma(G)$ is the strong product of two nontrivial graphs with vertex sets $A$ and $B$ (i.e., with $|A|>1$ and $|B|>1$). Then we can identify $G$ with the Cartesian product $A\times B$. Suppose that $(a,b)$ is the identity of $G$. Then $(a,b)$ is joined to all other vertices in the commuting graph.
It follows that $a$ is joined to every other vertex in $A$, and $b$ to every other vertex in $B$; therefore, for all $x\in A\setminus\{a\}$, $y\in B\setminus\{b\}$, the three vertices $(a,y)$, $(x,b)$ and $(x,y)$ are adjacent to each other in the commuting graph $\Gamma(G)$.
Choose $y\in B\setminus\{b\}$, and suppose that $\langle(a,y),(u,v)\rangle=G$. Now
  • (a)
    if $u=a$ then $(a,y)$ and $(a,v)$ are both joined to $(x,b)$ for any $x\in A\setminus\{a\}$;
  • (b)
    if $v=b$ then $(a,y)$ and $(u,b)$ are joined;
  • (c)
    if neither of the above, then $(a,y)$ and $(u,v)$ are both joined to $(u,b)$.
Each case is contradictory; so our assumption that $\Gamma(G)$ is the strong product of two nontrivial graphs is false, and the theorem is proved.$\qed$
Remark 2
The proof of Lemma 5.5 only requires that $Z(G)=1$ and that any non-identity element is contained in a $2$-element generating set. These assumptions are valid in any almost simple group $G$ with simple normal subgroup $S$ such that $G/S$ is cyclic [17].
As a consequence of Lemma 5.5, we obtain the following variant of Theorem 5.2.
Theorem 5.6
The problem of recognizing the commuting graph of a direct product of simple groups is polynomial-time reducible to the problem of recognizing the commuting graph of a simple group.
The algorithm for the reduction in the above theorem is much simpler and more efficient than in Theorem 5.2. For we only need to consider the prime factors $X_{i}$, in the factorization $X=\boxtimes_{i=1}^{k}X_{i}$, and not all the strong products $X_{S}$ for arbitrary subsets $S\subseteq[k]$.
6 Commuting graphs of semidirect products
In this section we explore whether we can recognize the commuting graphs of semidirect products $G=H\rtimes K$ if $H$ and $K$ are both from group classes whose commuting graphs are easily recognizable.
For example, consider the commuting graph of the dihedral group $D_{2n}$. Let $D_{2n}=\langle a,b\rangle$ where $a^{2}=1,b^{n}=1$ and $aba^{-1}=b^{-1}$. For $n$ odd, $1$ is the only dominating vertex, there is an $n$-clique corresponding to $\langle b\rangle$ and the $ab^{i}$ are pendant vertices. For $n$ even, it is a bit different with $b^{n/2}$ being the other dominating vertex.
More generally, consider the generalized dihedral group $\langle A,t\rangle$ defined earlier, where $A$ is abelian of twice odd order. Its commuting graph is isomorphic to that of the dihedral group of the same order. We conclude that we can recognise that a graph is the commuting graph of a generalized dihedral group, but cannot decide which one.
Seeking a generalization of this example we consider Frobenius groups.
6.1 Commuting graphs of Frobenius groups
In this subsection we demonstrate that we can recognise from its commuting graph that a group $G$ is a Frobenius group. These groups were defined earlier. Let $G$ be a Frobenius group with kernel $N$ and complement $H$.
Thompson [74]proved that the Frobenius kernel is nilpotent, and Zassenhaus [78] worked out the detailed structure of a Frobenius complement. Everything we need about Frobenius groups is explained in standard group theory texts  [59, 43, 67].
In the commuting graph of a group $G$, the identity is a dominating vertex (that is, joined to all others); indeed, any vertex in the centre is dominating, so if $Z(G)$ is non-trivial then the commuting graph is $2$-connected.
Lemma 6.1
  1. (a)
    A Frobenius kernel has non-trivial centre.
  2. (b)
    A Frobenius complement has non-trivial centre.
Proof 5
(a) By Thompson’s theorem, a Frobenius kernel is nilpotent, and nilpotent groups have a non-trivial centre.
(b) Let $H$ be a Frobenius complement. By [43, Theorem 6.3, Corollary 6.10], if $H$ has even order, then it has a unique element of order $2$, which lies in the centre; while, if $H$ has odd order, then it has a unique subgroup of order $p$ for every $p$ dividing $|H|$. In the latter case, choose $p$ to be the smallest prime divisor of $H$, and $P=\langle a\rangle$ the unique subgroup of order $p$. Then $P$ is a normal subgroup of $H$. For any $h\in H$, we have $a^{h}=a^{m}$ for some $m$ with $11\leq m\leq p-$ 1. If $m>1$, let $q$ be a prime divisor of $m$; then some power of $h$ has order $q$, a contradiction. So $a^{h}=a$ for all $h\in H$, whence $a\in Z(H)$$\qed$
Theorem 6.2
Let $G$ be a group of order $nk$, where $n,k>1$ and $\gcd(n,k)=1$, and let $\Gamma$ be the commuting graph of $G$. Then $G$ is a Frobenius group with Frobenius complement of order $k$ if and only if $\Gamma$ satisfies the following conditions:
  • (a)
    there is a dominating vertex $v$;
  • (b)
    $\Gamma\setminus\{v\}$ has a component of size $n-1$ and $n$ components of size $k-1$, and each component has a dominating vertex.
Proof 6
Suppose that $G$ is a Frobenius group with kernel $N$ of order $n$ and complement $H$ of order $k$. The identity is a dominating vertex. Also non-identity vertices in the kernel do not commute with non-identity vertices in a complement, and non-identity vertices in different complements do not commute with each other. So the components of $\Gamma\setminus\{1\}$ are as stated, and the lemmas above show that they have dominating vertices.
Conversely, suppose that the commuting graph $\Gamma$ of $G$ has properties (a) and (b). If $C$ is a component with a dominating vertex $c$, then the centralizer $C_{G}(c)$ is equal to $C\cup\{1\}$, which is thus a subgroup of $G$. Let $N$ be the subgroup containing the component of size $n-1$, and let $H_{1},\ldots,H_{n}$ be the subgroups containing the other components.
Since $\Gamma$ is invariant under automorphisms of $G$, we have an action of $G$ on the set $\Omega$ of components of $\Gamma\setminus\{v\}$ of size $k-1$ by conjugation. We show that this action satisfies the conditions for a Frobenius group given above.
Choose a prime $p$ dividing $k$. Since $p$ does not divide $n$, the subgroup $N$ cannot contain a Sylow $p$-subgroup; but each of $H_{1},\ldots,H_{n}$ contains such a subgroup. By the conjugacy part of Sylow’s theorem, $G$ acts transitively on $\Omega$.
A non-identity element of $G$, acting by conjugation, fixes itself, and so fixes the component containing it. We must show that it fixes no other component. Count fixed point of elements of $G$. The identity fixes $n$; non-identity elements of $N$ fix a positive number; and the remaining elements fix a non-negative number. So the sum of the fixed point numbers is at least $n+n(k-1)=nk$. But, by the Orbit-counting Lemma, this sum is equal to $|G|=nk$, since $G$ is transitive. So equality holds; non-identity elements of $N$ fix no point, and the remaining non-identity elements fix exactly one point each. So $G$ is a Frobenius group.$\qed$
Suppose $X$ is a graph given as input such that $X=\Gamma(G)$ for some group $G$. Then Theorem 6.2 gives us a polynomial-time test for checking if the group $G$ is a Frobenius group.
Remark 3
Above we assumed that our graph is a commuting graph. However, it is easy to construt a graph satisfying the conditions of Theorem 6.2 which is not a commuting graph. Can we find necessary and sufficient conditions for $\Gamma$ to be the commuting graph of a Frobenius group?
7 Recognizing commuting graphs of extraspecial groups
Our aim in this section is to design a polynomial-time algorithm that takes a simple undirected graph $X=(V,E)$, with $n=p^{2r+1}$ vertices, $r\geq 2$ and $p$ an odd prime, as input and determines if $X$ is the commuting graph of an extraspecial group $G$ of order $p^{2r+1}$. Moreover, the algorithm will also compute a certifying bijection $v\mapsto g_{v}$ labeling each vertex $v\in V$ by a unique group element $g_{v}\in G$ such that the commuting relations hold.
We recall that a $p$-group $G$ is extraspecial group $G$ if its centre $Z(G)$ is of order $p$ and coincides with its derived subgroup $G^{\prime}$ and the Frattini subgroup $\Phi(G)$ [43, Chapter 4, pp. 123]. Furthermore, it is known that there are exactly two non-isomorphic extraspecial groups of order $p^{2r+1}$, for each $r\geq 1$ and each prime $p$. These two groups, which we denote as $G_{1}$ and $G_{2}$, are described below. For a pair of elements $(x,y)$ in any group, let $[x,y]$ denote their commutator $x^{-1}y^{-1}xy$.
  • (a)
    The first group $G_{1}$ is generated by $12r+$ 1 elements $G_{1}=\langle z,x_{i},y_{i},1\leq i\leq r\rangle$ such that $[x_{i},x_{j}]=1$, $[y_{i},y_{j}]=1$ for all $i,j$, and $[x_{i},y_{j}]=1$ for all $i\neq j$. $[x_{i},y_{i}]=z$ for all $i$, and $x_{i}^{p}=y_{i}^{p}=z^{p}=1$ for all $i$. The center $Z(G_{1})=\langle z\rangle$.
  • (b)
    The second group $G_{2}$ also has $12r+$ 1 generators $G_{2}=\langle z,x_{i},y_{i},1\leq i\leq r\rangle$ such that $[x_{i},x_{j}]=1$, $[y_{i},y_{j}]=1$ for all $i,j$, and $[x_{i},y_{j}]=1$ for all $i\neq j$. And $[x_{i},y_{i}]=z$ for all $i$, and $x_{i}^{p}=1=z^{p}$, $y_{i}^{p}=z$ for all $i$. The center $Z(G_{2})=\langle z\rangle$.
Each element $g\in G_{1}\setminus\{1\}$ has order $p$. On the other hand, each element $g\in G_{2}\setminus\{1\}$ has order either $p$ or $p^{2}$. Hence, $G_{1}$ and $G_{2}$ are non-isomorphic.
7.1 Isoclinism
The lexicographic product of graphs $\Gamma_{1}$ and $\Gamma_{2}$ is the graph on the vertex set $V(\Gamma_{1})\times V(\Gamma_{2})$, in which $(v_{1},w_{1})$ and $(v_{2},w_{2})$ are joined if and only if either $v_{1}$ and $v_{2}$ are joined in $\Gamma_{1}$, or $v_{1}=v_{2}$ and $w_{1}$ and $w_{2}$ are joined in $\Gamma_{2}$. If $\Gamma_{2}$ is a complete graph, then the lexicographic product is obtained from $\Gamma_{1}$ by blowing up each vertex to a complete graph on $|V\Gamma_{2}|$ vertices.
Commutation in a group $G$ is a map from $G\times G$ to the derived group $G^{\prime}$; multiplying the arguments by elements of $Z(G)$ does not change the function value, so commutation induces a map from $G/Z(G)\times G/Z(G)$ to $G^{\prime}$. We say that groups $G$ and $H$ are isoclinic if there are isomorphisms $\theta:G/Z(G)\to H/Z(H)$ and $\phi:G^{\prime}\to H^{\prime}$ which intertwine the commutation map; that is, for all $g_{1},g_{2}\in G$,
$[\theta(g_{1}Z(G)),\theta(g_{2}Z(G))]=\phi([g_{1}Z(G),g_{2}Z(G)]).$
Now the commuting graph of a group $G$ is the lexicographic product of a graph on $G/Z(G)$ with $g_{1}Z(G)$ and $g_{2}Z(G)$ joined if $[g_{1},g_{2}]=1$ (this graph can also be described as the induced subgraph of the commuting graph of $G$ on a transversal for $Z(G)$ in $G$) with a complete graph on $Z(G)$; the graph on $G/Z(G)$ is determined by the isoclinism type. Hence, as observed by Pezzott and Nakaoka [60], we have:
Theorem 7.1
Isoclinic groups of the same order have isomorphic commuting graphs.
Next we observe that the two extraspecial $p$-groups of the same order $p^{2a+1}$ are isoclinic. In each case, $G/Z(G)$ is elementary abelian of order $p^{2a}$, which can be identified with a vector space $V$ of dimension $2a$ over the $p$-element field $F$; the center has order $p$ and can be identified with $F$. Commutation is a non-degenerate alternating bilinear form on $V$, that is, a map $b$ from $V\times V$ to $F$ which is bilinear and satisfies
  • (a)
    $b(v,v)=0$ for all $v\in V$;
  • (b)
    $b(v,w)=-b(w,v)$ for all $v,w\in V$;
  • (c)
    if $b(v,w)=0$ for all $w\in V$, then $w=0$.
The induced subgraph of $\Gamma_{G}$ on a transversal for $Z(G)$ can be identified with the graph on $V$, where $v$ and $w$ are joined if and only if $b(v,w)=0$. We call this the symplectic graph on $V$. (A symplectic form is a non-degenerate alternating bilinear form.)
Now there is a unique symplectic form on any even-dimensional vector space [73]. We deduce that the two extraspecial groups are isoclinic, and their commuting graphs are isomorphic.
Now we can recognise in polynomial time that a graph has $p^{2a+1}$ vertices and is the lexicographic product of a graph $\Gamma$ on $p^{2a}$ vertices (having a dominating vertex) with a complete graph on $p$ vertices as follows:
  • (a)
    Check that the number of vertices is $p^{2a+1}$.
  • (b)
    Compute the equivalence relation on the graph where two vertices are equivalent if they have the same closed neighbourhood. (Recall that the closed neighbourhood of $v$ consists of $v$ together with the vertices joined to $v$.)
  • (c)
    Check that there is one equivalence class of size $p$, consisting of the dominating vertices, and the other equivalence classes all have size $p^{2}-p$.
  • (d)
    Now we can compute the graph $\Gamma$ by choosing $|C|/p$ vertices from each equivalence class $C$. This gives a graph with $p^{2a}$ vertices.
It remains to recognise the symplectic graph. There are two possible strategies we could adapt. The first is to build, from the symplectic graph, a vector space with a bilinear form so that we have explicitly identified the unknown graph. We were able to produce a polynomial-time algorithm for this, which we hope to publish elsewhere; but in the interests of brevity we adopt a different strategy; our algorithm will simply decide if our unknown graph is isomorphic to a symplectic graph. The strategy is to go from the vector space to the projective space and use the Buekenhout–Shult theorem [16] to recognise the graph. For reference, we state the theorem in the finite case.
Theorem 7.2
Let $(P,L)$ be a finite point-line geometry, with $P$ and $L$ both non-empty. Assume
  1. (a)
    two points lie on at most one line;
  2. (b)
    no point is collinear with all others;
  3. (c)
    if a point $p$ is not on a line $l$, then $p$ is joined to one or all points of $l$.
Then either $(P,L)$ is a generalized quadrangle, or it is the geometry of points and totally isotropic (or totally singular) lines in a classical polar space over a finite field.
We refer to the paper for the undefined terms in this theorem.
Now we proceed in a number of steps.
Step 1
The case $a=1$. A non-abelian group of order $p^{3}$ is the union of $p+1$ maximal abelian subgroups of order $p^{2}$ intersecting pairwise in the identity; so our original graph is the commuting graph of a non-abelian group of order $p^{3}$ if and only if the graph $\Gamma$ consists of $p+1$ pairwise disjoint cliques of size $p-1$ together with a dominating vertex.
Step 2
Moving to projective space. For this we remove the dominating vertex (which would correspond to the zero vector) and choose one from each equivalence class of size $p-1$ for the remaining vertices. We obtain a graph on $(p^{2a}-1)/(p-1)$ vertices.
Step 3
Construct the lines. If $u$ and $v$ are adjacent vertices of $\Gamma$, the line joining them is
$L(u,v)=\{u,v\}\cup\{w:\bar{N}(u)\cap\bar{N}(v)\subseteq\bar{N}(w)\}.$
If a line contains fewer than $p+1$ points, the algorithm terminates and returns false.
Step 4
Check the Buekenhout–Shult hypotheses. Return false if they fail.
We are left with the two possible conclusions of Buekenhout and Shult. The first of these is that the geometry is a generalized quadrangle. (This will only occur if $a=2$.) In this case, the points of the quadrangle are all the points of the $3$-dimensional projective space over the $p$-element field $F$. The closed neighbourhoods are the planes of the projective space. It is clear that the closed neighbourhood of a vertex contains the line through any two of its points. So it suffices to show that all closed neighbourhoods have size $p^{2}+p+1$ to identify the projective space. If so, then the map $v\leftrightarrow\bar{N}(v)$ is a null polarity of the projective space, and so defines a symplectic polarity; the algorithm can return true.
In the other case, there is still one problem to resolve. Inspecting the list of classical polar spaces, we see that there is another with the same parameters as the symplectic space on a $2k$-dimensional vector space: the orthogonal space, defined by a non-singular quadratic form on a $(2k+1)$-dimensional vector space. In characteristic $2$, they are isomorphic, so there is no more to do. In odd characteristic, the orthogonal space contains a $3\times 3$ grid. Indeed, over the $p$-element field $F$, it contains a $(p+1)\times(p+1)$ grid, the quadric defined by a hyperbolic form (one of Witt index $2$) on a $4$-dimensional vector space. (This is also realised as a regulus in projective $3$-space together with the opposite regulus.) But the symplectic space contains no such grid. For we can take a $4$-cycle with vertices spanned by the vectors $e_{1},e_{2},f_{1},f_{2}$ (where $b(e_{i},e_{j})=b(f_{i},f_{j})=0$ and $b(e_{i},f_{j})=\delta_{ij}$). Choose arbitrary points on the edges of the cycle so that opposite points are collinear: say the points spanned by $e_{1}+ae_{2}$, $f_{2}-af_{1}$, $e_{2}+bf_{1}$, $e_{1}+bf_{2}$. The determinant of their coefficients is $2ab\neq 0$, so the two lines are skew.
Thus we can simply check whether such a grid exists. Transitivity of the group of the symplectic space on $4$-cycles shows that only one $4$-cycle needs to be checked.
We leave to the reader the verification that all these tests can be done in polynomial time.
8 A quasipolynomial time algorithm in the general case
Assuming CFSG (the classification of finite simple groups), in this section we present a $2^{O(\log^{3}n)}$ time algorithm for checking if a given $n$-vertex graph is the commuting graph of some $n$-element group and, if so, labeling the vertices of the graph by the group elements consistent with all the commuting pairs.
  1. (a)
    The main idea is the notion of description $\mathop{\mathrm{desc}}(G)$ for any finite group $G$. It is essentially an encoding of the group $G$ as a binary string of length at most $c\cdot\log^{3}|G|$, for some fixed constant $c>0$, from which $G$ can be recovered efficiently. More precisely, we will describe an algorithm $\mathcal{D}$ that takes $\mathop{\mathrm{desc}}(G)$ as input and computes the multiplication table of $G$ as output in time polynomial in $|G|$.
  2. (b)
    This algorithm $\mathcal{D}$ can be used to generate the multiplication tables of all groups $G$ of order $n$ as follows. We enumerate all possible binary strings $w$ of length at most $c\cdot\log^{3}n$. For each $w$ we run the algorithm $\mathcal{D}$ on it. If $w=\mathop{\mathrm{desc}}(G)$ for some $G$ then $\mathcal{D}$ will output its multiplication table. We compute $\Gamma(G)$ from $G$ and check if $X$ is isomorphic to $\Gamma(G)$ in time $2^{O(\log^{3}n)}$ using Babai’s graph isomorphism algorithm [4].
Thus, the main task of this section is to define $\mathop{\mathrm{desc}}(G)$ and describe the algorithm $\mathcal{D}$. We will first define $\mathop{\mathrm{desc}}(G)$ for finite simple groups, along with how the algorithm $\mathcal{D}$ will compute the multiplication table for $G$ given $\mathop{\mathrm{desc}}(G)$ as input. We will then extend the notion of $\mathop{\mathrm{desc}}(G)$ to all groups by using a composition series for $G$ combined with $\mathop{\mathrm{desc}}(N)$ for each composition factor $N$ of $G$.
Remark 4
We emphasize here that our notion of “short” description is entirely different from the well-studied notion of generator-relator presentations of groups (see e.g. [5]). Indeed, $\mathop{\mathrm{desc}}(G)$ and its construction will be significantly simpler. It suffices for the purpose of recovering the multiplication table of the group $G$ in time polynomial in $|G|$.
8.1 Short descriptions for finite simple groups assuming CFSG
We first briefly summarize the classification theorem of finite simple groups (CFSG) [39].
  • (a)
    Cyclic groups of prime order.
  • (b)
    Alternating groups $A_{n},n\geq 5$.
  • (c)
    Finite simple groups of Lie type.
  • (d)
    Sporadic finite simple groups.
As there are only finite many sporadic finite simple groups, for the algorithm $\mathcal{D}$ they can be trivially dealt with. The cyclic groups of prime order are also trivial.
Lemma 8.1
If $G$ is cyclic of prime order $p$ then $\mathop{\mathrm{desc}}(G)$ can be defined as the binary encoding $\mathop{\mathrm{bin}}(p)$ of $p$ which is of length $\log p$ bits. We can include $1$ as the generator with the additive group $\mathbb{Z}_{p}$ as the group it encodes.
For several of the remaining cases, the following simple lemma gives a framework for defining $\mathop{\mathrm{desc}}(G)$ and computing $G$’s multiplication table from it. We use the well-known fact that any finite simple group can be generated with two generators (see [55]).
Lemma 8.2
Let $H$ be a finite group and $N\lhd K\leq H$ be subgroups such that
  • (a)
    $\mathop{\mathrm{desc}}(H)$ is of length $m$.
  • (b)
    The multiplication table for $H$ can be computed from $\mathop{\mathrm{desc}}(H)$ in time polynomial in $|H|$.
  • (c)
    We can test membership of elements in $N$ in time polynomial in $|H|$.
Suppose $G=K/N$ is simple. Then $\mathop{\mathrm{desc}}(G)$ has size $O(m+\log|H|)$ and in time polynomial in $|H|$ we can compute the multiplication table for $G$.
Proof 7
As $G=K/N$ is simple it can be generated by two elements $Ng_{1},Ng_{2}\in K/N$ where $g_{1},g_{2}\in K$. Let $\mathop{\mathrm{desc}}(G)$ be the indices of these two elements in $H$ along with $\mathop{\mathrm{desc}}(H)$. (By “indices” here we mean the positions of these elements in the description.) Clearly, it is of size $O(\log|H|)$. The claimed algorithm will first use $\mathop{\mathrm{desc}}(H)$ to compute the multiplication table for $H$. It can enumerate $H$ and use the indices to identify $g_{1}$ and $g_{2}$. Using the multiplication table for $H$ it can generate the subgroup $K=\langle g_{1},g_{2}\rangle$, listing down all its elements in time polynomial in $|H|$. Next, the algorithm can identify the normal subgroup $N$ (as given in the hypothesis) in time polynomial in $|H|$ and obtain the multiplication table for $G=K/N$$\qed$
We will apply the above lemma to the remaining cases of finite simple groups.
Lemma 8.3
If $G=A_{n}$ for some $n\geq 5$ then we can define $\mathop{\mathrm{desc}}(G)$ of size $O(\log|G|)$.
Proof 8
We can define $\mathop{\mathrm{desc}}(A_{n})$ as the encoding of two generators of $A_{n}$, by giving the index of the two generating permutations, say $g_{1},g_{2}$ in $S_{n}$ in the natural order of enumerating $S_{n}$. Hence, $\mathop{\mathrm{desc}}(G)$ will be of size $2\log n!=O(\log|A_{n}|)$. We can now apply Lemma 8.2 with $H=S_{n}$, $K=A_{n}$ and $N=\{1\}$ to complete the proof.$\qed$
We now consider the finite simple groups of Lie type. It is convenient to deal with them in two categories: classical and exceptional. We will follow the (concrete) constructions of these groups as explained in Wilson’s text [77].
8.1.1 Classical Lie type
Lemma 8.4
For each finite simple group $G$ that is of classical Lie type, there is a description $\mathop{\mathrm{desc}}(G)$ of $O(\log|G|)$ size such that the multiplication table for $G$ can be computed from $\mathop{\mathrm{desc}}(G)$ in time polynomial in $|G|$.
Proof 9
These finite simple groups are categorized as follows:
  1. (a)
    $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$ for either $n>2$ or $q>3$. These are the projective special linear group arising from the matrix group $\mathop{\mathrm{GL}}\nolimits_{n}(q)$ (the finite group of $n\times n$ invertible matrices over the finite field $\mathbb{F}_{q}$).
  2. (b)
    $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$ which is the projective symplectic group, obtained as quotient of the symplectic group $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$, and where $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ itself is a subgroup of $\mathop{\mathrm{GL}}\nolimits_{2m}(q)$.
  3. (c)
    $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ which is the projective unitary group, obtained as a quotient of the special unitary group $\mathop{\mathrm{SU}}\nolimits_{n}(q)$. The general unitary group $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ itself is a subgroup of the matrix group $\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})$.
  4. (d)
    Finally, the orthogonal groups which are also obtained essentially as quotients of matrix groups, but there are different subcases depending on both $n$ and $q$.
Consider $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$. The special linear group $\mathop{\mathrm{SL}}\nolimits_{n}(q)=\{A\in\mathop{\mathrm{GL}}\nolimits_{n}(q)\mid\det(A)=1\}$ is a subgroup of $\mathop{\mathrm{GL}}\nolimits_{n}(q)$ of index $q-1$. $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$ is obtained as the quotient $\mathop{\mathrm{SL}}\nolimits_{n}(q)/Z$ where $Z$ consists of all “scalar matrices” $\lambda I_{n}$ of determinant $1$. Thus, $|\mathop{\mathrm{PSL}}\nolimits_{n}(q)|\geq|\mathop{\mathrm{GL}}\nolimits_{n}(q)|/q^{2}$. We can apply Lemma 8.2, letting $H=\mathop{\mathrm{GL}}\nolimits_{n}(q)$, $K=\mathop{\mathrm{SL}}\nolimits_{n}(q)$ and $N=Z$ in it. Notice that $\mathop{\mathrm{GL}}\nolimits_{n}(q)$ has a description of size $O(\log n+\log q)$. Given $n$ and $q$ we need to enumerate $\mathop{\mathrm{GL}}\nolimits_{n}(q)$ in time polynomial in $|\mathop{\mathrm{GL}}\nolimits_{n}(q)|$. The algorithm will enumerate all $n\times n$ matrices $M$ over $\mathbb{F}_{q}$, one by one, run a polynomial time algorithm for checking if $M$ is invertible (say, using Gaussian elimination with $O(n^{3})$ many field operations) and if so include it in $\mathop{\mathrm{GL}}\nolimits_{n}(q)$. Since a large fraction of matrices are invertible (for $n$ more than some constant), this procedure will run in time polynomial in $|GL_{n}(q)|$. Thus, by Lemma 8.2 we can obtain the desired description for $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$.
Next, consider $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$. Recall that the symplectic group $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ is the subgroup of $\mathop{\mathrm{GL}}\nolimits_{2m}(q)$ consisting of all isometries of a nonsingular alternating bilinear form on the vector space $\mathbb{F}_{q}^{2m}$. It turns out that $|\mathop{\mathrm{Sp}}\nolimits_{2m}(q)|>q^{m^{2}}$. As $|\mathop{\mathrm{GL}}\nolimits_{2m}(q)|=q^{4m^{2}}$, clearly
$|\mathop{\mathrm{GL}}\nolimits_{n}(q)|\leq|\mathop{\mathrm{Sp}}\nolimits_{2m}(q)|^{4}.$
Thus, the description of the group $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ is easy to construct. Given $2m$ and $q$ in binary, in time polynomial in $m$ and $q$ we can define a nonsingular alternating bilinear form $\langle$,$\rangle$ , ⟩ on $\mathbb{F}_{q}^{2m}$. Then, as done for the previous case, we can enumerate the elements of $\mathop{\mathrm{GL}}\nolimits_{2m}(q)$, one by one, and put each into $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ if it is an isometry for $\langle$,$\rangle$ , ⟩. Now, the simple group $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$ by quotienting out the subgroup of scalar matrices (which has size at most $2$). Clearly, Lemma 8.2 can be applied to obtain the desired description for $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$ along with the algorithm for obtaining its multiplication table.
For the next case $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$, we start with the group $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ which is the subgroup of $\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})$ that preserves the sesquilinear form $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}^{q}$. From there on the process of obtaining $\mathop{\mathrm{SU}}\nolimits_{n}(q)$ and then $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ is exactly like in the case of $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$. In order to obtain the desired description for $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ it suffices to generate $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ in time polynomial in $|\mathop{\mathrm{PSU}}\nolimits_{n}(q)|$ given $n$ and $q$. Clearly, $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ can be generated in time polynomial in $|\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})|$ by looking all these matrices. However, $|\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})|<|\mathop{\mathrm{GU}}\nolimits_{n}(q)|^{3}$. Thus, the naive algorithm for generating all of $\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})$ and picking those which preserve the sesquilinear form will again work. Finally, the group $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ is only smaller than $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ by a small power of $q$.
The fourth category are the simple groups arising from the orthogonal groups. There are different cases to consider here.
When $q$ is odd, there are exactly two non-singular $n$-variate symmetric bilinear forms over $\mathbb{F}_{q}$. If $f$ is one of them then $\mathop{\mathrm{GO}}\nolimits_{n}(f)$ is the subgroup of $GL_{n}(q)$ that leaves $f$ invariant. If $n=2m+1$ is odd, it turns out that there is only one orthogonal group which we may denote by $\mathop{\mathrm{GO}}\nolimits_{2m+1}(q)$. We can then construct $\mathop{\mathrm{SO}}\nolimits_{2m+1}(q)$ and $\mathop{\mathrm{PSO}}\nolimits_{2m+1}(q)$ analogous to the earlier cases. But there is a further subgroup of index $2$ in $\mathop{\mathrm{PSO}}\nolimits_{2m+1}(q)$ denoted $\mathop{\mathrm{P}\Omega}\nolimits_{2m+1}(q)$. As before, we can apply Lemma 8.2 to obtain the desired description of $\mathop{\mathrm{P}\Omega}\nolimits_{2m+1}(q)$ from which the group’s multiplication table can be obtained in polynomial time.
For $n=2m$, there are two orthogonal groups $\mathop{\mathrm{GO}}\nolimits^{+}_{2m}(q)$ and $\mathop{\mathrm{GO}}\nolimits^{-}_{2m}(q)$. As in the previous case we can obtain the groups $\mathop{\mathrm{P}\Omega}\nolimits^{\epsilon}_{2m}(q)$ for $\epsilon\in\{+,-1\}$. For odd $q$ these are always simple and Lemma 8.2 is again easily applicable.
The characteristic $2$ case of orthogonal groups requires different techniques. However, Lemma 8.2 is still applicable in defining $\mathop{\mathrm{desc}}(G)$ for the simple groups $G$ arising in this case. We omit the details.$\qed$
8.1.2 Exceptional Lie type
We will closely follow Wilson’s exposition [77, Chapter 4]. These finite simple groups come in ten families:
  • (a)
    The five families of Chevalley groups $G_{2}(q)$, $F_{4}(q)$, and $E_{n}(q),n\in\{6,7,8\}$.
  • (b)
    The two families of Steinberg–Tits–Herzig groups ${}^{3}{D}_{4}(q)$ and ${}^{2}{E}_{6}(q)$.
  • (c)
    The Suzuki and Ree groups $Sz(2^{2n+1})$, $R_{1}(3^{2n+1})$ and $R_{2}(2^{2n+1})$ (also known as ${}^{2}B_{2}(q)$, ${}^{2}G_{2}(q)$ and ${}^{2}F_{4}(q)$ for the appropriate $q$).
Unlike the classical case, these families have only one parameter which is the field size $q$. It turns out that the following lemma (analogous to Lemma 8.2) captures how these simple groups arise and suggest a suitable $\mathop{\mathrm{desc}}(G)$ for each $G$ along with the corresponding algorithm for recovering $G$ from $\mathop{\mathrm{desc}}(G)$.
Lemma 8.5
Let $G$ be finite simple group from any of the ten families listed above. Then we can define a description $\mathop{\mathrm{desc}}(G)$ of size $O(\log q)$ such that $G$’s multiplication table can be obtained from it in time polynomial in $|G|$.
Proof 10
We will briefly describe how the groups are constructed for each of these ten families. The easy descriptions and algorithms for generating their multiplication tables will suggest themselves.
  1. (a)
    $\mathop{\mathrm{Sz}}\nolimits(2^{2n+1})$:  Let $q=2^{2n+1}$. The Suzuki group $Sz(q)$ is a subgroup of the symplectic group $\mathop{\mathrm{Sp}}\nolimits_{4}(2^{2n+1})$. It arises as the centralizer of a particular automorphism of order $2$ in $\mathop{\mathrm{Sp}}\nolimits_{4}(2^{2n+1})$. As $|\mathop{\mathrm{Sz}}\nolimits(q)|=(q^{2}+1)q^{2}(q-1)$ $|\mathop{\mathrm{GL}}\nolimits_{4}(q)|\leq|\mathop{\mathrm{Sz}}\nolimits(q)|^{4}$. Thus, we can cycle through all elements of $\mathop{\mathrm{Sp}}\nolimits_{4}(q)$ looking for the elements of this centralizer. The automorphism is also easily described (in size independent of $q$).
  2. (b)
    $G_{2}(q)$:  It turns out that $G_{2}(q)$ arises as the automorphism group of the octonion algebra over $\mathbb{F}_{q}$. The entire octonion algebra can be listed in time $O(q^{8})$. As the automorphisms permute fix the identity in the octonion algebra, $G_{2}(q)$ is essentially a subgroup of $\mathop{\mathrm{GO}}\nolimits_{7}(q)$. We can enumerate $\mathop{\mathrm{GO}}\nolimits_{7}(q)$ and pick out the automorphisms which will give us $G_{2}(q)$. As $|G_{2}(q)|$ is $q^{6}(q^{6}-1)(q^{2}-1)$, this will be polynomial time in the group size. We can define the $\mathop{\mathrm{desc}}(G_{2}(q))$ of size $O(\log q)$. It is essentially just the encoding of $q$.
  3. (c)
    $R_{1}(3^{2n+1})={}^{2}G_{2}(3^{2n+1})$: This can be handled quite analogously to $G_{2}(q)$ [77, Section 4.5].
  4. (d)
    ${}^{3}{D}_{4}(q)$:  This group arises as centralizer of a certain order $3$ automorphism of $\mathop{\mathrm{P}\Omega}\nolimits^{+}_{8}(q^{3})$. Because of the special nature of this automorphism (product of an $\mathbb{F}_{q}$ automorphism and the triality automorphism), it has a short description of size $O(\log q)$. As $\mathop{\mathrm{P}\Omega}\nolimits^{+}_{8}(q^{3})$ can be enumerated in time polynomial in $q$, we can pick out the centralizer elements in time polynomial in $q$.
  5. (e)
    $F_{4}(q)$:  when $q$ is not a power of $2$ or $3$, $F_{4}(q)$ arises as automorphisms of the finite (non-associative) Albert algebra which is $27$-dimensional. Thus, $F_{4}(q)$ can be picked out from $GL_{27}(q)$ by enumerating it and checking each element for being an automorphism. The finite Albert algebra can be defined in $O(\log q)$ size by defining the multiplication of its basis elements.
  6. (f)
    $R_{2}(2^{2n+1})={}^{2}F_{4}(2^{2n+1})$:  This case is quite similar to the previous one [77, Section 4.9].
  7. (g)
    $E_{6}(q)$:  This group can be defined as automorphisms of a certain trilinear form over the Albert algebra. After quotienting out the scalar subgroup from this automorphism group we will obtain $E_{6}(q)$. The same algorithm of enumerating the bigger group and selecting the automorphisms and then quotienting out the scalars will work.
  8. (h)
    $E_{7}(q)$ and $E_{8}(q)$:  Both these arise as automorphisms of a finite Lie algebra. For $E_{8}(q)$ we have to consider a $248$ dimensional Lie algebra. Thus we can look for $E_{8}(q)$ inside $GL_{248}(q)$ by enumerating all its elements and picking out the automorphisms. All we need is that from a $O(\log q)$ size description we should be able to obtain the Lie algebra. For that we only need to express the multiplication of basis elements as an $\mathbb{F}_{q}$ linear combination of the basis (which is of size $248$). $E_{7}(q)$ is similarly dealt with as automorphisms of a smaller dimensional Lie algebra.$\qed$
The above lemmas immediately imply the following.
Theorem 8.6
There is an algorithm that, given as input $\mathop{\mathrm{desc}}(G)$ for a finite simple group $G$, outputs its multiplication table in time polynomial in $|G|$.
8.2 Short descriptions for finite groups
We first show that every finite group $G$ has a description $\mathop{\mathrm{desc}}(G)$ of bit-length $O(\log^{3}|G|)$. Then we describe the algorithm that given as input a bit string $y$, if $y=\mathop{\mathrm{desc}}(G)$ for some finite group $G$ then the algorithm runs in time $2^{O(\log^{3}|G|)}$ and outputs the multiplication table of a group $G^{\prime}$ isomorphic to $G$.
Building on the construction of $\mathop{\mathrm{desc}}(G)$ for finite simple groups $G$, described in Section 8.1, we obtain $\mathop{\mathrm{desc}}(G)$ for all finite groups with a straightforward construction that uses a composition series for $G$. It is convenient to use Erdős-Rényi cube generating sequences [32].
Definition 2 (Erdős-Rényi cube generating sequences)
Let $G$ be a finite group. A cube generating sequence (abbreviated as c.g.s.) for $G$ is a sequence $(g_{1},g_{2},\ldots,g_{k})$ of elements from $G$ such that
$G=\{g_{1}^{\epsilon_{1}}g_{2}^{\epsilon_{2}}\cdots g_{k}^{\epsilon_{k}}\mid\epsilon_{i}\in\{0,1\}\hbox{ for each }i\}.$
Thus, given the cube generating sequence $(g_{1},g_{2},\ldots,g_{k})$, each element of $G$ can be defined by the bit vector $(\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{k})$. We note that this bit vector need not be unique for a group element.
Theorem 8.7
[32] Every finite group $G$ has a cube generating sequence of size at most $2\log|G|$.
Let $X=(g_{1},g_{2},\ldots,g_{k})$ be cube generating sequence for a group $G$. Its length $k$ is denoted $|X|$. Abusing notation, by $g\in X$ we will mean that $g=g_{i}$ for some $i$.
We note that cube generating sequences can be computed in polynomial time.
Lemma 8.8
There is a polynomial-time algorithm that takes a group $G$ by its multiplication table as input and outputs a cube-generating sequence of size $2\log|G|$ for $G$.
Proof 11
The simple algorithm is directly based on a “cube doubling” trick due to Babai and Szemerédi [6]. Suppose we have picked the sequence $(g_{1},g_{2},\ldots,g_{i})$ of group elements from $G$. Let $C_{i}$ be the cube it generates. That is,
$C_{i}=\{g_{1}^{\epsilon_{1}}g_{2}^{\epsilon_{2}}\cdots g_{i}^{\epsilon_{k}}\mid\epsilon_{j}\in\{0,1\}\hbox{ for each }j\}.$
For any $g\in G$ let $C_{i}g=\{xg\mid x\in C_{i}\}$. We can enumerate all the elements of $C_{i}$ and $C_{i}g$ with at most $2i\cdot 2^{i}$ group multiplications, for $i\leq\log|G|$, which is is polynomial in $|G|$.
Now, if $C_{i}g\cap C_{i}\neq\emptyset$ for $g\in G$ then $g$ is in the cube generated by the sequence $(g_{1}^{-1},g_{2}^{-1},\ldots,g_{i}^{-1},g_{1},\ldots,g_{i})$. Thus, if for all $g\in G$ $C_{i}g\cap C_{i}\neq\emptyset$ then we have a cube generating sequence of length $2i$. Otherwise, there is a $g_{i+1}\in G$ such that
$C_{i}g_{i+1}\cap C_{i}=\emptyset,$
and the cube generated by $(g_{1},g_{2},\ldots,g_{i+1})$ is twice the size of $C_{i}$ by a simple counting argument. Clearly, we can find $g_{i+1}$ in time polynomial in $|G|$ by searching through $G$$\qed$
We now explain the construction of $\mathop{\mathrm{desc}}(G)$ for a finite group $G$, where $G$ is given by its multiplication table. Compute a composition series for $G$ (which can be done in time polynomial in $|G|$ by computing normal closures of elements):
$1=N_{0}\lhd N_{1}\lhd\cdots\lhd N_{r}=G.$
For each quotient group $H_{i}=N_{i}/N_{i-1}$ for $1\leq i\leq r$, we can compute a cube generating sequence $Y_{i}=(N_{i-1}g_{i,1},N_{i-1}g_{i,2},\ldots,N_{i-1}g_{i,k_{i}})$ of length $k_{i}\leq 2\log|H_{i}|$ in polynomial time, by Lemma 8.8. For each $i$, let $X_{i}=(g_{i,1},g_{i,2},\ldots,g_{i,k_{i}})$ be the corresponding sequence of coset representatives.
Proposition 8.9
The sequence $X=(X_{1},X_{2},\cdots,X_{i})$ obtained by concatenating the sequences $X_{j},1\leq j\leq i$, defined above, in that order is a cube generating sequence for the subgroup $N_{i}$ of length at most $\sum_{j=1}^{i}2\log|H_{j}|=2\log|N_{i}|$, for $1\leq i\leq r$. In particular, $X=(X_{1},X_{2},\cdots,X_{r})$ is a cube generating sequence for $G$ of length at most $2\log|G|$.
Proof 12
Suppose $H\lhd G$ and $Y=\{y_{1},y_{2},\ldots,y_{\ell}\}$ is a cube generating sequence for $H$ and $Z=\{Hz_{1},Hz_{2},\ldots,Hz_{m}\}$ is a cube generating sequence for $G/H$. We note that any element $Hg\in G/H$ can be expressed as
$Hg=Hz_{1}^{\epsilon_{1}}z_{2}^{\epsilon_{2}}\cdots z_{m}^{\epsilon_{m}},\epsilon_{i}\in\{0,1\},$
which implies that any element of $G$ can be expressed as
$y_{1}^{\delta_{1}}y_{2}^{\delta_{2}}\cdots y_{\ell}^{\delta_{\ell}}z_{1}^{\epsilon_{1}}z_{2}^{\epsilon_{2}}\cdots z_{m}^{\epsilon_{m}},\,\epsilon_{i},\delta_{j}\in\{0,1\}.$
Now, building on this, an easy inductive argument yields the proposition.$\qed$
For $1\leq i\leq r$, let $Y_{i}=(X_{1},X_{2},\ldots,X_{i})$. By the above proposition, $Y_{i}$ is a cube generating sequence for the group $N_{i}$ of length at most $2\log|N_{i}|$. As $N_{i-1}\lhd N_{i}$, for each $x\in X_{i}$ we have $xN_{i-1}=N_{i-1}x$, $1<i\leq r$. In particular, for $y\in Y_{i}$ there is a unique element $w(x,y)\in N_{i-1}$ such that we have
$xy=w(x,y)x,$ (6)
where $w[x,y]$ is a word over the cube generating sequence $Y_{i-1}$. Thus, $w(x,y)$ is a word of length at most $2\log|N_{i-1}|$, expressed using the c.g.s. $Y_{i-1}$. Let $R_{i}$ consist of the set of all such equations defined by Equation (6) for $y\in Y_{i-1}$ and $x\in X_{i}$. Notice that $R_{i}$ has $|Y_{i-1}|\cdot|X_{i}|$ equations in it and each word $w(x,y)$ over $Y_{i-1}$ is defined by the bit vector of exponents in the cube product over $Y_{i-1}$. Thus, the size of $R_{i}$ is
$\mathop{\mathrm{size}}(R_{i})=O(|Y_{i-1}|\cdot|X_{i}|\cdot|Y_{i-1}|)=O((\log|N_{i-1}|)^{2}\cdot\log|H_{i}|).$
We are now ready to define $\mathop{\mathrm{desc}}(G)$ for the finite group $G$. It is defined as an efficient [^2] We can encode an $r$-tuple of $3$-tuples in binary by using some standard encoding which is polynomial time computable and its inverse is also polynomial time computable. binary encoding of the $r$-tuple defined below
$\mathop{\mathrm{desc}}(G)~{}=~{}[(\mathop{\mathrm{desc}}(H_{1}),X_{1},R_{1}),(\mathop{\mathrm{desc}}(H_{2}),X_{2},R_{2}),\ldots,(\mathop{\mathrm{desc}}(H_{r}),X_{r},R_{r})],$
where $R_{1}$ is the empty set. The size of $\mathop{\mathrm{desc}}(G)$ is clearly
$O(\sum_{i=1}^{r}\log~{}|H_{i}|)+O(\sum_{i=1}^{r}\mathop{\mathrm{size}}(X_{i}))+O(\sum_{i=1}^{r}\mathop{\mathrm{size}}(R_{i})),$
where $\mathop{\mathrm{size}}(X_{i})$ and $\mathop{\mathrm{size}}(R_{i})$ denote the bit length of their encodings described above. Notice that $\sum_{i=1}^{r}\log~{}|H_{i}|=\log|G|$. It follows that $\sum_{i=1}^{r}\mathop{\mathrm{size}}(X_{i})\leq O(\log^{2}|G|)$, because $|X_{i}|\leq 2\log|H_{i}|\leq 2\log|G|$ and each element of $X_{i}$ requires at most $\log|H_{i}|$ bits to encode it. We now bound $\sum_{i=1}^{r}\mathop{\mathrm{size}}(R_{i})$.
$\sum_{i=1}^{r}\mathop{\mathrm{size}}(R_{i})$ $=$ $O(\sum_{i=1}^{r}(\log^{2}|N_{i-1}|\cdot\log|H_{i}|)$
$\leq$ $\sum_{i=1}^{r}O(\log^{2}|G|)\cdot\log~{}|H_{i}|$
$=$ $O(\log^{2}|G|)\cdot\sum_{i=1}^{r}\log~{}|H_{i}|$
$=$ $O(\log^{3}|G|).$
The above discussion has the following immediate consequence.
Lemma 8.10
Given as input the multiplication table of a finite group $G$, there is an algorithm that computes in time polynomial in $|G|$ its description $\mathop{\mathrm{desc}}(G)$.
However, we are interested in the converse statement, about generating $G$ from its description $\mathop{\mathrm{desc}}(G)$ in time polynomial in $|G|$. The next theorem, which shows this, will allow us to enumerate the multiplication tables of all groups of order $n$ upto isomorphism in time $2^{O(\log^{3}n)}$.
Theorem 8.11
There is an algorithm that given as input $\mathop{\mathrm{desc}}(G)$, for a finite group $G$, outputs its multiplication table in time polynomial in $|G|$. As a consequence, there is an algorithm that outputs the list of multiplication tables of all distinct groups of order $n$ in time $2^{O(\log^{3}n)}$.
Proof 13
As argued above, every finite group $G$ has a description $\mathop{\mathrm{desc}}(G)=[(\mathop{\mathrm{desc}}(H_{1}),X_{1},R_{1}),(\mathop{\mathrm{desc}}(H_{2}),X_{2},R_{2}),\ldots,(\mathop{\mathrm{desc}}(H_{r}),X_{r},R_{r})]$ where the $r$-tuple is encoded in a binary string of length bounded by $c\cdot\log^{3}|G|$ for some constant $c>0$ (which is independent of $G$).
Let $w$ be a binary string. If it encodes $\mathop{\mathrm{desc}}(G)$ for some group $G$ such that $|G|=n$, then
  • (a)
    $|w|\leq c\log^{3}n$ and $w$ decodes to an $r$-tuple of $3$-tuples
    $[(\mathop{\mathrm{desc}}(H_{1}),X_{1},R_{1}),(\mathop{\mathrm{desc}}(H_{2}),X_{2},R_{2}),\ldots,(\mathop{\mathrm{desc}}(H_{r}),X_{r},R_{r})].$
  • (b)
    Moreover, for some composition series $1=N_{0}\lhd N_{1}\lhd\cdots\lhd N_{r}=G$ of $G$, $\mathop{\mathrm{desc}}(N_{i})$ is the binary encoding of the $i$-tuple
    $\mathop{\mathrm{desc}}(N_{i})=[(\mathop{\mathrm{desc}}(H_{1}),X_{1},R_{1}),(\mathop{\mathrm{desc}}(H_{2}),X_{2},R_{2}),\ldots,(\mathop{\mathrm{desc}}(H_{i}),X_{i},R_{i})],$
    for $1\leq i\leq r$, where $H_{i}=N_{i}/N_{i-1}$.
The algorithm will compute the multiplication table for $N_{i}$, inductively for increasing values of $i,1\leq i\leq r$. For $i=1$, it can be obtained from $\mathop{\mathrm{desc}}(H_{1})$ by Theorem 8.6.
Given the multiplication table for $N_{i-1}$ and the triple $(\mathop{\mathrm{desc}}(H_{i}),X_{i},R_{i})$ we first obtain $H_{i}$ by Theorem 8.6, and the actual elements of the c.g.s. $X_{i}=(x_{1},x_{2},\ldots,x_{\ell})$ for $\ell\leq 2\log|H_{i}|$. The c.g.s. for $N_{i-1}$ is $Y_{i}=(X_{1},X_{2},\ldots,X_{i-1})=(y_{1},y_{2},\ldots,y_{k})$ for some $k\leq 2\log|N_{i-1}|$. Applying Equation (6) we have as part of $R_{i}$
$x_{\mu}y_{\nu}=w(x_{\mu},y_{\nu})x_{\mu},$ (7)
for each $\mu\in[\ell]$ and each $\nu\in[k]$. Here, note that $w(x_{\mu},y_{\nu})$ is expressed as a product $y_{1}^{e_{1}}y_{2}^{e_{2}}\cdots y_{k}^{e_{k}}$ for $(e_{1},e_{2},\ldots,e_{k})\in\{0,1\}^{k}$.
We now explain how to obtain the multiplication table for the subgroup $N_{i}$ from the above data in time polynomial in $|N_{i}|$. More precisely, the data comprises of the multiplication table for $N_{i-1}$, the multiplication table for the quotient group $H_{i}=N_{i}/N_{i-1}$ and $R_{i}$ (Equation (7) for all $y_{\nu}$ and $x_{\mu}$).
We first note that $N_{i}=\cup_{x}N_{i-1}x$, where $x$ is in the cube
$C=\{\prod_{i=1}^{\ell}x_{j}^{\varepsilon_{j}}\mid\varepsilon_{j}\in\{0,1\}\textrm{ and }x_{j}\in X_{i}\}.$
Thus, $N_{i}=N_{i-1}C$. So, any element in $N_{i}$ is of the form $ab,a\in N_{i-1},b\in C$. We need to express the product of any two elements $a_{1}b_{1},a_{2}b_{2}\in N_{i}$, where $a_{1},a_{2}\in N_{i-1}$ and $b_{1},b_{2}\in C$ in the form $a_{3}b_{3},a_{3}\in N_{i-1},b_{3}\in C$.
Claim 8.12
For any $x_{\mu}\in X_{i}$ and $a\in N_{i-1}$ we can find $a^{\prime}\in N_{i-1}$ such that $x_{\mu}a=a^{\prime}x_{\mu}$ in time polynomial in $|N_{i-1}|$.
Proof of Claim 1
We first express $a$ using the c.g.s. $Y_{i}$ for $N_{i-1}$ to write it as
$a=y_{1}^{\beta_{1}}y_{2}^{\beta_{2}}\cdots y_{\ell}^{\beta_{\ell}}.$
Then we can use Equation (7) $\ell$ times to write $x_{\mu}a=x_{\mu}\prod_{j=1}^{\ell}y_{j}^{\beta_{j}}$ as the product $x_{\mu}a=(\prod_{j=1}^{\ell}w(x_{\mu},y_{j})^{\beta_{j}})x_{\mu}$. We can compute $a^{\prime}=\prod_{j=1}^{\ell}w(x_{\mu},y_{j})^{\beta_{j}}$ using the multiplication table for $N_{i-1}$ as each $w(x_{\mu},y_{j})$ is in $N_{i-1}$.
Using the above claim repeatedly we can express products of the form $xa$ for $x\in C,a\in N_{i-1}$ as $a^{\prime}x$ for som $a^{\prime}\in N_{i-1}$. Finally, we can express the product of any two elements $a_{1}b_{1},a_{2}b_{2}\in N_{i}$, where $a_{1},a_{2}\in N_{i-1}$ and $b_{1},b_{2}\in C$ in the form $a_{3}b_{3},a_{3}\in N_{i-1},b_{3}\in C$. The algorithm runs in time polynomial in $|N_{i-1}|$ and $|X_{i}|$ and hence in time polynomial in $|N_{i}|$.
Applied for increasing values of $i$, finally, for $i=r$ we obtain the table for $N_{r}=G$. The overall algorithm clearly runs in polynomial in $|G|$ time.
The algorithm for enumerating all order $n$ group multiplication tables is easy to obtain from the above. We run through all strings $w$ of length at most $c\log^{3}n$, where $n=|G|$, and apply the above algorithm to generate the multiplication for $G$ if $w$ encodes $\mathop{\mathrm{desc}}(G)$.
Notice that this enumeration may contain duplicates because the same group can have different composition series and hence different descriptions. We can eliminate the duplicates by running Miller’s group isomorphism test [57]. Each such test takes $n^{\log n}$. Since there are at most $2^{\log^{3}n}$ groups being listed the running time to detect multiplicities takes time at most $2^{\log^{3}n}\cdot n^{\log n}\cdot 2^{\log^{3}n}=2^{O(\log^{3}n)}$$\qed$
Remark 5
Some notes concerning the preceding results:
  1. (a)
    Computing a composition series for $G$ can be done in time polynomial in $|G|$ by computing normal closures of elements:
    $1=N_{0}\rhd N_{1}\rhd\cdots\rhd N_{r}=G.$
    More precisely, for each $g\in G$ we compute its normal closure by starting with the cyclic subgroup $N=\langle g\rangle$ and repeatedly replacing it with the (larger) subgroup generated by the set $\{xyx^{-1},x\in G,y\in N\}$ until $N$ does not grow any further. By construction, the final $N$ is a normal subgroup of $G$ and its computation takes time polynomial in $|G|$. If $G$ is not simple, then for some $g\in G$ this construction will yield a nontrivial normal subgroup $N$. We can now recursively find composition series for $N$ and $G/N$ (we have multiplication tables for both) and put them together to obtain a composition series for $G$ in time polynomial in $|G|$.
  2. (b)
    Let $H_{i}=N_{i}/N_{i-1}$ for $1\leq i\leq r$ be the simple quotient groups. Let $n_{i}=|H_{i}|$ for $1\leq i\leq r$ denote their respective orders.
  3. (c)
    For each group $N_{i}$ we compute a cube generating sequence $X_{i}$ of size at most $k_{i}\leq 2\log|N_{i}|$. This computation is done in time polynomial in $|N_{i}|$, by Lemma 8.8. As part of the description of $G$, for each $1\leq i\leq r$ we obtain a set of relations $R_{i}$ as defined in Equation (6).
This algorithm is broadly based on a result of McIver and Neumann [56] that bounds the number of $n$ element groups by $2^{O(\log^{3}n)}$. The theorem is sharpened by subsequent work of Babai et al. [5] on short presentations for finite groups. They show that groups of order $n$ (that do not have a specific finite simple group type, called the Ree groups, as section) have short generator-relator presentations $\langle Y|R\rangle$ of size $O(\log^{3}n)$. The crux of their proof is that all simple groups of $n$ elements (except the Ree groups) have generator-relator presentations of size $O(\log^{2}n)$. Combined with the fact that $n$ element groups have composition series of $\log n$ length gives the $O(\log^{3}n)$ bound.
Theorem 8.13
There is a $2^{O(\log^{3}n)}$ time algorithm that recognizes if a given $n$-vertex graph is the commuting graph of a group of order $n$.
Proof 14
Let $X$ be an $n$-vertex graph. The algorithm cycles through all multiplication tables of distinct groups of order $n$, listed by the $2^{O(\log^{3}n)}$ time algorithm described in Theorem 8.11. For each group $G$ that is listed we can compute its commuting graph $\Gamma(G)$ in polynomial time. We can test if $\Gamma(G)$ is isomorphic to the input graph $X$ using Babai’s $2^{O(\log^{3}n)}$ time algorithm [4].
Thus, the overall computation takes $2^{O(\log^{3}n)}$ time.$\qed$
Remark 6
This algorithm is not very interesting because it does not reveal any new relationship between the group and its commuting graph. Can we do better?
9 Isoclinism of groups and commuting graphs
A property of extraspecial groups that we exploited in the algorithm is that extraspecial groups of the same order are isoclinic. We saw that isoclinic groups of the same order have isomorphic commuting graphs.
What about the converse? That is, if two groups have isomorphic commuting graphs, must they be isoclinic? This holds for various classes of groups, such as abelian groups, nonabelian simple groups, and extraspecial groups (as we have seen).
However, it is not true in general; we recycle an example taken from [23] to show this. Let $G$ be the group of order $64$ which is SmallGroup(64,182) in the SmallGroups library in GAP [36]. The Schur multiplier of $G$ has order $2$, so a Schur cover (a group $H$ of maximal order subject to having a subgroup $Z\leq H^{\prime}\cap Z(H)$ such that $H/Z\cong G$) has order $128$. Moreover, the Bogomolov multiplier of $G$ is equal to the Schur multiplier, which implies that the Schur covers $H$ are commutation-preserving: that is, two elements $a,b\in H$ commute if and only if their projections $Za,Zb\in G$ commute. This implies that the commuting graph of a Schur cover is obtained from the commuting graph of $G$ by replacing each vertex with a clique of size $2$, with all edges between cliques corresponding to adjacent vertices. This procedure also describes the commuting graph of $G\times C_{2}$. On the other hand, it is easy to verify computationally that the derived groups of $H$ and $G\times C_{2}$ are not isomorphic, so these groups cannot be isoclinic.
We note that the group $G$ has nilpotency class $3$, as do all of its Schur covers (these are SmallGroup(128,$i$) for $i\in\{789,790,791,815,816,817\}$ in the GAP library). So the following question is still open:
Conjecture
Is it true that a nilpotent group of class $2$ is determined up to isoclinism by its commuting graph?
Part II Forbidden subgraphs
10 Preliminaries
A number of important graph classes can be defined either structurally or in terms of forbidden induced subgraphs. Here are structural descriptions.
  1. (a)
    $\Gamma$ is called a cograph if it can be built from $1$-vertex graphs by the operations of disjoint union and complementation.
  2. (b)
    $\Gamma$ is called a chordal graph provided that $\Gamma$ contains no induced cycles of length greater than $3$; in other words, any cycle of length at least $4$ contains a chord.
The characterization by forbidden subgraphs is a composite of results from [35, 27, 70, 26].
Proposition 10.1
  1. (a)
    A graph is a cograph if and only if it forbids $P_{4}$.
  2. (b)
    A graph is chordal if and only if it forbids $\{C_{n}:n\geq 4\}$.
Since a cycle of length $5$ or more contains a $4$-vertex path, we have the following observation.
Observation 2
If a graph is $\{P_{4},C_{4}\}$-free, then it is chordal.
We conclude this section with a general remark. In the commuting graph $\Gamma(G)$ of a group $G$, elements of the center $Z(G)$ are dominating vertices; this is the reason why, to investigate questions about connectedness, they are excluded. However, in our case, it doesn’t matter:
Proposition 10.2
  1. (a)
    The class of groups whose commuting graph is $\mathcal{F}$-free is subgroup-closed.
  2. (b)
    If no graph in $\mathcal{F}$ has a dominating vertex, then the class of groups with $\mathcal{F}$-free commuting graph is independent of whether vertices in $Z(G)$ are included or excluded.
Proof 15
(a) This is immediate from the fact that, if $H$ is a subgroup of $G$, then $\Gamma(H)$ is an induced subgraph of $\Gamma(G)$. (We remark that the same result holds for various other graphs such as the power graph, enhanced power graph, nilpotency and solvability graphs. But it does not hold for yet others such as the generating graph or the deep commuting graph.)
(b) If $X$ is an induced subgraph of $\Gamma(G)$, then vertices of $X\cap Z(G)$ would be dominating vertices in $X$. $\qed$
In the cases we are interested in (chordal graphs and cographs), the excluded subgraphs are paths and cycles with more than three vertices. None of these have dominating vertices, so we can take the vertex set of the commuting graph to be all of $G$, but to decide if it has any of these properties we can delete vertices in $Z(G)$.
Also to prove main results of this paper, we need the following assertion which generalizes a result obtained by Wielandt [76] in 1939.
Proposition 10.3
Let $G$ be a finite group. Assume that one of the following statements holds:
  1. (a)
    $Z(G)=1$;
  2. (b)
    $G=G^{\prime}$.
Then $C_{\mathop{\mathrm{Aut}}(G)}(\mathop{\mathrm{Inn}}(G))=1$.
Proof 16
Let $\phi$ be an automorphism of $G$ which commutes with each inner automorphism of $G$. Then for each $g,h\in G$, we have
$\phi(h)^{g}=\phi(h^{g})=\phi(h)^{\phi(g)},$
i. e. $[g,\phi]=g^{-1}\phi(g)\in Z(G)$ and $\phi(g)=g\cdot z_{g}$ for some $z_{g}\in Z(G)$ (depending of $g$).
If $Z(G)=1$, then $z_{g}=1$ for each $g$, therefore $\phi$ is trivial and $C_{\mathop{\mathrm{Aut}}(G)}(\mathop{\mathrm{Inn}}(G))=1$, as required.
Now assume that $G$ is perfect, i.e. $G^{\prime}=G$. Note that $\phi([x,y])=[\phi(x),\phi(y)]=[x,y]$ for each $x,y\in G$. Since $G$ is generated by commutators, we have $z_{g}=1$ for each $g$. Thus, again $\phi$ is trivial and so, $C_{\mathop{\mathrm{Aut}}(G)}(\mathop{\mathrm{Inn}}(G))=1$. $\qed$
11 A structure of groups whose commuting graphs are cographs or chordal graphs
In this section, we will investigate the group-theoretical structure of a group whose commuting graph is a cograph or a chordal graph. In particular, in some sense, the class of such groups generalizes the class of finite solvable groups.
First, we investigate special cases of groups like direct and central products and Frobenuis groups whose commuting graph is a cograph or a chordal graph, and then prove the general structure theorem.
11.1 Direct and central products
Observation 3
Let $A$ be an abelian group. Then $\Gamma(A\times G)$ is a cograph (resp. a chordal graph) if and only if $\Gamma(G)$ is a cograph (resp. a chordal graph).
Lemma 11.1
Let $H$ and $K$ be non-abelian groups. Then $\Gamma(H\times K)$ has an induced subgraph isomorphic to $P_{4}$ and an induced subgraph isomorphic to $C_{4}$.
Proof 17
Now let $a$ and $b$ be non-commuting elements of $G$, and $u$ and $v$ non-commuting elements of $H$. Then
$(a,u)\sim(e_{H},u)\sim(b,e_{K})\sim(e_{H},v)$
is an induced path with four vertices, and
$(a,e_{K})\sim(e_{H},u)\sim(b,e_{K})\sim(e_{H},v)\sim(a,e_{K})$
is an induced cycle of length $4$, as desired. $\qed$
By Lemma 11.1 and Observation 3, the following result holds.
Proposition 11.2
Let $H$ and $K$ be groups. Then
  1. (a)
    $\Gamma(H\times K)$ is a cograph if and only if one of $H$ and $K$ is abelian and the commuting graph of the other is $P_{4}$-free;
  2. (b)
    $\Gamma(H\times K)$ is chordal if and only if one of $H$ and $K$ is abelian and the other is $C_{n}$-free for all $n\geq 4$.
Recall that a finite nilpotent group is the direct product of its Sylow subgroups. Thus, Proposition 11.2 implies the following result.
Corollary 11.3
Let $G$ be a non-abelian nilpotent group. Then $\Gamma(G)$ is a cograph (resp. a chordal graph) if and only if there exists exactly one non-abelian Sylow subgroup in $G$ whose commuting graph is $P_{4}$-free (resp. $C_{n}$-free for all $n\geq 4$), while the other Sylow subgroups are abelian.
We saw in the first part the definition of the strong product of graphs and the fact that the commuting graph of $G\times H$ is $\Gamma(G)\boxtimes\Gamma(H)$ (Proposition 5.1).
We can extend this to central products as follows. Let $\mathfrak{N}_{2}$ be the class of all finite nilpotent groups of nilpotency class at most $2$ (thus, the class consists of finite abelian groups and finite non-abelian groups of nilpotency class $2$).
Lemma 11.4
Let $H$ and $K$ be two groups not in the class $\mathfrak{N}_{2}$. Then $\Gamma(H\circ K)$ is neither a cograph nor a chordal graph.
Proof 18
By assumption, we can choose $a,b\in H$ such that the commutator $[a,b]$ is not in the centre of $H$, and similarly $u,v\in K$. Then
$(a,u)\sim(a,e_{K})\sim(e_{H},v)\sim(b,v)$
carries a subgroup $P_{4}$ of $\Gamma(H\times K)$ such that the projection to $H\circ K$ is an isomorphism, and also
$(a,e_{K})\sim(e_{H},u)\sim(b,e_{K})\sim(e_{H},v)\sim(a,e_{K})$
is a $C_{4}$ subgraph which projects isomorphically to $H\circ K$. $\qed$
11.2 Frobenius groups
In this sections we investigate when the commuting graph of a Frobenius group is a cograph or a chordal graph. We prove the following assertion.
Proposition 11.5
Let $G$ be a Frobenius group with Frobenius complement $H$ and Frobenius kernel $K$ and $\mathcal{F}$ be a collection of graphs such that no graph in $\mathcal{F}$ has a dominating vertex. Then $\Gamma(G)$ is $\mathcal{F}$-free if and only if $\Gamma(K)$ is $\mathcal{F}$-free and $\Gamma(H)$ is $\mathcal{F}$-free.
Proof 19
This is clear from our remarks about dominating vertices and the fact (observed in Theorem 6.2) that $\Gamma(G\setminus\{1\})$ is the disjoint union of $\Gamma(H\setminus\{1\})$ and $|H|$ copies of $\Gamma(K\setminus\{1\})$.
Corollary 11.6
Let $G$ be a Frobenius group with Frobenius complement $H$ and Frobenius kernel $K$. Then $\Gamma(G)$ is a cograph (resp. chordal graph) if and only if $\Gamma(K)$ is a cograph (resp. a chordal graph) and $\Gamma(H)$ is a cograph (resp. a chordal graph).
Question 3
For which Frobenius groups is the commuting graph a cograph or a chordal graph?
Remark 7
Since $K$ is nilpotent, Corollary 11.3 gives a criterion when $\Gamma(K)$ is a cograph (resp. a chordal graph).
11.3 A general structure theorem
The main result of this section is the following theorem.
Theorem 11.7
Let $G$ be a finite group such that $\Gamma(G)$ is a cograph (resp. a chordal graph), and let $F(G)$ be the Fitting subgroup of $G$ (the largest nilpotent normal subgroup of $G$).
  1. (a)
    One of the following statements holds:
    1. (i)
      $C_{G}(F(G))=Z(F(G))$;
    2. (ii)
      $F(G)$ is of nilpotent class at most $2$ and $G/F(G)$ is almost simple; in particular, $G$ has a unique non-abelian composition factor.
  2. (b)
    If $\mathop{\mathrm{Soc}}(G)$ is non-abelian, then $G$ has a normal subgroup $N=A\times S$, where $A$ is abelian, $S$ is simple and $\Gamma(S)$ is a cograph (resp. a chordal graph), such that $G/N$ is isomorphic to a subgroup of $\mathop{\mathrm{Out}}(S)$ and, therefore, $G/N$ is solvable.
Proof 20
Let $F^{*}(G)$ be the generalized Fitting subgroup of $G$. By [3, 31.12],
$F^{*}(G)=F(G)\circ E(G),$ (8)
where $E(G)$ is the subgroup of $G$ generated by the set $\mathop{\mathrm{Comp}}(G)$ of all components (subnormal quasisimple subgroups) of $G$.
Let $H_{1}$ and $H_{2}$ be distinct elements from $\mathop{\mathrm{Comp}}(G)$. By [3, 31.5], $H_{1}$ and $H_{2}$ commute, therefore $\langle H_{1},H_{2}\rangle=H_{1}\circ H_{2}$ is a subgroup of $G$. By Lemma 11.4, $\Gamma(G)$ is neither a cograph nor a chordal graph, a contradiction to Proposition 10.2. Thus, $|\mathop{\mathrm{Comp}}(G)|\leq 1$.
If $|\mathop{\mathrm{Comp}}(G)|=0$, then $F^{*}(G)=F(G)$, and because we have $C_{G}(F^{*}(G))\leq F^{*}(G)$ by [3, 31.13], statement (a)(i) holds.
Assume that $|\mathop{\mathrm{Comp}}(G)|=1$. Then $G$ has a quasisimple characteristic subgroup $N=E(G)$. Consider $C=C_{G}(N)$, which is a normal subgroup of $G$. Let $H=\langle N,C\rangle=N\circ C$. By Proposition 10.2 and Lemma 11.4, $C$ is nilpotent of nilpotency class at most $2$, therefore $C\leq F(G)$. On the other hand, $F(G)\leq C$ by (8). Thus, $C=F(G)$ and, therefore $G/F(G)$ is a subgroup of $\mathop{\mathrm{Aut}}(N)$ which has a simple normal subgroup $H/C\cong N/Z(N)$.
Now we show that $\mathop{\mathrm{Aut}}(N)$ is an almost simple group and $\mathop{\mathrm{Soc}}(\mathop{\mathrm{Aut}}(N))\cong N/Z(N)$. Indeed, $N/Z(N)\cong\mathop{\mathrm{Inn}}(N)$ is a normal subgroup of $\mathop{\mathrm{Aut}}(N)$ and $N/Z(N)$ is simple. By Proposition 10.3, we have
$\mathop{\mathrm{Aut}}(N)\cong\mathop{\mathrm{Aut}}(N)/C_{\mathop{\mathrm{Aut}}(N)}(\mathop{\mathrm{Inn}}(N))$
is a subgroup of $\mathop{\mathrm{Aut}}(\mathop{\mathrm{Inn}}(N))\cong\mathop{\mathrm{Aut}}(N/Z(N))$. So, we have that $\mathop{\mathrm{Aut}}(N)$ is a subgroup of an almost simple subgroup $\mathop{\mathrm{Aut}}(N/Z(N))$ which contains a normal subgroup $\mathop{\mathrm{Inn}}(N)\cong N/Z(N)$. Thus, $\mathop{\mathrm{Aut}}(N)$ is an almost simple group with socle $N/Z(N)$, as required.
So, $G/F(G)$ is isomorphic to a subgroup of $\mathop{\mathrm{Aut}}(N)$ which is almost simple with socle isomorphic to $\mathop{\mathrm{Inn}}(N)$ and contains a subgroup $H/F(G)\cong\mathop{\mathrm{Inn}}(N)$, therefore $G/F(G)$ is almost simple. Thus, statement (a)(ii) holds.
Now we prove statement (b). Assume that $L=\mathop{\mathrm{Soc}}(G)$ is non-abelian. Then $L=A\times S$, where $A$ is abelian and $S$ is the direct product of non-abelian simple groups, moreover, both $A$ and $S$ are normal subgroups of $G$. By Propositions 10.2 and 11.2, $S$ is simple and $\Gamma(S)$ is a cograph (resp. a chordal graph). If $C=C_{G}(S)$, then $C\unlhd G$ and $C\cap S=1$ since $S$ is simple. Thus, $N=SC=S\times C$ is a normal subgroup of $G$. By Proposition 10.2, $\Gamma(N)$ is also a cograph (resp. a chordal graph). Using Proposition 11.2, we conclude that $C$ is abelian. Moreover, $G/C$ is isomorphic to a subgroup of $\mathop{\mathrm{Aut}}(S)$ which is almost simple, therefore $G/N$ is isomorphic to a subgroup of $\mathop{\mathrm{Out}}(S)$. Thus, by the Schreier Conjecture, $G/N$ is solvable, and therefore statement (b) holds. $\qed$
Remark 8
If $\Gamma(G)$ is a cograph, then in statement $(b)$ of Theorem 11.7 the group $G/N$ is cyclic. This result follows from the classification of finite simple groups whose commuting graphs are cographs, see Theorem 12.15 below.
The following problem is of interest.
Problem 1
Classify finite simple groups whose commuting graphs are chordal.
12 Simple groups whose commuting graph is a cograph or a chordal graph
In this section, we investigate when the commuting graph of a simple sporadic or alternating group is a cograph or a chordal graph, and then obtain a complete classification of finite simple groups whose commuting graph is a cograph.
12.1 CA-groups and AC-groups
A group $G$ is a CA-group if the centralizer of every non-identity element of $G$ is abelian. Another name is CT-groups, since they are the groups for which the commutation relation on non-identity elements is transitive. (If $x\sim y\sim z$ in the CA-group $G$, then $x,z\in C_{G}(y)$, so $x$ and $z$ commute. The converse is similar.)
If $G$ is a CA-group, then the induced subgraph $\Gamma(G)$ on non-identity elements is a disjoint union of complete graphs. As a result, we see:
Proposition 12.1
Let $G$ be a CA-group. Then $\Gamma(G)$ is a cograph, chordal, and perfect.
CA-groups played an important historical role in developing techniques for attacking problems on finite groups; they are now well-understood. From results of Weisner [75], Brauer, Suzuki and Wall [13] and Suzuki [71], we obtain:
Proposition 12.2
A finite CA-group $G$ satisfies one of the following:
  1. (a)
    $G$ is abelian;
  2. (b)
    $G$ is a Frobenius group with abelian Frobenius kernel and cyclic Frobenius complement; or
  3. (c)
    $G\cong\mathop{\mathrm{PSL}}\nolimits(2,q)$ for $q$ a power of $2$.
More generally, $G$ is said to be an AC-group if $C_{G}(x)$ is abelian for every element $x\in G\setminus Z(G)$. This is a weaker condition, though clearly a group with trivial centre is an AC-group if and only if it is a CA-group.
AC-groups investigated by Schmidt [63] and Rocke [62]. (Schmidt investigated groups in which two maximal centralizers intersect just in $Z(G)$; it is not hard to see that, for finite groups, this is equivalent to the AC property.) Abdollahi et al. [2] investigated the connection of this property with the non-commuting graph, the complement of the commuting graph. The classification of the AC-groups is not as complete as that of CA-groups, but the cited authors have proved a number of results about these groups.
From our point of view, AC-groups are as good as CA-groups:
Proposition 12.3
Let $G$ be an AC-group. Then $\Gamma(G)$ is a cograph and a chordal graph.
Proof 21
By Proposition 10.2, it suffices to consider the induced subgraph on the set $G\setminus Z(G)$; and this graph is a disjoint union of complete graphs. $\qed$
Here is a simple general result which we will use.
Proposition 12.4
Let $G$ be a finite group with the property that every non-identity element of $G/Z(G)$ has cyclic centralizer. Then $G$ is an AC-group.
Proof 22
Take $a\in G\setminus Z(G)$. Then $Z(G)\leq C_{G}(a)$, and $C_{G}(a)/Z(G)\leq C_{G}(aZ(G))$ (in $G/Z(G)$); so $C_{G}(a)/Z(G)$ is cyclic, whence $C_{G}(a)$ is abelian. $\qed$
12.2 Generalized dihedral groups and generalized quaternion groups
In this subsection we show that these groups are AC-groups, so that their commuting graphs are cographs as well as chordal graphs.
Proposition 12.5
Let $D(A)$ be the generalized dihedral group as presented in Equation (3). Then $D(A)$ is an AC-group. Hence $\Gamma(D(A))$ is a cograph as well as a chordal graph. In particular, the commuting graph of a dihedral group is a cograph as well as a chordal graph.
Proof 23
The centre of $D(A)$ is the subgroup $T$ of $A$ consisting of elements of orders $1$ or $2$. If we remove it, then $A\setminus T$ is a clique, and the elements outside $A$ fall into cliques of the form $\langle t,T\rangle\setminus T$ for $t\notin A$; there are no edges between these cliques. So the induced subgraph is a disjoint union of complete graphs. Now Proposition 10.2 gives the result.$\qed$
Proposition 12.6
Let $Q_{4m}$ be the generalized quaternion group as presented in Equationi (4). Then $Q_{4m}$ is an AC-group. Hence $\Gamma(Q_{4m})$ is a cograph as well as a chordal graph.
Proof 24
The proof is similar to the preceding theorem. We find that, with $G=Q_{4m}$, we have $Z(G)=\{e,x^{m}\}$; for $11\leq i\leq 2m-$ 1 and $i\neq m$, we find that $C_{G}(x^{i})=\langle x\rangle$, while for $10\leq i\leq 2m-$ 1 we have $C_{G}(x^{i}y)=\langle x^{i}y\rangle=\{e,x^{i}y,x^{m},x^{m+i}y\}$. So again the induced subgraph on $G\setminus Z(G)$ consists of a complete graph of size $22m-$ 2 together with $m$ disjoint edges. $\qed$
12.3 Symmetric groups and alternating groups
The symmetric group, denoted by $\mathop{\mathrm{Sym}}\nolimits(n)$ or $S_{n}$, is the group of permutations on $n$ letters. As we known, the symmetric group is important in many different areas of mathematics, including combinatorics and group theory. Note that $S_{n}$ is abelian if and only if $n\leq 2$.
Proposition 12.7
$\Gamma(S_{n})$ is a cograph if and only if $n\leq 3$.
Proof 25
Proposition cog-subg shows that it is enough to show that $\Gamma(S_{3})$ is a cograph but $\Gamma(S_{4})$ is not (since $S_{n-1}\leq S_{n}$ for any $n$). By Proposition 12.5, we have that $\Gamma(S_{3})$ is a cograph as $S_{3}\cong D_{6}$. i As $S_{n-1}\leq S_{n}$, Proposition 10.2 shows that if $n\leq 3$, then $\Gamma(S_{n})$ is a cograph. In $\Gamma(S_{4})$,
$(12)\sim(12)(34)\sim(13)(24)\sim(13)$
is an induced path isomorphic to $P_{4}$$\qed$
Proposition 12.8
$\Gamma(S_{n})$ is a chordal graph if and only if $n\leq 4$.
Proof 26
It is easy to see that
$(12)\sim(34)\sim(15)\sim(24)\sim(35)\sim(12)$
is an induced cycle isomorphic to $C_{5}$ in $\Gamma(S_{5})$. As in Proposition 12.7, it suffices to prove that $\Gamma(S_{4})$ is a chordal graph. Note that in $\Gamma(S_{4})$, for any non-central element, the cycle type of this element must be one of
$\{(ij),(ij)(kl),(ijk),(ijkl)\},$ (9)
where $\{i,j,k,l\}=\{1,2,3,4\}$. Moreover, it is easy to see that
$C_{S_{4}}((ij))=\{e,(ij),(kl),(ij)(kl)\},C_{S_{4}}((ij))=\{e,(ij),(kl),(ij)(kl)\}$
$C_{S_{4}}((ijkl))=\langle(ijkl)\rangle$ (10)
and
$C_{S_{4}}((ij)(kl))=\{e,(ij),(kl),(ij)(kl),(ik)(jl),(il)(jk),(ikjl),(iljk)\}.$
Suppose, for the sake of contradiction, that $\Gamma(S_{4})$ contains an induced subgraph isomorphic to $C_{n}$ with $n\geq 4$, say $\Delta$. Let $a,b,c,d$ be $4$ distinct vertices of $\Delta$. Then $C_{S_{4}}(x)$ is non-abelian for any $x\in\{a,b,c,d\}$. Note that every centralizer in 10 is abelian. It follows from 9 that every element in $\{a,b,c,d\}$ has cycle type $(ij)(kl)$, which is impossible because $S_{4}$ has precisely three elements that are the product of two disjoint cycles. $\qed$
The group consisting of all even permutations on a set of length $n$ is called the alternating group and is denoted by $\mathop{\mathrm{Alt}}\nolimits(n)$ or $A_{n}$. Note that $A_{n}$ is a simple group for each $n\geq 5$.
Proposition 12.9
The following are equivalent:
$(a)$ $\Gamma(A_{n})$ is a cograph;
$(b)$ $\Gamma(A_{n})$ is a chordal graph;
$(c)$ $n\leq 5$.
Proof 27
The group $A_{5}$ is a CA-group (it is easily checked that all its non-abelian subgroups have trivial centre), so the result for $n=5$ follows from Proposition 12.2. Moreover, $A_{n}\leq A_{5}$ for $n\leq 5$, so the result follows for these groups as well.
On the other hand, in $\Gamma(A_{6})$, it is easy to verify that
$(12)(56)\sim(12)(34)\sim(13)(24)\sim(13)(56)$
is an induced path isomorphic to $P_{4}$. Also, it is not hard to verify that
$(13)(45)\sim(14)(35)\sim(14)(26)\sim(12)(46)\sim(12)(35)\sim(13)(25)\sim(13)(46)$
$\sim(14)(36)\sim(14)(25)\sim(12)(45)\sim(12)(36)\sim(13)(26)\sim(13)(45)$
is an induced cycle isomorphic to $C_{12}$, so $\Gamma(A_{6})$ is not chordal. Now $A_{6}\leq A_{n}$ for $n>6$, so $\Gamma(A_{n})$ is neither a cograph nor chordal for $n\geq 6$, by Proposition 10.2. $\qed$
Remark 9
The graph $\Gamma(A_{6})$ has no induced cycles of length $n$, where $123<n<$ 12. However, $\Gamma(A_{7})$ has an induced cycle isomorphic to $C_{6}$, such as $(123)\sim(456)\sim(127)\sim(345)\sim(126)\sim(457)\sim(123)$. Also, $\Gamma(A_{8})$ has an induced cycle isomorphic to $C_{4}$, such as $(12)(34)\sim(567)\sim(348)\sim(12)(56)\sim(12)(34)$.
12.4 Sporadic groups
In this subsection, we resolve the question for the sporadic simple groups.
Proposition 12.10
If $G$ is a sporadic simple group, then $\Gamma(G)$ is neither a cograph nor a chordal graph.
Proof 28
There are $26$ sporadic simple groups. Observe that $S_{5}$ is a subgroup of the Mathieu group $M_{11}$ by using information in the $\mathbb{ATLAS}$ of finite groups [28]. Also, note that the fact that the Mathieu group $M_{11}$ is a subgroup of all the other sporadic simple groups except $M_{22}$, $J_{1}$, $J_{2}$, $J_{3}$, $He$, $Ru$ and $Th$. Thus, by Proposition 10.2 and Propositions 12.7 and 12.8, the commuting graph of a sporadic simple group containing $M_{11}$ is neither a cograph nor a chordal graph.
Now by [28] again, it is easy to observe that $A_{7}\leq M_{22}$, $D_{6}\times D_{10}\leq J_{1}$, $A_{4}\times A_{5}\leq J_{2}$, $A_{6}\leq J_{3}$, $S_{6}\leq He$, $A_{8}\leq Ru$, $S_{5}\leq Th$. Thus, it follows from Proposition 10.2, Lemma 11.1 and Proposition 12.9 that the commuting graphs of these $7$ sporadic simple groups are neither cographs nor chordal graphs, as desired. $\qed$
12.5 Simple groups of Lie type whose commuting graph is a cograph
Now we turn to simple groups of Lie type. In this analysis, we only consider cographs, not chordal graphs.
Proposition 12.11
Let $G=\mathop{\mathrm{PSL}}\nolimits(2,q)$, with $q\geq 4$. Then $\Gamma(G)$ is a cograph if and only if either $q$ is a power of $2$ or $q=5$.
Note that $\mathop{\mathrm{PSL}}\nolimits(2,5)\cong\mathop{\mathrm{PSL}}\nolimits(2,4)$.
Proof 29
The subgroup structure of $\mathop{\mathrm{PSL}}\nolimits(2,q)$ was determined by Dickson [31], and can be found, for example, in [14].
Suppose first that $q\geq 4$ is a power of $2$. Then every non-trivial element of $\mathop{\mathrm{PSL}}\nolimits(2,q)$ has order $2$ or a prime divisor of $q-1$ and $q+1$. Also, the centralizers of involutions of $\mathop{\mathrm{PSL}}\nolimits(2,q)$ are elementary abelian groups with order $q$, and the centralizers of other non-trivial elements of $\mathop{\mathrm{PSL}}\nolimits(2,q)$ are cyclic of order $q-1$ or $q+1$. It follows that in this case $\mathop{\mathrm{PSL}}\nolimits(2,q)$ is a CA-group, and so by Proposition 12.2, $\Gamma(\mathop{\mathrm{PSL}}\nolimits(2,q))$ is a cograph.
Suppose next that $q$ is odd at least $5$. Let $\epsilon\in\{+1,-1\}$ be chosen so that $q+\epsilon 1$ is divisible by $4$. Note that $\mathrm{PSL}(2,q)$ contains a Klein group such that the centralizers of its involutions are Dihedral groups $D_{2(q+\epsilon 1)}$ which contain disjoint cyclic groups of order $(q+\epsilon)/2$. If $q>5$, then $(q+\varepsilon)/2>2$. So two involutions together with generators of cyclic groups of order $(q+\epsilon)/2$ in their centralizers induce a path of length $4$. (The list of subgroups shows that these generators do not commute.) So $\Gamma(\mathop{\mathrm{PSL}}\nolimits(2,q))$ is not a cograph. $\qed$
Remark 10
It is well-known that $\mathop{\mathrm{PSL}}\nolimits(2,2)\cong S_{3}$ and $\mathop{\mathrm{PSL}}\nolimits(2,3)\cong A_{4}$, therefore their commuting graphs are cographs by Propositions 12.7 and 12.9, respectively.
Proposition 12.12
Let $G$ be either $\mathop{\mathrm{PSL}}\nolimits(3,q)$ or a group of Lie type of BN-rank at least $3$. Then $\Gamma(G)$ is not a cograph.
Proof 30
We begin by showing that, if $G=\mathop{\mathrm{SL}}\nolimits(3,q)$, then $\Gamma(G)$ is not a cograph.
Suppose first that $q\notin\{2,4\}$. Choose a non-zero element $a\in\mathop{\mathrm{GF}}\nolimits(q)$ such that $a^{3}\neq 1$, and let $b=a^{-2}$, so that $b\neq a$.
Consider the four elements $g_{1},\ldots,g_{4}$ given by
$\left(\begin{array}[]{ccc}1&1&0\\ 0&1&0\\ 0&0&1\end{array}\right),\,\left(\begin{array}[]{ccc}a&0&0\\ 0&a&0\\ 0&0&b\end{array}\right),\,\left(\begin{array}[]{ccc}b&0&0\\ 0&a&0\\ 0&0&a\end{array}\right),\,\left(\begin{array}[]{ccc}1&0&0\\ 0&1&1\\ 0&0&1\end{array}\right).$
It is clear that all of them belong to $G$, and that $g_{1}\sim g_{2}\sim g_{3}\sim g_{4}$ in $\Gamma(G)$. A short calculation gives the commutators of the other three pairs; none of them is trivial. So we have an induced path in $\Gamma(G)$.
But the argument gives more. In fact none of the commutators is a scalar matrix, so if $Z$ is any subgroup of the centre of $G$ (the group of scalar matrices), then we also have an induced path in $\Gamma(G/Z)$, so that $\Gamma(G/Z)$ is not a cograph. In particular, $\Gamma(\mathop{\mathrm{PSL}}\nolimits(3,q))$ is not a cograph.
Now we can cover the remaining two cases. Since $\mathop{\mathrm{PSL}}\nolimits(3,2)\cong\mathop{\mathrm{PSL}}\nolimits(2,7)$, the commuting graph of $\mathop{\mathrm{PSL}}\nolimits(3,2)$ is not a cograph. Also, $\mathop{\mathrm{PSL}}\nolimits(3,2)$ is a subgroup of both $\mathop{\mathrm{SL}}\nolimits(3,4)$ and $\mathop{\mathrm{PSL}}\nolimits(3,4)$, so their commuting graphs are not cographs either.
For the remainder of the proof, we use the Levi decomposition of parabolic subgroups of groups with BN-pairs (see [25, Section 8.5]). The Borel subgroup $H$ of such a group is generated by the root subgroups corresponding to the positive roots with respect to a fundamental system $\Pi$ of roots. Parabolic subgroups $P_{\Phi}$ are those which contain $H$; they are generated by $H$ together with the root subgroups corresponding to the negatives of a subset $\Phi$ of the negatives of the fundamental group. According to the Levi decomposition, $P_{\Phi}$ is the semidirect product of its unipotent radical by the Levi factor $L_{\Phi}$, a group with BN-pair corresponding to the roots $\Phi$. In the covering group, $L_{\Phi}$ may have a centre, which may or may not be killed when we factor out the centre of $G$. (For example, $\mathop{\mathrm{SL}}\nolimits(n,q)$ has a parabolic subgroup corresponding to the first two roots, generated by all upper unitriangular matrices together with the lower unitriangular matrices with non-zero off-diagonal elements in positions $1,2,3$ only. These generate $\mathop{\mathrm{SL}}\nolimits(3,q)$; its image in $\mathop{\mathrm{PSL}}\nolimits(n,q)$ is either $\mathop{\mathrm{SL}}\nolimits(3,q)$ or $\mathop{\mathrm{PSL}}\nolimits(3,q)$ depending on the congruence of $n$ and $q$ mod $3$.)
Now the Coxeter–Dynkin diagram of any simple group $G$ of Lie type with BN-pair of rank at least $3$ contains a single edge, corresponding to a root system of type $A_{2}$; so $G$ has a parabolic subgroup whose Levi factor is a central quotient of $\mathop{\mathrm{SL}}\nolimits(3,q)$, as in the example above. From Proposition 10.2, we conclude that $\Gamma(G)$ is not a cograph. $\qed$
Proposition 12.13
Let $G$ be a central quotient of $\mathop{\mathrm{SU}}\nolimits(3,q)$, with $q>2$. Then $\Gamma(G)$ is not a cograph.
Proof 31
Suppose that $q>2$. We take the Hermitian form $H(x_{1},x_{2},x_{3})=x_{1}^{q+1}+x_{2}^{q+1}+x_{3}^{q+1}$. A diagonal matrix with diagonal entries $(a,b,c)$ preserves this form if and only if
$a^{q+1}=b^{q+1}=c^{q+1}=1,$
and belongs to $\mathop{\mathrm{SU}}\nolimits(3,q)$ if additionally $abc=1$. So the set of such matrices forms a subgroup of $\mathop{\mathrm{SU}}\nolimits(3,q)$ isomorphic to $\mathbb{Z}_{q+1}^{2}$, and its quotient in $\mathop{\mathrm{PSU}}\nolimits(3,q)$ is $\mathbb{Z}_{q+1}\times\mathbb{Z}_{(q+1)/\gcd(q+1,3)}$. We observe that $(-1)^{q+1}=1$ (this is clear if $q$ is even and holds if $q$ is odd since then $(-1)^{2}=1$).
Let $a\in GF(q^{2})$ satisfy $a^{q+1}=1$, $a^{3}\neq 1$, and let $b=a^{-2}$, so that $a\neq b$. Now consider the four matrices $g_{1}$, …, $g_{4}$ given by
$\left(\begin{array}[]{ccc}0&1&0\\ 1&0&0\\ 0&0&-1\end{array}\right),\,\left(\begin{array}[]{ccc}a&0&0\\ 0&a&0\\ 0&0&b\end{array}\right),\,\left(\begin{array}[]{ccc}b&0&0\\ 0&a&0\\ 0&0&a\end{array}\right),\,\left(\begin{array}[]{ccc}-1&0&0\\ 0&0&1\\ 0&1&0\end{array}\right).$
It is readily checked that all four matrices belong to $\mathop{\mathrm{SU}}\nolimits(3,q)$, and that $g_{2}$ commutes with $g_{1}$ and $g_{3}$ while $g_{3}$ commutes with $g_{4}$. The other pairs fail to commute; indeed in each case $g_{i}g_{j}$ is not a scalar multiple of $g_{j}g_{i}$. So the four matrices induce a $P_{4}$ subgraph of the commuting graph, which remains as a $P_{4}$ subgraph when the scalar matrices in $\mathop{\mathrm{SU}}\nolimits(3,q)$ are factored out. $\qed$
Among the groups of Lie type, this leaves only those of ranks $1$ or $2$, and we have already handled $\mathop{\mathrm{PSL}}\nolimits(2,q)$, $\mathop{\mathrm{PSL}}\nolimits(3,q)$, and $\mathop{\mathrm{PSU}}\nolimits(3,q)$ for $q>2$. The remaining groups are:
  1. (a)
    $\mathrm{PSp}(4,q)$;
  2. (b)
    $\mathrm{PSU}(n,q)$ for $n\in\{4,5\}$;
  3. (c)
    $G_{2}(q)$, ${}^{2}B_{2}(q)\cong\mathop{\mathrm{Sz}}\nolimits(q)$, ${}^{2}G_{2}(q)$, ${}^{2}F_{4}(q)$, and ${}^{3}D_{4}(q)$.
We have omitted the orthogonal groups here. For these satisfy
  1. (a)
    $\mathop{\mathrm{P}\Omega}\nolimits(3,q)\cong\mathop{\mathrm{PSL}}\nolimits(2,q)$;
  2. (b)
    $\mathop{\mathrm{P}\Omega}\nolimits^{+}(4,q)\cong\mathop{\mathrm{PSL}}\nolimits(2,q)\times\mathop{\mathrm{PSL}}\nolimits(2,q)$ and $\mathop{\mathrm{P}\Omega}\nolimits^{-}(4,q)\cong\mathop{\mathrm{PSL}}\nolimits(2,q^{2})$;
  3. (c)
    $\mathop{\mathrm{P}\Omega}\nolimits(5,q)\cong\mathop{\mathrm{PSp}}\nolimits(4,q)$;
  4. (d)
    $\mathop{\mathrm{P}\Omega}\nolimits^{+}(6,q)\cong\mathop{\mathrm{PSL}}\nolimits(4,q)$;
  5. (e)
    $\mathop{\mathrm{P}\Omega}\nolimits^{-}(6,q)\cong\mathop{\mathrm{PSU}}\nolimits(4,q)$
(see the $\mathbb{ATLAS}$ of Finite Groups [28]). So (a), (b) andn (d) are covered by our earlier results, and (c) and (e) will be treated as symplectic or unitary groups.
Also $G_{2}(q)$ contains a subgroup $\mathop{\mathrm{PSL}}\nolimits(3,q)$, and ${}^{3}D_{4}(q)$ contains $G_{2}(q)$ [29, 48, 49], so their commuting graphs are not cographs, by Proposition 10.2 and Proposition 12.12.
We now finish off the classical groups.
The group $\mathop{\mathrm{PSp}}\nolimits(4,q)$ contains the central product of two copies of $\mathop{\mathrm{SL}}\nolimits(2,q)$. (To see this, note that the $4$-dimensional vector space can be written as the orthogonal direct sum of two $2$-dimensional non-degenerate subspaces; so $\mathop{\mathrm{Sp}}\nolimits(4,q)$ contains $\mathop{\mathrm{SL}}\nolimits(2,q)\times\mathop{\mathrm{SL}}\nolimits(2,q)$. If $q$ is odd, take the quotient by {±$\{\pm I\}$ to obtain a central product.) So its commuting graph is not a cograph, by Lemma 11.4.
For unitary groups, the $\mathop{\mathrm{PSU}}\nolimits(3,2)$ is not simple; so for $q=2$ it remains to deal with $\mathop{\mathrm{PSU}}\nolimits(4,2)$ and $\mathop{\mathrm{PSU}}\nolimits(5,2)$. The first is isomorphic to $\mathop{\mathrm{PSp}}\nolimits(4,3)$ [28], which we have dealt with. The second contains $\mathop{\mathrm{PSL}}\nolimits(2,11)$, see [28]. For $q>2$, by Proposition 12.13 central quotients of $\mathop{\mathrm{SU}}\nolimits(3,q)$ have commuting graphs which are not cographs. For $n=5$, $\mathop{\mathrm{PSU}}\nolimits(n,q)$ contains such a central quotient, obtained as above by writing the $5$-dimensional unitary space as the sum of non-degenerate subspaces of dimensions $2$ and $3$. For $n=4$, the space is the sum of two non-degenerate $2$-dimensional subspaces, so $\mathop{\mathrm{PSU}}\nolimits(4,q)$ contains a central product of two copies of $\mathop{\mathrm{SU}}\nolimits(2,q)$. So the commuting graphs of these groups are not cographs either.
This concludes dealing with the classical groups.
We show next that the commuting graphs of the Suzuki groups are cographs.
Proposition 12.14
For each $q$, the commuting graph of $\mathop{\mathrm{Sz}}\nolimits(q)$ is a cograph.
Proof 32
It is well-known that $\mathop{\mathrm{Sz}}\nolimits(2)\cong 5:4$ is a Frobenius groups with cyclic kernel and cyclic complement. Therefore, by Proposition 11.5, $\Gamma(\mathop{\mathrm{Sz}}\nolimits(2))$ is a cograph. Thus, we can assume that $q>2$.
It suffices to show that the graph obtained by removing the identity is a cograph. Now $\mathop{\mathrm{Sz}}\nolimits(q)$ has maximal cyclic subgroups of orders $q-1$ and $q\pm\sqrt{2q}+1$, and every element of odd order lies in one of these; so the non-identity elements of odd order carry a disjoint union of complete graphs, with no edges to the remaining elements. (The information used in this paragraph and the next is taken from Suzuki’s original paper [72]; it can also be found in [77, Section 4.2].) The remaining elements have order $2$ or $4$, and since $\mathop{\mathrm{Sz}}\nolimits(q)$ is a CIT-group (one in which the centralizer of any involution is a $2$-group), the remaining vertices lie in a disjoint union of copies of the commuting graph on $P\setminus\{e\}$, where $P$ is a Sylow $2$-subgroup. Now $|P|=q^{2}$ and $|Z(P)|=q$; elements of $Z(P)$ are joined to every element of $P$ so we may remove them. The remaining elements of $P$ have order $4$. If $g$ is one of these, then $|C_{P}(g)|=2q$ and $C_{P}(g)$ consists of two cosets of $Z(P)$, so it is abelian; thus the elements of order $4$ fall into $q-1$ sets of size $q$, each set a complete graph with no edges between different sets. Thus the graph is indeed a cograph. $\qed$
Next, the small Ree group ${}^{2}G_{2}(q)$, with $q$ an odd power of $3$, is not simple for $q=3$; for larger $q$, it contains $\mathop{\mathrm{PSL}}\nolimits(2,q)$ (the centralizer of an involution in ${}^{2}G_{2}(q)$ is $\mathbb{Z}_{2}\times\mathop{\mathrm{PSL}}\nolimits(2,q)$ [49]).
The Tits group ${}^{2}F_{4}(2)^{\prime}$ contains a subgroup isomorphic to $\mathop{\mathrm{PSL}}\nolimits(2,25)$ (this can be found in the $\mathbb{ATLAS}$ of Finite Groups [28]); so its commuting graph is not a cograph. This group is a subgroup of the large Ree group ${}^{2}F_{4}(2^{d})$ for all odd $d$ [52].
To summarize our results, combined with the earlier results on alternating and sporadic groups, and noting that $\mathop{\mathrm{Alt}}\nolimits(5)\cong\mathop{\mathrm{PSL}}\nolimits(2,4)$:
Theorem 12.15
Let $G$ be a non-abelian finite simple group. Then the commuting graph of $G$ is a cograph if and only if $G$ is isomorphic to $\mathop{\mathrm{PSL}}\nolimits(2,q)$, with $q$ a power of $2$ and $q>2$, or to $\mathop{\mathrm{Sz}}\nolimits(q)$, with $q$ an odd power of $2$ and $q>2$.
13 Appendix
In this section, we consider small groups whose commuting graphs are cographs or chordal graphs.
Lemma 13.1
If $|G/Z(G)|=pqr$ with primes $p,q,r$ (not necessary distinct), then $G$ is an AC-group.
Proof 33
Let $a\in G\setminus Z(G)$. It suffices to prove that $C_{G}(a)$ is abelian. Note that $Z(G)\subseteq Z(C_{G}(a))\subseteq C_{G}(a)<G$. Since $|G/Z(G)|=pqr$, we have that $|C_{G}(a)/Z(G)|<pqr$ and $|C_{G}(a)/Z(G)|$ is a divisor of $pqr$. By the argument of Proposition 12.4, it suffices to consider the case where $C_{G}(a)/Z(G)$ is non-cyclic and has order $pq$.
Then $Z(G)a$ has order $p$ or $q$. Thus, $C_{G}(a)/Z(G)$ has an element, say $Z(G)b$, such that $C_{G}(a)/Z(G)=\langle Z(G)b,Z(G)a\rangle$. Note that $[b,a]=e$. Hence, $C_{G}(a)=\langle b,a,Z(G)\rangle$ is an abelian group, as desired. $\qed$
A similar but easier proof shows the following.
Corollary 13.2
If $|G/Z(G)|=pq$ or $p$ where $p,q$ are primes, then $G$ is an AC-group. In particular, If $|G|=pq$ or $pqr$ with primes $p,q,r$, then $G$ is either an abelian group or an AC-group.
Note that these primes $p,q,r$ in Lemma 13.1 and Corollary 13.2 are not necessarily distinct. Now combining Corollary 13.2, Lemma 13.1 and Proposition 12.3, we have the following result.
Lemma 13.3
$\Gamma(G)$ is a cograph as well as a chordal graph if one of the following holds:
  1. (a)
    $G$ is abelian;
  2. (b)
    $|G|=pqr$ or $pq$, where $p,q,r$ are primes;
  3. (c)
    $|G/Z(G)|=p$, $pq$ or $pqr$, where $p,q,r$ are primes (not necessary distinct).
Proposition 13.4
If $G$ is a group of minimal order such that $\Gamma(G)$ is not a cograph, then $|G|=24$. In particular, for $|G|\leq 24$, $\Gamma(G)$ is a cograph if and only if $G\ncong S_{4}$.
Proof 34
Proposition 12.7 implies that $\Gamma(S_{4})$ is not a cograph. Thus, it suffices to show that if $|G|\leq 24$, then $\Gamma(G)$ is a cograph if and only if $G\ncong S_{4}$. Now suppose that $|G|\leq 24$. It follows from Lemma 13.3 that if $|G|\notin\{16,24\}$, then $\Gamma(G)$ is a cograph. If $|G|=16$, since a $2$-group has non-trivial center, we have that $|G/Z(G)|=2^{k}$ with $k\leq 3$, and so Lemma 13.3(c) implies that $\Gamma(G)$ is a cograph. Finally, we suppose that $|G|=24$. Note that the fact that if $G\ncong S_{4}$, then $G$ has non-trivial center (cf. Burnside [18]). So, by Lemma 13.3(c), we have that if $G\ncong S_{4}$, then $\Gamma(G)$ is a cograph, as desired. $\qed$
In this paper, we identify a finite group by the unique identifier of this finite group in SmallGroups library [10] in the GAP system [36]. To be specific, the $m$-th group of order $n$ is identified as SmallGroup$[n,m]$, where $[n,m]$ is called the GAP ID of the group.
By Lemma 11.1, $\Gamma(S_{3}\times S_{3})$ is not a cograph. One can use the GAP system [36] to classify all finite groups $G$ satisfying $|G|\leq 36$ and $\Gamma(G)$ is not a cograph; these groups are presented in Table 1. In particular, by Lemma 11.1, $\Gamma(S_{3}\times S_{3})$ is not a cograph.
Table 1: All groups $G$ such that $\Gamma(G)$ is not a cograph and $|G|\leq 36$
GAP ID Structure
$[24,12]$ $S_{4}$
$[32,6]$ $(\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2})\rtimes\mathbb{Z}_{4}$
$[32,7]$ $(\mathbb{Z}_{8}\rtimes\mathbb{Z}_{2})\rtimes\mathbb{Z}_{2}$
$[32,8]$ $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})_{\cdot}(\mathbb{Z}_{4}\times\mathbb{Z}_{2})$
$[32,43]$ $\mathbb{Z}_{8}\rtimes(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$
$[32,44]$ $(\mathbb{Z}_{2}\times Q_{8})\rtimes\mathbb{Z}_{2}$
$[32,49]$ $(\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2})\rtimes(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$
$[32,50]$ $(\mathbb{Z}_{2}\times Q_{8})\rtimes\mathbb{Z}_{2}$
$[36,10]$ $S_{3}\times S_{3}$
For the groups of order $32$, there are precisely two extra-special $2$-groups $D_{8}\circ D_{8}$ and $D_{8}\circ Q_{8}$ as follows:
$D_{8}\circ D_{8}$ $=$ $\langle x,y,z,w|x^{2}=y^{2}=z^{2}=w^{2}=[x,y]^{2}=[z,w]^{2}=e,$ (12)
$[x,z]=[z,y]=[y,w]=[w,x]=e\rangle$
where $[x,y]=[z,w]$ is an involution belonging to $Z(D_{8}\circ D_{8})$;
$D_{8}\circ Q_{8}$ $=$ $\langle x,y,z,w|x^{2}=y^{2}=z^{4}=w^{4}=[x,y]^{2}=[z,w]^{2}=e,$ (14)
$[x,z]=[z,y]=[y,w]=[w,x]=e\rangle$
where $z^{2}=w^{2}=[x,y]=[z,w]$ is an involution belonging to $Z(D_{8}\circ Q_{8})$.
Proposition 13.5
If $G$ is a group of minimal order such that $\Gamma(G)$ is not a chordal graph, then $|G|=32$. In particular, if $G\cong D_{8}\circ D_{8}$ or $D_{8}\circ Q_{8}$, then $\Gamma(G)$ is not chordal.
Proof 35
Suppose that $G\cong D_{8}\circ D_{8}$ or $D_{8}\circ Q_{8}$. Then by (12) and (14), it is easy to see that the induced subgraph of $\Gamma(G)$ by the set $\{x,y,z,w\}$ is isomorphic to $C_{4}$, which implies that $\Gamma(G)$ is not chordal.
Now, it suffices to show that if $|G|<32$, then $\Gamma(G)$ is a chordal graph. Suppose that $|G|\leq 31$. Note that a $p$-group has non-trivial center where $p$ is a prime. From Lemma 13.3, it follows that if $|G|\neq 24$, then $\Gamma(G)$ is a chordal graph. Now consider $|G|=24$. If $G\cong S_{4}$, then by Proposition 12.8, $\Gamma(G)$ is chordal. If $G\ncong S_{4}$, then by Lemma 13.3 and the proof of Proposition 13.4, $\Gamma(G)$ is chordal, as desired. $\qed$
We have taken these computations further and determined, using GAP with the package GRAPE [68], the groups of order up to $128$ whose commuting graphs are not cographs. The numbers of groups are given in Table 2.
Table 2: Small groups whose commuting graph is not a cograph
$\begin{array}[]{|r|r|r|r|r|r|}\hline\cr\hbox{Order}&\hbox{Groups}&\hbox{Order}&\hbox{Groups}&\hbox{Order}&\hbox{Groups}\\ \hline\cr 24&1&32&7&36&1\\ 48&10&54&2&60&2\\ 64&115&72&11&80&12\\ 84&1&96&112&100&2\\ 108&10&112&8&120&15\\ 126&2&128&1539&&\\ \hline\cr\end{array}$
Acknowledgements
The authors of this paper contributed equally to this work. The authors are ordered with respect to alphabet ordering in English.
This work was initiated during the online series of Research Seminars on “Groups and Graphs” in March–August 2021, run by Ambat Vijayakumar and Aparna Lakshmanan, Cochin Univ. of Science and Technology.
The second author acknowledges the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Groups, representations and applications: new perspectives (supported by EPSRC grant no. EP/R014604/1), where he held a Simons Fellowship.
The third author was supported by National Natural Science Foundation of China (Grant Nos. 11801441 and 12326333) and Shaanxi Fundamental Science Research Project for Mathematics and Physics (Grant No. 22JSQ024).
The fourth author is thankful to Prof. Andrey V. Vasil’ev for his helpful comments which improved this text. Also the fourth author started thinking on this project during her visit to the Steklov Institute of Mathematics, and she is very thankful to colleagues from this institution for hospitality and moral support.