### v4

# Aspects of the commuting graph [^1] To the memory of Richard Parker

Abstract
The commuting graph of a group $G$ is the simple graph whose vertices are the elements of $G$, two distinct vertices joined if they commute.
Our purpose in this paper is twofold.
  1. (a)
    We discuss the computational problem of deciding whether a given graph is the commuting graph of a finite group; we give a quasipolynomial algorithm, and a polynomial algorithm for the case when the group is an extraspecial $p$-group for $p$ an odd prime.
  2. (b)
    We give new results on the question of whether the commuting graph of a given group is a cograph or a chordal graph, two classes of graphs defined by forbidden subgraphs.
The problems are not unrelated, since there are a number of cases where hard computational problems on graphs are easier when restricted to special classes of graphs; we conjecture that the recognition problem is polynomial for cographs and chordal graphs.
keywords:
finite group , commuting graph , forbidden subgraph , cograph , chordal graph , graph isomorphism , quasipolynomial algorithm , extraspecial group
MSC:
[2020] 05C25 , 05C85 , 20D60
1]arvind@imsc.res.in 2]pjc20@st-andrews.ac.uk 3]xuanlma@xsyu.edu.cn 4]butterson@mail.ru
1 Introduction
Graphs associated with some algebraic structures, such as the power graph of a group or semigroup, have been actively investigated (cf. [1, 20, 22, 24]) Of course the famous Cayley graphs have a long history. Moreover, graphs from algebraic structures have valuable applications (cf. [47]): for example, Cayley graphs as classifiers for data mining [48].
Let $G$ be a group and $S$ be a subset of $G$. The commuting graph of $G$ on $S$ is a simple graph with vertex set $S$, and two distinct elements are adjacent in this graph if they commute in $G$. In particular, if $S=G$, then we denote this commuting graph on $G$ by $\Gamma(G)$. In 1955, Brauer and Fowler [12] first introduced the commuting graph of a group and showed that there are only finitely many simple groups with even order such that they have a prescribed centralizer. In fact, in the seminal paper [12], they considered the induced subgraph of the commuting graph on the set of all non-trivial elements of the group. Since then, the commuting graphs have received considerable attention and have been well studied in the literature. For example, Segev et al. [66, 67, 68] established some longstanding conjectures on division algebras by using some combinatorial parameters of commuting graphs. Britnell and Gill [15] classified all finite quasi-simple groups whose commuting graph is perfect, where the vertex set of the commuting graph is the set consisting of all non-central elements. Giudici and Parker [39] showed that for every positive integer $d$, there is a finite $2$-group $G$ such that the commuting graph of $G$ on the set of all non-central elements has diameter greater than $d$; further results are found in [30]. Morgan and Parker [60] proved that, for a finite group $G$ with trivial centre, every connected component of the commuting graph of $G$ on the set of all non-trivial elements has diameter at most $10$. Recently, Lewis et al. [9] extended some results by Morgan and Parker [60] about commuting graphs. Moreover, some further applications of commuting graphs in group theory can be found in [21].
The first part of this paper is devoted to the following algorithmic question.
Question 1
Given a graph $\Gamma$, is it isomorphic to the commuting graph of a group?
We give a quasi-polynomial algorithm in general, and a polynomial-time algorithm in the case of extraspecial $p$-groups with $p$ odd.
A number of important graph classes, including perfect graphs, cographs, chordal graphs, split graphs, and threshold graphs, can be defined either structurally or in terms of forbidden induced subgraphs. Forbidden subgraphs of power graphs were studied in [56]. In [21], the second author described various graphs that can be associated with a given group $G$. The vertex set of such a graph is the set of all elements of $G$, and its edge set reflects the group structure of $G$ in some way. He proposed the following question.
Question 2
([21, Question 14]) For which finite groups is the commuting graph a perfect graph, or a cograph, or a chordal graph, or a split graph, or a threshold graph?
Britnell and Gill [15] studied the finite groups whose commuting graph is perfect. The second and third authors [53] completely classified the finite groups whose commuting graph is a split graph or a threshold graph.
In Part II of the paper (Sections 10, 11, 12, and 13) , we address the question of when the commuting graph is a cograph or a chordal graph. We prove that in some sense, the class of such groups generalizes the class of finite solvable groups (see Theorem 11.7 in Subsection 11.3). Also we determine all finite simple groups whose commuting graph is a cograph (see Problem 1 in Subsection 11.3 and Theorem 12.15 at the end of Section 12). The results partly answer Question 2.
These issues are relevant to the recognition question (Question 1). There are extra hypotheses which could make the problem of recognition of commuting graphs easier and yield faster algorithms. One is to restrict the groups to a known class, such as nilpotent groups. Another is to restrict the graphs to a class in which computation is easier. For example, the result of the second and third authors [53] describes precisely the split or threshold graphs which are commuting graphs; these are the complete graphs, together with the graphs consisting of an $n$-vertex complete graph (where $n$ is odd), with $n$ pendant edges added at one of its vertices. Clearly these are easily recognised. We have not pursued this issue further.
We note that some works, especially on connectedness and diameter, define the commuting graph to have vertex set $G\setminus Z(G)$ (that is, omitting dominating vertices). This has little effect on our results. For the recognition problem, if we are given the commuting graph on $G\setminus Z(G)$, then the commuting graph on $G$ is found by adding $|Z(G)|$ dominating vertices to the graph, and this number is a divisor of $|G\setminus Z(G)|$ so there are not many possibilities. In the second part of the paper too, as mentioned before Proposition 10.2, we can define the commuting graph without including the centre $Z(G)$ and the results essentially remain the same.
We conclude the introduction with a suggestion for further research, which would link the two parts of the paper together more closely.
Question 3
Suppose that $\Gamma$ is a perfect graph, a cograph or a chordal graph. Is there a polynomial algorithm to decide whether $\Gamma$ is the commuting graph of a group?
2 Preliminaries
In this section, we will introduce some notation and recall some definitions on groups and graphs.
Every group considered is finite. Throughout, $G$ always denotes a finite group and $1_{G}$ denotes the identity element of $G$, and we denote $1_{G}$ simply by $1$ if the considered group is unambiguous. The order of the group $G$ is its cardinality, denoted $|G|$. For an element $g\in G$, the order $o(g)$ of $g$ in $G$ is the size of the cyclic subgroup $\langle g\rangle$ generated by $g$. The element $g$ is called an involution if $o(g)=2$. The centre of $G$, denoted by $Z(G)$, is defined as the subgroup
$\{x\in G:gx=xg\hbox{ for each $g\in G$}\}.$
Also, the centralizer $C_{G}(g)$ of $g$ in $G$ is the subgroup of all elements that commute with $g$, that is,
$C_{G}(g)=\{x\in G:gx=xg\}.$
The commutator of two elements $x,y$ is denoted as $[x,y]$, that is, $[x,y]=x^{-1}y^{-1}xy$. Note that $x$ and $y$ commute if and only if $[x,y]=1$. The derived group, or commutator subgroup, $G^{\prime}$ of $G$ is the subgroup generated by all commutators. The derived series is the series of subgroups $G^{(n)}$ of $G$ defined by $G^{0}=G$ and $G^{(n+1)}=(G^{(n)})^{\prime}$. The group $G$ is solvable of $G^{(n)}=1$ for some $n$.
Denote by $\mathbb{Z}_{n}$ the cyclic group with order $n$. For a prime $p$, a $p$-group is a group whose order is a power of $p$. An elementary abelian $p$-group is a $p$-group that is a direct product of cyclic groups of order $p$. A Sylow $p$-subgroup of $G$ is a $p$-subgroup whose index in $G$ is coprime to $p$.
An automorphism of $G$ is a permutation $\alpha$ of $G$ preserving the group operation: $(gh)^{\alpha}=g^{\alpha}h^{\alpha}$. An inner automorphism is a map $\phi(g):x\mapsto g^{-1}xg$. It is readily checked that the automorphisms of $G$ form a group $\mathop{\mathrm{Aut}}(G)$, while the inner automorphisms form a subgroup $\mathop{\mathrm{Inn}}(G)$ of $\mathop{\mathrm{Aut}}(G)$.
A subgroup $N$ of $G$ is a normal subgroup of $G$, denoted $N\lhd G$, if $gN=Ng$ for all $g\in G$. The quotient $G/N$ is the group whose elements are the cosets of $N$ in $G$ with the (well-defined) group operation $(Ng_{1})(Ng_{2})=N(g_{1}g_{2})$. For example,
  • (a)
    the centre $Z(G)$ is a normal subgroup of $G$, and $G/Z(G)$ is isomorphic to $\mathop{\mathrm{Inn}}(G)$;
  • (b)
    the inner automorphism group $\mathop{\mathrm{Inn}}(G)$ is a normal subgroup of the automorphism group $\mathop{\mathrm{Aut}}(G)$; the quotient is the outer automorphism group $\mathop{\mathrm{Out}}(G)$ of $G$;
  • (c)
    the derived group $G^{\prime}$ is a normal subgroup of $G$; it is the smallest normal subgroup $N$ for which $G/N$ is abelian.
A subgroup of $G$ is characteristic if it is mapped to itself by all automorphisms of $G$. Since a subgroup is normal if it is mapped to itself by all inner automorphisms, we see that each characteristic subgroup is normal.
A group $G$ is simple if its only normal subgroups are $G$ and $\{1\}$. In this paper, we will use the Classification of the Finite Simple Groups (CFSG), which will be explained when it is used in Section 8.1. Here we note one consequence of CFSG, the proof of the Schreier conjecture:
Theorem 2.1
If $G$ is a non-abelian finite simple group, then $\mathop{\mathrm{Out}}(G)$ is solvable.
Each minimal (non-trivial) normal subgroup of a finite group is a direct product of isomorphic simple groups [32, Theorem 4.3A(iii)]. The socle of $G$, denoted $\mathop{\mathrm{Soc}}(G)$, is the product of the minimal normal subgroups of $G$; it is a direct product of simple groups (not necessarily isomorphic). If the socle is simple, we say that $G$ is almost simple.
A normal series for a group $G$ is a collection of normal subgroups $N_{i}\lhd G,0\leq i\leq r$ such that
$\{1\}=N_{0}\subset N_{1}\subset N_{2}\subset\cdots\subset N_{r}=G.$
This series is a central series for $G$ if $N_{i}/N_{i-1}\subseteq Z(G/N_{i-1})$ for $1\leq i\leq r$. A group $G$ is nilpotent if it has a central series. If $G$ is nilpotent, and we build the upper central series by taking $N_{i}/N_{i-1}=Z(G/N_{i-1})$, the number of steps in the series is the nilpotence class of $G$.
Theorem 2.2
[45, Theorem 1.26] If $G$ is a finite group, the following conditions are equivalent:
  1. (a)
    $G$ is nilpotent.
  2. (b)
    Every Sylow subgroup of $G$ is normal in $G$.
  3. (c)
    $G$ is the direct product of its Sylow subgroups.
A central product $H\circ K$ of two groups $H$ and $K$ with respect to an isomorphism $\theta$ from a central subgroup $Z$ of $H$ to a central subgroup of $K$ is defined to be
$(H\times K)/\{(z,z\theta):z\in Z\}.$
By $F(G)$ we will denote the Fitting subgroup of $G$, i.e. the largest normal nilpotent subgroup of $G$.
A component of $G$ is a subgroup $H$ of $G$ which is quasisimple (that is, $H=H^{\prime}$ and $H/Z(H)$ is simple) and subnormal in $G$ (that is, a term in some normal series of $G$). Let $\mathop{\mathrm{Comp}}(G)$ be the set of components of $G$. Then the subgroup $E(G)$ is the subgroup generated by $\mathop{\mathrm{Comp}}(G)$, and the generalized Fitting subgroup $F^{*}(G)$ of $G$ is the central product
$F^{*}(G)=F(G)\circ E(G),$ (1)
see [3, 31.12].
Given a positive integer $n$, we use $D_{2n}$ to denote the dihedral group with order $2n$. A presentation of $D_{2n}$ is given by
$D_{2n}=\langle a,b:a^{n}=b^{2}=1,bab=a^{-1}\rangle.$ (2)
Note that $D_{2n}$ is an abelian group if and only if $n\in\{1,2\}$, and for all $1\leq i\leq n$,
$o(a^{i}b)=2,~{}~{}D_{2n}=\langle a\rangle\cup\{b,ab,a^{2}b,\ldots,a^{n-1}b\}.$ (3)
Suppose that $A$ is an abelian group. The generalized dihedral group corresponding to $A$, denoted by $D(A)$, is the external semidirect product $A\rtimes T$ of $A$ with the cyclic group $T=\langle t\rangle$ of order two, with the non-identity element $t$ acting as the inverse map on $A$, that is,
$x^{t}=x^{-1}\textrm{ for all }x\in A.$
Viewing this as an internal semidirect product, $A$ is an abelian normal subgroup of $D(A)$ of index two. A presentation for $D(A)$ is as follows:
$D(A)=\langle A,t:o(t)=2,tgt=g^{-1}\hbox{ for each $g\in A$}\rangle.$ (4)
Observation 1
Let $D(A)$ be the generalized dihedral group as presented in (4). Then we have the following.
  • (i)
    Every element of $D(A)\setminus A$ is an involution;
  • (ii)
    $D(A)$ is abelian if and only if $A$ is an elementary abelian $2$-group, if and only if $D(A)$ is an elementary abelian $2$-group;
  • (iii)
    If $A$ is cyclic of order $n$, then $D(A)\cong D(\mathbb{Z}_{n})\cong D_{2n}$.
Proof 1
As $A$ is a normal subgroup of $D(A)$ of index two, every element $g\in D(A)\setminus A$ is in the coset $At$ and is of the form $g=at$. Hence, $g^{2}=atat=aa^{t}=aa^{-1}=1$ which means $g$ is an involution. Furthermore, $D(A)$ is commutative if and only if for all $a\in A$, $at=ta$. That is, $tat=a^{-1}=a$ which implies $a^{2}=1$ for all $a\in A$, implying that $A$ is an elementary abelian $2$-group which, in turn, implies that $D(A)$ is an elementary abelian $2$-group. The last part follows directly from the definition of dihedral groups. $\qed$
Observation 1(ii) means that $D(A)$ is non-abelian if and only if $A$ has an element of order at least $3$. Moreover, Observation 1(iii) implies that the dihedral groups are special cases of generalized dihedral groups. As an example, one can check that $D(\mathbb{Z}_{2}\times\mathbb{Z}_{6})\cong\mathbb{Z}_{2}\times D_{12}$.
Given a positive integer $m$ at least $2$, Johnson [46, pp. 44–45] introduced the generalized quaternion group $Q_{4m}$ of order $4m$, that is,
$Q_{4m}=\langle x,y:x^{m}=y^{2},x^{2m}=1,y^{-1}xy=x^{-1}\rangle.$ (5)
Observation 2
For any $1\leq i\leq 2m$, $o(x^{i}y)=4$ and $Z(Q_{4m})=\{1,x^{m}\}$.
Proof 2
By definition, $x^{2m}=y^{4}=1$, which implies that $o(y)=4$, as there are no other relations. Furthermore, as $xy=yx^{-1}$ we can write $x^{i}y=yx^{-i}$ for any positive integer $i$. It follows that $x^{i}yx^{i}y=yx^{-i}x^{i}y=y^{2}$ for all $i$. Hence, $(x^{i}y)^{4}=y^{4}=1$ for all $i$, implying that $o(x^{i}y)=4$ for $1\leq i\leq 2m$. We can express every element in $Q_{4m}$ as $x^{i}y^{j}$ for $10\leq i\leq 2m-$ 1 and $0\leq j\leq 3$. Now, the element $x^{i}y^{j}$ is in the centre $Z(Q_{4m})$ if and only if it commutes with both $y$ and $x$. Clearly, $yx^{i}y^{j}=x^{i}y^{j}y$ if and only if $yx^{i}=x^{i}y$ if and only if $yx^{i}=yx^{-i}$, which is equivalent to $x^{2i}=1$. This forces $i\in\{0,m\}$. Thus, $x^{m}$ is in the centre $Z(Q_{4m})$. If $j\in\{0,2\}$ then $x^{i}y^{j}\in\{1,x^{m}\}$ which is contained in the centre $Z(Q_{4m})$. Consider the case $j\in\{1,3\}$. Then $x^{m}y^{j}x=x^{m+1}y^{j}$ if and only if $y^{j}x=xy^{j}$, which is equivalent $y^{j}x=y^{j}x^{-1}$ as $xy^{j}=y^{j}x^{-1}$ since $j$ is odd. But that implies $x^{2}=1$ which is a contradiction as $o(x)=2m>2$. Hence $Z(Q_{4m})=\{1,x^{m}\}$. $\qed$
The group of all permutations of $\{1,2,\ldots,n\}$ is the symmetric group $S_{n}$. For $n\geq 2$, its derived group is the alternating group $A_{n}$ consisting of all even permutations. The alternating group is simple for $n\geq 5$.
A group $G$ is called a Frobenius group if there is a proper subgroup $H$ of $G$ such that $H\cap H^{g}=1$ whenever $g\in G\setminus H$. Let
$K=\{1_{G}\}\cup(G\setminus(\cup_{g\in G}H^{g}))$
be the Frobenius kernel of $G$. It is well-known (see, for example, [3, 35.24 and 35.25]) that $K\trianglelefteq G$, $G=K\rtimes H$, $C_{G}(h)\leq H$ for each $h\in H$, and $C_{G}(k)\leq K$ for each $k\in K$. Moreover, by the Thompson theorem on finite groups with fixed-point-free automorphisms of prime order [76, Theorem 1], $K$ is nilpotent. Frobenius groups are also characterised as transitive non-regular permutation groups in which non-identity elements fix at most one point.
The Frattini subgroup $\Phi(G)$ of a finite group $G$ is the intersection of all maximal proper subgroups of $G$. A $p$-group $G$ is extraspecial group $G$ if its centre $Z(G)$ is of order $p$ and coincides with its derived subgroup $G^{\prime}$ and the Frattini subgroup $\Phi(G)$ [45, Chapter 4, pp. 123]. Furthermore, there are exactly two non-isomorphic extraspecial groups of order $p^{2r+1}$, for each $r\geq 1$ and each prime $p$.
Every graph considered is simple: that is, it is undirected, with no loops and no multiple edges. Given a graph $\Gamma$, the vertex set and edge set of $\Gamma$ are denoted by $V(\Gamma)$ and $E(\Gamma)$, respectively. For a vertex subset $U\subset V(\Gamma)$ the induced subgraph $\Gamma[U]$ has vertex set $U$ and edge set $\{\{u,v\}\in E(\Gamma)\mid u,v\in U\}$. In other words, the subgraph induced by $U$ has vertex set $U$ and its edge set consists of all edges of $\Gamma$ between vertices in $U$.
Let $\mathcal{F}$ be a collection of finite graphs. If no induced subgraph of $\Gamma$ is isomorphic to a graph in $\mathcal{F}$, then $\Gamma$ is called $\mathcal{F}$-free; and if =$\mathcal{F}=\{\Delta\}$, we say that $\Gamma$ is $\Delta$-free. Sometimes we say instead that $\Gamma$ forbids $\mathcal{F}$ or $\Delta$. Denote by $P_{n}$ and $C_{n}$ the path of length $n-1$ and the cycle of length $n$, respectively. For distinct $x,y\in V(\Gamma)$, if $x$ is adjacent to $y$ in $\Gamma$, then we write $x\sim_{\Gamma}y$ to denote this edge $\{x,y\}$, or just by $x\sim y$ when $\Gamma$ is clear from the context. In particular, if $P_{n}$ is the subgraph of $\Gamma$ induced by the vertex subset $\{x_{1},x_{2},\ldots,x_{n}\}$ with adjacencies $x_{i}\sim x_{i+1},1\leq i\leq n-1$ we denote the induced path $P_{n}$ by
$x_{1}\sim x_{2}\sim\cdots\sim x_{n}.$
Similarly, in $\Gamma$, if the $n$-cycle $C_{n}$ is the subgraph induced by the subset of vertices $\{x_{1},x_{2},\ldots,x_{n}\}$, where $\{x_{1},x_{n}\}\in E(\Gamma)$ and $\{x_{i},x_{i+1}\}\in E(\Gamma)$ for all $11\leq i\leq n-$ 1, we denote it by
$x_{1}\sim x_{2}\sim\cdots\sim x_{n}\sim x_{1}.$
The neighbourhood of a vertex $v$ is $N(v)=\{w\in V(\Gamma):w\sim v\}$, and the closed neighbourhood of $v$ is $\bar{N}(v)=\{v\}\cup N(v)$. A vertex $v$ of a graph $\Gamma$ is said to be dominating if $v\sim w$ for all $w\in V(\Gamma)\setminus\{v\}$.
Part I Recognising the commuting graph
3 The Recognition Problem
The main focus of this part is the recognition problem of commuting graphs. It is basically an algorithmic question. Given an undirected graph $X=(V,E)$ as input, the aim is to design an algorithm that efficiently checks if there is a group $G$ with $|V|$ elements such that the input graph $X$ is isomorphic to $\Gamma(G)$, and if so to efficiently determine such an isomorphism.
We summarize our results.
  • (a)
    We obtain a deterministic polynomial-time algorithm for the case of extraspecial $p$-groups for odd primes $p$. Extraspecial $p$-groups are a special case of $p$-groups of nilpotence class $2$. A formal definition is in Section 7.
  • (b)
    For the general case, we obtain a quasipolynomial-time algorithm. The algorithm is essentially based on a simple $2^{O(\log^{3}n)}$ time algorithm that enumerates the multiplication tables of all distinct groups of order $n$ (where by distinct we mean mutually non-isomorphic groups of order $n$). This enumeration algorithm combined with Babai’s quasipolynomial time algorithm for graph isomorphism [4] yields a simple $2^{O(\log^{3}n)}$ time algorithm to check if a given $n$-vertex graph $X$ is the commuting graph of a group of order $n$: for each $n$-element group $G$ that is output by the enumeration algorithm we check if $\Gamma(G)$ is isomorphic to $X$ by using Babai’s isomorphism test [4].
  • (c)
    A natural question in connection with our algorithm for recognizing the commuting graphs of extraspecial groups is whether groups with the same commuting graphs are isoclinic. This holds for extraspecial groups which is exploited by the algorithm, and it is natural to conjecture that this property holds for all groups of nilpotence class 2 (of which extraspecial groups are a subclass). We present counter-examples for class-3 nilpotent groups and conjecture that the property holds for class-2 nilpotent groups.
  • (d)
    Additionally, we also show an efficient reduction of the commuting graph recognition problem to the recognition problem of commuting graphs of indecomposable groups. We also note recognizing the commuting graph of generalized dihedral groups along with an extension to Frobenius groups.
Some related work.  There is a result by Giudici and Kuzma [38] that shows the following: every $n$-vertex graph $X$ with at least two vertices of degree $n-1$ is realizable as the commuting graph of a semigroup. It is easy to see that their construction actually gives a polynomial-time algorithm for finding a semigroup with $n$ elements and a bijection from it to the vertex set of $X$ such that the edges of $X$ realize the commuting relation of the semigroup. It is a nearly complete answer to the question in the semigroups setting.
Solomon and Woldar [71] have shown that the commuting graph $\Gamma(G)$ of a finite simple group $G$ uniquely determines the group. More precisely, they have shown that for any finite group $H$ and any finite simple group $G$, their commuting graphs $\Gamma(H)$ and $\Gamma(G)$ are isomorphic if and only if the groups $G$ and $H$ are isomorphic. An interesting question is whether or not the proof of this result [71] yields a polynomial-time algorithm for checking if a given graph $X$ is the commuting graph of a simple group.
4 Basic properties
We being with some preliminary observations that are well-known in the literature (see, e.g., the survey [21]).
Let $G$ be a finite group. What are the cliques of $\Gamma(G)$? If a vertex subset $S$ induces a clique in $\Gamma(G)$ then $S$ is a commuting subset of elements of $G$. Conversely, every commuting subset of elements of $G$ forms a clique in $\Gamma(G)$. Which cliques of $\Gamma(G)$ correspond to subgroups of $G$? Although we cannot directly infer the group multiplication from $\Gamma(G)$, we can observe the following.
Lemma 4.1
Let $G$ be a finite group and $\Gamma(G)$ its commuting graph. A vertex subset $S$ is a maximal clique of $\Gamma(G)$ if and only if $S$ is a maximal abelian subgroup of $G$.
Proof 3
Suppose $H$ is a maximal abelian subgroup of $G$. Clearly, $H$ is a clique in $X$. If $x\notin H$ is adjacent to all of $H$ then $x$ commutes with all of $H$ implying that $\langle H\cup\{x\}\rangle$ is an abelian subgroup of $G$ strictly larger than $H$. Hence the clique induced by $H$ is maximal.
Conversely, suppose $S\subset G$ is a maximal clique in $\Gamma(G)$. Then any two elements $x,y\in S$ commute with each other. Hence, the subgroup $H=\langle S\rangle$ generated by $S$ is an abelian subgroup of $G$ which means that the vertex subset $H$ is a clique in $\Gamma(G)$ containing $S$. As $S$ is maximal, by assumption, it follows that $S=H$ which means $S$ is a maximal abelian subgroup. $\qed$
4.1 The vertex degrees of $\Gamma(G)$ and conjugacy classes of $G$
Let $X=(G,E)$ be the commuting graph of a finite group $G$. Let $x^{G}$ denote the conjugacy class of $x\in G$:
$x^{G}=\{g^{-1}xg\mid g\in G\},$
which is the orbit of $x$ under the conjugation action of $G$ on itself. Let $\deg(v)$ denote the degree of a node $v$ in the graph $X$. The orbit-stabilizer lemma [19] directly implies the following.
Proposition 4.2
For each $x\in G$ its centralizer $C_{G}(x)$ is the closed neighborhood $\bar{N}(x)$ of $x$ in $\Gamma(G)$, and
$|C_{G}(x)|=1+\deg(x)={\frac{|G|}{|x^{G}|}},\textrm{ for all }x\in G.$
Let $n=|G|$, let $m$ denote the number of edges in the commuting graph $\Gamma(G)$, and let $k$ denote the number of conjugacy classes. As $\sum_{x\in G}\deg(x)=2m$, we have:
$2m+n=\sum_{x\in G}{\frac{|G|}{|x^{G}|}}=|G|\cdot\sum_{x\in G}{\frac{1}{|x^{G}|}}=n\cdot k.$ (6)
Thus the number of conjugacy classes of $G$ is $k=(2m+n)/n$, which, by the above equation, can be inferred from the commuting graph. Thus, for example, the only regular commuting graphs are the complete graphs, which are the commuting graphs of abelian groups.
Let $f(n)$ denote the minimum number of conjugacy classes in a group of order $n$. The function $f(n)$ is quite well studied. Landau [52] showed that $f(n)\to\infty$ as $n\to\infty$. The first lower bound was obtained by Brauer [11] and Erdős and Turán [34], who showed that $f(n)\geq\log\log n$ (logarithms to base $2$). This was improved to $\epsilon\log n/(\log\log n)^{8}$ by Pyber [63]. The exponent $8$ in the previous expression was reduced to $7$ by Keller [49], and then to $3+\epsilon$ by Baumeister, Maróti and Tong-Viet [8], for any $\epsilon>0$. It is conjectured that a bound of the form $f(n)\geq C\log n$ holds for some constant $C$.
This is relevant to the recognition problem because of the following observation, which follows immediately from equation (6).
Proposition 4.3
An $n$-vertex graph $X$ with fewer than $\frac{n(f(n)-1)}{2}$ edges cannot be the commuting graph of a group, where $f(n)$ is the minimum number of conjugacy classes that a group of order $n$ can have.
At the other extreme, we can rule out very dense incomplete graphs by the $5/8$-theorem [43] for finite groups. More precisely, for any finite non-abelian group $G$ with $n$ elements:
$|C|=|\{(x,y)\mid x,y\in G\textrm{ and }xy=yx\}|\leq\frac{5}{8}\cdot n^{2},$
where $C\subset G\times G$ denotes the set of ordered pairs $(x,y)$ of commuting elements. Then, in the commuting graph $\Gamma(G)=(G,E)$ the number of edges
$|E|=\frac{|C|-n}{2}\leq\frac{5}{16}\cdot n^{2}-\frac{n}{2}\leq\frac{5}{8}\cdot{n\choose 2}.$
Putting it together, we have:
Proposition 4.4
An incomplete graph $X=(V,E)$ cannot be the commuting graph of a group if $|E|>\frac{5}{8}\cdot{|V|\choose 2}$.
4.2 The commuting graph and maximal abelian subgroups
For a finite group $G$, let $\mathcal{M}\subseteq 2^{G}$ denote the set of all maximal abelian subgroups of $G$. Associated with $G$ is the natural hypergraph $(G,\mathcal{M})$, where the hyperedges are precisely the maximal abelian subgroups of $G$.
Proposition 4.5
The commuting graph of a finite group determines the hypergraph of its maximal abelian subgroups, and, conversely the maximal abelian subgroups hypergraph of the group determines its commuting graph.
Proof 4
Let $G$ be a finite group. The maximal abelian subgroups are the maximal cliques in $\Gamma(G)$; and the edges of $\Gamma(G)$ are the $2$-element subsets of maximal abelian subgroups.$\qed$
Remark 1
Since the commuting graph $\Gamma(G)$ of a finite simple group $G$ uniquely determines the group [71], by Proposition 4.5 it follows that the set of maximal abelian groups of a finite simple group $G$ uniquely determines $G$.
For a finite group $G$, the number of maximal abelian subgroups is bounded by (${|G|}\choose{\log|G|}$ because every subgroup of $G$ has a generating set of size bounded by $\log|G|$.
This simple bound is tight apart from a constant in the exponent. To see this, consider the extraspecial group $G$ of order $p^{2n+1}$ and exponent $p$, where $p$ is an odd prime. The centre has order $p$, and $G/Z(G)$ is isomorphic to a $2n$-dimensional vector space $V$ over the field $F$ of $p$ elements, with the bilinear form from $V\times V$ to $F$ corresponding to the commutation map from $G/Z(G)\times G/Z(G)$ to $Z(G)$. Maximal abelian subgroups contain the centre, and correspond to maximal totally isotropic subspaces of $V$. It is known that the number of such subspaces is $\prod_{i=1}^{n}(p^{i}+1)$ (see [75]). As
$\prod_{i=1}^{n}(p^{i}+1)>p^{n(n+1)/2}>p^{n(2n+1)/4}=|G|^{n/4},$
it follows that the number of maximal abelian subgroups of the extraspecial $p$-group $G$ is $|G|^{\Omega(\log|G|)}$ for any odd prime $p$.
5 Commuting graphs of product groups
A natural approach to the problem of recognizing commuting graphs is to examine the connection between product decompositions of the commuting graph $\Gamma(G)$ and the direct product decomposition of the group $G$.
Let $G$ and $H$ be finite groups. Consider the commuting graph $\Gamma(G\times H)$ of their direct product. We recall the definition of strong products of graphs.
Definition 1
[44, Chapter 4, Page 35] Let $X=(V,E)$ and $X^{\prime}=(V^{\prime},E^{\prime})$ be simple undirected graphs. Their strong product, denoted $X\boxtimes X^{\prime}$, is a simple undirected graph with the cartesian product $V\times V^{\prime}$ as its vertex set and edges defined as follows: distinct pairs $(u,u^{\prime})$ and $(v,v^{\prime})$ are adjacent if and only if one of the following holds:
  • (a)
    $u=v$ and $(u^{\prime},v^{\prime})\in E^{\prime}$,
  • (b)
    $u^{\prime}=v^{\prime}$ and $(u,v)\in E$,
  • (c)
    $(u,v)\in E$ and $(u^{\prime},v^{\prime})\in E^{\prime}$.
The following proposition is immediate from the definition.
Proposition 5.1
For finite groups $G$ and $H$
$\Gamma(G\times H)=\Gamma(G)\boxtimes\Gamma(H).$
The strong product is both commutative and associative [44, Chapter 4], in the sense that $X\boxtimes Y\cong Y\boxtimes X$ and $(X\boxtimes Y)\boxtimes Z\cong X\boxtimes(Y\boxtimes Z)$.
By a non-trivial graph we mean a graph with more than one vertex. Given a graph $X$, a factorization $X\cong Y\boxtimes Z$ is non-trivial if both $Y$ and $Z$ are non-trivial. A graph $X$ is said to be a prime graph if it does not have a non-trivial factorization. (We warn the reader that this term is also used for graphs defined by arithmetic parameters of a finite group; see for example [55].)
It turns out that any connected graph $X$ has a unique factorization (up to isomorphism) into a strong product of prime graphs [44, Theorem 7.14, Pg 85]. Moreover, this prime factorization can be computed in polynomial time [35]. As a consequence, we obtain the following polynomial-time reduction.
Recall the a group $G$ is said to be indecomposable if it cannot be expressed as the direct product of two non-trivial groups.
Theorem 5.2
The problem of recognizing the commuting graphs of groups is polynomial-time reducible to the problem of recognizing the commuting graphs of indecomposable groups.
Proof 5
Suppose we have an algorithm $\mathcal{A}$ for recognizing the commuting graphs of indecomposable groups. Using $\mathcal{A}$ as subroutine, we present a polynomial-time algorithm for recognizing the commuting graphs of all finite groups.
Let $X=(V,E)$ be an undirected graph on $n$ vertices which is a purported commuting graph.
First, using the polynomial-time algorithm of Feigenbaum and Schäffer [35] we can factorize $X$ as
$X=X_{1}\boxtimes X_{2}\boxtimes\cdots\boxtimes X_{k},$
where each $X_{i}$ is a prime graph on at least two vertices. It follows that $k\leq\log n$. Now, for each subset $S\subseteq[k]$ of the prime graphs we combine them by taking the strong direct product to define the graph
$X_{S}=\boxtimes_{i\in S}X_{i}.$
Notice that any order in which the strong product of these graphs $X_{i}$ is computed yields the same graph, up to isomorphism.
Thus, in time which is polynomial in $n$, we have computed graphs $X_{S}$ for each subset $S$ of $[k]$. Now, we invoke the subroutine $\mathcal{A}$ that checks if $X_{S}$ is the commuting graph of an indecomposable group $G_{S}$, and if so, finds a labeling of the vertices of $X_{S}$ with elements of $G_{S}$ consistent with the commuting relations.
We can now check if the input graph $X$ is the commuting graph of a group with a straightforward dynamic programming strategy based on the following claim which directly follows from Proposition 5.1.
Claim 5.3
For any two disjoint subsets $S,S^{\prime}$ of $[k]$ such that $X_{S}$ and $X_{S^{\prime}}$ are the commuting graphs of groups $G_{S}$ and $G_{S^{\prime}}$ the graph $X_{S}\boxtimes X_{S^{\prime}}$ is the commuting graph of the direct product group $G_{S}\times G_{S^{\prime}}$.
Now, the algorithm works in stages, computing subsets $S$ of $[k]$ along with the graph $X_{S}$ and group $G_{S}$ such that $X_{S}=\Gamma(G_{S})$.
  1. (a)
    for stages 00 to $k$ do
  2. (b)
    Stage 00   we have subsets $S$ such that $G_{S}$ is an indecomposable group. We mark all such subsets $S$ as true. We mark the remaining subsets as false.
  3. (c)
    Stage $i+1$   For each pair of disjoint subsets $S$ and $S^{\prime}$ marked true in Stages $1,2,\ldots,i$, such that $S\cup S^{\prime}$ is marked false, we mark $S\cup S^{\prime}$ true and compute $X_{S\cup S^{\prime}}=X_{S}\boxtimes X_{S^{\prime}}$ and $G_{S\cup S^{\prime}}=G_{S}\times G_{S^{\prime}}$.
  4. (d)
    end-for
  5. (e)
    If $[k]$ is marked true then the input $X$ is the commuting graph of the group $G_{[k]}$ computed above.
The above description checks if $X$ is the commuting graph of a group with at most $2^{k}\leq n$ calls to the subroutine $\mathcal{A}$ and the running time of the remaining computation is clearly polynomially bounded in $n$$\qed$
From Theorem 5.2 we can easily deduce from the Solomon–Woldar theorem [71] that direct products of simple groups too have uniquely determined commuting graphs.
Corollary 5.4
If $G$ is a direct product of simple groups, then $G$ is uniquely determined by its commuting graph $\Gamma(G)$.
We do not know if the commuting graph of an indecomposable group $G$ is necessarily a prime graph. However, if $G$ is simple then it turns out, as we show below, that $\Gamma(G)$ is a prime graph.
Lemma 5.5
The commuting graph of a finite simple group is a prime graph under the strong product.
Proof 6
We use the fact that, if $G$ is a non-abelian simple group and $g\in G\setminus\{1\}$, then there exists $h\in G$ such that $\langle g,h\rangle=G$ [42]. Now if $\langle g,h\rangle=G$, then
  • (a)
    $g$ and $h$ are nonadjacent in the commuting graph (since $G$ is non-abelian);
  • (b)
    $g$ and $h$ have no non-identity common neighbour in the commuting graph (since a common neighbour would belong to $Z(G)$, but $Z(G)=\{1\}$.
Now suppose for a contradiction that $\Gamma(G)$ is the strong product of two non-trivial graphs with vertex sets $A$ and $B$ (i.e., with $|A|>1$ and $|B|>1$). Then we can identify $G$, as the vertex set of $\Gamma(G)$, with the Cartesian product $A\times B$. Suppose that $(a,b)$ is the identity of $G$. Then $(a,b)$ is joined to all other vertices in the commuting graph.
It follows that $a$ is joined to every other vertex in $A$, and $b$ to every other vertex in $B$; therefore, for all $x\in A\setminus\{a\}$, $y\in B\setminus\{b\}$, the three vertices $(a,y)$, $(x,b)$ and $(x,y)$ are adjacent to each other in the commuting graph $\Gamma(G)$.
Choose $y\in B\setminus\{b\}$, and suppose that $\langle(a,y),(u,v)\rangle=G$. Now
  • (a)
    if $u=a$ then $(a,y)$ and $(a,v)$ are both joined to $(x,b)$ for any $x\in A\setminus\{a\}$;
  • (b)
    if $v=b$ then $(a,y)$ and $(u,b)$ are joined;
  • (c)
    if neither of the above, then $(a,y)$ and $(u,v)$ are both joined to $(u,b)$.
Each case is contradictory; so our assumption that $\Gamma(G)$ is the strong product of two non-trivial graphs is false, and the theorem is proved.$\qed$
Remark 2
The proof of Lemma 5.5 only requires that $Z(G)=1$ and that any non-identity element is contained in a $2$-element generating set. These assumptions are valid in any almost simple group $G$ with simple normal subgroup $S$ such that $G/S$ is cyclic [17].
As a consequence of Lemma 5.5, we obtain the following variant of Theorem 5.2.
Theorem 5.6
The problem of recognizing the commuting graph of a direct product of simple groups is polynomial-time reducible to the problem of recognizing the commuting graph of a simple group.
The algorithm for the reduction in the above theorem is much simpler and more efficient than in Theorem 5.2. For we only need to consider the prime factors $X_{i}$, in the factorization $X=\boxtimes_{i=1}^{k}X_{i}$, and not all the strong products $X_{S}$ for arbitrary subsets $S\subseteq[k]$.
6 Commuting graphs of semidirect products
In this section we explore whether we can recognize the commuting graphs of semidirect products $G=H\rtimes K$ if $H$ and $K$ are both from group classes whose commuting graphs are easily recognizable.
For example, consider the commuting graph of the dihedral group $D_{2n}$. Let $D_{2n}=\langle a,b\rangle$ where $a^{n}=b^{2}=1$ and $bab=a^{-1}$. For $n$ odd, $1$ is the only dominating vertex, there is an $n$-clique corresponding to $\langle a\rangle$ and the $ba^{i}$ correspond to pendant vertices. Thus, the commuting graph of $D_{2n}$ is easily recognizable. For $n$ even, it is a bit different with $b^{n/2}$ being the other dominating vertex.
More generally, consider the generalized dihedral group $\langle A,t\rangle$ defined earlier, where $A$ is abelian of order $2n$, $n$ odd. Its commuting graph is isomorphic to that of the dihedral group $D_{2n}$ of the same order. Hence, as observed above, we can recognize that a graph is the commuting graph of a generalized dihedral group, but cannot decide which one.
Seeking a generalization of this example we consider Frobenius groups.
6.1 Commuting graphs of Frobenius groups
In this subsection we demonstrate that we can recognise from its commuting graph that a group $G$ is a Frobenius group. These groups were defined earlier. Let $G$ be a Frobenius group with kernel $N$ and complement $H$.
Thompson [76] proved that the Frobenius kernel is nilpotent, and Zassenhaus [80] worked out the detailed structure of a Frobenius complement. Everything we need about Frobenius groups is explained in standard group theory texts  [61, 45, 69].
In the commuting graph of a group $G$, the identity is a dominating vertex (that is, joined to all others); indeed, any vertex in the centre is dominating, so if $Z(G)$ is non-trivial then the commuting graph is $2$-connected.
Lemma 6.1
  1. (a)
    A Frobenius kernel has non-trivial centre.
  2. (b)
    A Frobenius complement has non-trivial centre.
Proof 7
(a) By Thompson’s theorem [76, Theorem 1], a Frobenius kernel is nilpotent, and nilpotent groups have a non-trivial centre.
(b) Let $H$ be a Frobenius complement. By [45, Theorem 6.3, Corollary 6.10], if $H$ has even order, then it has a unique element of order $2$, which lies in the centre; while, if $H$ has odd order, then it has a unique subgroup of order $p$ for every $p$ dividing $|H|$. In the latter case, choose $p$ to be the smallest prime divisor of $H$, and $P=\langle a\rangle$ the unique subgroup of order $p$. Then $P$ is a normal subgroup of $H$. For any $h\in H$, we have $a^{h}=a^{m}$ for some $m$ with $11\leq m\leq p-$ 1. If $m>1$, let $q$ be a prime divisor of $m$; then some power of $h$ has order $q$, a contradiction. So $a^{h}=a$ for all $h\in H$, whence $a\in Z(H)$$\qed$
Theorem 6.2
Let $G$ be a group of order $nk$, where $n>1$, $k>1$, and $n$ and $k$ are relatively prime. Then $G$ is a Frobenius group with Frobenius complement of order $k$ if and only if its commuting graph $\Gamma(G)$ satisfies the following conditions:
  • (a)
    there is a unique dominating vertex $1$ in $\Gamma(G)$, where $1$ corresponds to the identity element of $G$.
  • (b)
    $\Gamma(G)\setminus\{1\}$ has a component of size $n-1$ and $n$ components of size $k-1$, and each component has a dominating vertex.
Proof 8
Suppose that $G$ is a Frobenius group with kernel $N$ of order $n$ and complement $H$ of order $k$. The identity is a dominating vertex. Also non-identity vertices in the kernel do not commute with non-identity vertices in a complement, and non-identity vertices in different complements do not commute with each other. Thus the components of $\Gamma(G)\setminus\{1\}$ are as stated, and Lemma 6.1 shows that they have dominating vertices.
Conversely, suppose that the commuting graph $\Gamma(G)$ has properties (a) and (b). If $C$ is a component with a dominating vertex $c$, then the centralizer $C_{G}(c)$ is equal to $C\cup\{1\}$, which is thus a subgroup of $G$. Let $N$ be the subgroup containing the component of size $n-1$, and let $H_{1},\ldots,H_{n}$ be the subgroups containing the other components. Note that $|N|=n$ and $|H_{i}|=k,1\leq i\leq n$.
Since $\Gamma(G)$ is invariant under automorphisms of $G$, we have an action of $G$ on the set $\Omega$ of components of $\Gamma(G)\setminus\{1\}$ of size $k-1$ by conjugation. We show that this action satisfies the conditions for a Frobenius group given above.
Choose a prime $p$ dividing $k$. Since $n$ and $k$ are relatively prime, $p$ does not divide $n$. So the subgroup $N$ cannot contain a Sylow $p$-subgroup; but each of $H_{1},\ldots,H_{n}$ contains such a subgroup. By the conjugacy part of Sylow’s theorem, $G$ acts transitively on $\Omega$.
For $1\leq i\leq n$, each element of $H_{i}\setminus\{1\}$ acting by conjugation, fixes itself and so fixes the component induced by $H_{i}\setminus\{1\}$. That is, a non-identity element of $G\setminus N$, acting by conjugation, fixes itself, and so fixes the component containing it. We must show that it fixes no other component. To that end, we count the fixed points of elements of $G$ acting on $\Omega$. The identity fixes $n$; each non-identity element of $N$ fixes some non-negative number; and each of the remaining elements $G\setminus N$ fix a positive number of elements. So the sum of the fixed point numbers is at least $n+n(k-1)=nk$. But, by the Orbit-counting Lemma, this sum is equal to $|G|=nk$, since $G$ is transitive. Consequently, non-identity elements of $N$ do not fix any point of $\Omega$, and the remaining non-identity elements fix exactly one point each. So $G$ is a Frobenius group.$\qed$
Suppose $X$ is a graph given as input such that $X=\Gamma(G)$ for some group $G$. Then Theorem 6.2 gives us a polynomial-time test for checking if the group $G$ is a Frobenius group.
Remark 3
Above we assumed that our graph is a commuting graph. However, it is easy to construct a graph satisfying the conditions of Theorem 6.2 which is not a commuting graph. Can we find necessary and sufficient conditions for $\Gamma$ to be the commuting graph of a Frobenius group?
7 Recognizing commuting graphs of extraspecial groups
In this section we design a polynomial-time algorithm that takes a simple undirected graph $X=(V,E)$, with $n=p^{2r+1}$ vertices, $r\geq 2$ and $p$ an odd prime, as input and determines if $X$ is the commuting graph of an extraspecial group $G$ of order $p^{2r+1}$ (see Section 2) for definition). The algorithm will also compute a certifying bijection $v\mapsto g_{v}$ labeling each vertex $v\in V$ by a unique group element $g_{v}\in G$ such that the commuting relations hold.
There are exactly two non-isomorphic extraspecial groups of order $p^{2r+1}$, for each $r\geq 1$ and each prime $p$. These two groups, which we denote as $G_{1}$ and $G_{2}$, are described below.
  • (a)
    The first group $G_{1}$ is generated by $12r+$ 1 elements $G_{1}=\langle z,x_{i},y_{i},1\leq i\leq r\rangle$ such that $[x_{i},x_{j}]=1$, $[y_{i},y_{j}]=1$ for all $i,j$, and $[x_{i},y_{j}]=1$ for all $i\neq j$, $[x_{i},y_{i}]=z$ for all $i$, and $x_{i}^{p}=y_{i}^{p}=z^{p}=1$ for all $i$. The centre $Z(G_{1})=\langle z\rangle$.
  • (b)
    The second group $G_{2}$ also has $12r+$ 1 generators $G_{2}=\langle z,x_{i},y_{i},1\leq i\leq r\rangle$ such that $[x_{i},x_{j}]=1$, $[y_{i},y_{j}]=1$ for all $i,j$, and $[x_{i},y_{j}]=1$ for all $i\neq j$. And $[x_{i},y_{i}]=z$ for all $i$, and $x_{i}^{p}=1=z^{p}$, $y_{i}^{p}=z$ for all $i$. The centre $Z(G_{2})=\langle z\rangle$.
Each element $g\in G_{1}\setminus\{1\}$ has order $p$. On the other hand, each element $g\in G_{2}\setminus\{1\}$ has order either $p$ or $p^{2}$. Hence, $G_{1}$ and $G_{2}$ are non-isomorphic.
We refer to Gorenstein [40, Chapter 5] for more about extraspecial groups.
We now analyze the commuting graph $\Gamma(G)$ for $G\in\{G_{1},G_{2}\}$ of order $n=p^{2r+1}$. The vertices of $\Gamma(G)$ corresponding to the centre $Z(G)=Z=\langle z\rangle$ are precisely the $p$ vertices of degree $p^{2r+1}-1$ each. In either case ($G\in\{G_{1},G_{2}\}$) we have:
  • (a)
    The quotient group $G/Z$ is an elementary abelian $p$-group of order $p^{2r}$.
  • (b)
    The centre $Z$ can be identified with the finite field $\mathbb{F}_{p}$ and $G/Z$ can be identified with the $2r$-dimensional vector space $V(2r,p)=\mathbb{F}_{p}^{2r}$.
Furthermore, as $Z=G^{\prime}$, for $x,y\in G$ we can write their commutator $[x,y]=z^{i_{x,y}}$ for some $10\leq i_{x,y}\leq p-$ 1. Hence, the commutation map
$(xZ(G),yZ(G))\mapsto[x,y]$
from $G/Z\times G/Z\to Z$ can be identified with a bilinear form $\langle,\rangle:V(2r,p)\times V(2r,p)\to\mathbb{F}_{p}$ that satisfies the following:
  • (a)
    It is symplectic: We have $\langle x,y\rangle=-\langle y,x\rangle$ for all $x,y\in V(2r,p)$. In particular, $\langle x,x\rangle=0$ for all $x\in V(2r,p)$.
  • (b)
    It is non-degenerate: there is no non-zero $u\in V(2r,p)$ such that $\langle u,v\rangle=0$ for all $v\in V(2r,p)$.
Definition 2
The vector space $V(2r,p)$ equipped with a symplectic form $\langle$,$\rangle$ , ⟩ is a symplectic space.
If $\langle u,v\rangle=0$ we say $u$ is orthogonal to $v$. For a subset $S$ of the symplectic space $V(2r,p)$ let
$S^{\perp}=\{u\in V(2r,p)\mid\langle u,v\rangle=0\text{ for all }v\in S\}$
denote the subspace of $V(2r,p)$ consisting of vectors orthogonal to each vector in $S$.
Remark 4
We will describe an elementary self-contained polynomial-time algorithm for recognising commuting graphs of extraspecial groups. We have an alternative algorithm, based on the Buekenhout-Shult theorem [16] classifying certain point-line geometries, that exploits the finite geometry structure of the quotient group $G/Z(G)$. However, we present the former in the interest of keeping the proof entirely self-contained in the context of this paper.
Lemma 7.1
Given as input the multiplication table of an extraspecial group $G$ of order $p^{2r+1}$, in time polynomial in $p^{2r+1}$ we can construct its associated symplectic form $\langle$,$\rangle$ , ⟩, as an explicit bilinear map from $\mathbb{F}_{p}^{2r}\times\mathbb{F}_{p}^{2r}\to\mathbb{F}_{p}$.
Proof 9
Let $n=p^{2r+1}$. First, we compute the centre $Z=\langle z\rangle$ of $G$ from the multiplication table. In order to compute $G/Z$ we can find its multiplication table from the given multiplication table of $G$. We can find a decomposition of $G/Z$ as the direct product of $2r$ cyclic groups of order $p$ by finding independent generators for $G/Z$ using a simple algorithm. Suppose $g_{1},g_{2},\ldots,g_{j}\in G/Z$ are independent (i.e. the identity can be expressed as a product of their powers only if the exponents are all zero). If $j<2r$ then the subgroup $H_{j}=\langle g_{1},g_{2},\ldots,g_{j}\rangle$ which is of size $p^{j}$ can be listed in time polynomial in $n$. We can pick $g_{j+1}$ as any element in $G\setminus H_{j}$ and continue the process until we pick $2r$ such independent generators which will generate $G/Z$. That gives us an efficient identification of $G/Z$ with $V(2r,p)$. For $xZ,yZ\in G/Z$ we will have corresponding vectors $v_{x},v_{y}\in V(2r,p)$. If the commutator $[x,y]=z^{i_{x,y}}$ then we can define the symplectic form as $\langle v_{x},v_{y}\rangle=i_{x,y}$, where $i_{x,y}\in\mathbb{F}_{p}$. This completes the proof. $\qed$
Definition 3
The symplectic graph of the symplectic space $V(2r,p)$ (equipped with symplectic form $\langle$,$\rangle$ , ⟩) is an undirected simple graph with $V(2r,p)$ as vertex set such that for each pair of distinct vectors $u,v\in V(2r,p)$, $(u,v)$ is an undirected edge in the graph precisely when $\langle u,v\rangle=0$.
In particular, let $S_{1}$ and $S_{2}$ denote the associated symplectic graph of the extraspecial groups $G_{1}$ and $G_{2}$, respectively.
Lemma 7.2
  1. (a)
    The commuting graphs $\Gamma(G_{1})$ and $\Gamma(G_{2})$ are isomorphic.
  2. (b)
    The associated symplectic graphs $S_{1}$ and $S_{2}$ are isomorphic.
Proof 10
Recall, as defined at the beginning of this section, that both $G_{1}=\langle x_{1},x_{2},\ldots,x_{r},y_{1},y_{2},\ldots,y_{r},z\rangle$ and $G_{2}=\langle x_{1},x_{2},\ldots,x_{r},y_{1},y_{2},\ldots,y_{r},z\rangle$. The commuting relations between the generators are identical. The only difference is that in $G_{2}$ each $y_{i}$ is of order $p^{2}$, and in $G_{1}$ they are all order $p$.
Because of the commutation relation $[x_{i},y_{i}]=z$ in both groups, where $Z=\langle z\rangle$ is the centre, each element in both groups can be expressed as products of the form
$x_{1}^{e_{1}}x_{2}^{e_{2}}\cdots x_{r}^{e_{r}}y_{1}^{f_{1}}y_{2}^{f_{2}}\cdots y_{r}^{f_{r}}\cdot z^{c},$
where $10\leq c,e_{i},f_{i}\leq p-$ 1 for all $i$. Suppose now we have two elements $x=\prod_{i=1}^{r}x_{i}^{e_{i}}\cdot\prod_{i=1}^{r}y_{i}^{f_{i}}\cdot z^{c}$ and $y=\prod_{i=1}^{r}x_{i}^{e^{\prime}_{i}}\cdot\prod_{i=1}^{r}y_{i}^{f^{\prime}_{i}}\cdot z^{c^{\prime}}$. Then, using the commutation relation $[x_{i},y_{i}]=z$, which holds in both $G_{1}$ and $G_{2}$, we can check that $x$ and $y$ commute if and only if $\sum_{i=1}^{r}e_{i}f^{\prime}_{i}=\sum_{i=1}^{r}f_{i}e^{\prime}_{i}(\operatorname{mod}p)$. It follows that $(x,y)$ is an edge in $\Gamma(G_{1})$ if and only if it is an edge in $\Gamma(G_{2})$. Hence the commuting graphs $G_{1}$ and $G_{2}$ are isomorphic.
Let $V_{1}$ and $V_{2}$ denote the symplectic spaces corresponding to the groups $G_{1}$ and $G_{2}$, respectively. For $G\in\{G_{1},G_{2}\}$, as already observed $G/Z$ is isomorphic to $V(2r,p)$ (as an additive group). Let $x_{i}Z\mapsto\bar{x}_{i}$ and $y_{i}Z\mapsto\bar{y}_{i}$, $1\leq i\leq r$, under this isomorphism, for $\bar{x}_{i},\bar{y}_{i}\in V(2r,p),1\leq i\leq r$. Then, $xZ$ for
$x=\prod_{i=1}^{r}x_{i}^{e_{i}}\cdot\prod_{i=1}^{r}y_{i}^{f_{i}}\cdot z^{c}$
maps to $\bar{x}=\sum_{i=1}^{r}(e_{i}\bar{x}_{i}+f_{i}\bar{y}_{i})$. Likewise, $y=\prod_{i=1}^{r}x_{i}^{e^{\prime}_{i}}\cdot\prod_{i=1}^{r}y_{i}^{f^{\prime}_{i}}\cdot z^{c^{\prime}}$ maps to $\bar{y}=\sum_{i=1}^{r}(e^{\prime}_{i}\bar{x}_{i}+f^{\prime}_{i}\bar{y}_{i})$. Therefore, we can evaluate $\langle\bar{x},\bar{y}\rangle$ using the fact that $\langle\bar{x}_{i},\bar{x_{j}}\rangle=0$, $\langle\bar{y}_{i},\bar{y_{j}}\rangle=0$, for all $i\neq j$ and for all $i$ we have $\langle\bar{x}_{i},\bar{y}_{i}\rangle=1$ and $\langle\bar{y}_{i},\bar{x_{i}}\rangle=-1$.
$\langle\bar{x},\bar{y}\rangle=\sum_{i=1}^{r}(e_{i}f^{\prime}_{i}-e^{\prime}_{i}f_{i})$
It follows that the vertices $\bar{x}$ and $\bar{y}$ are adjacent in the symplectic graph $S_{1}$ if and only if they are adjacent in $S_{2}$. $\qed$
We now show that the symplectic graph of an extraspecial group $G$ can be computed in polynomial time, given just its commuting graph $\Gamma(G)$ as input.
Lemma 7.3
Given as input the commuting graph $\Gamma(G)$, of an extraspecial group $G$ of order $p^{2r+1}$, $p$ an odd prime, we can compute in polynomial time an undirected graph $X_{s}=(V_{s},E_{s})$ with $|V_{s}|=p^{2r}$ such that $X_{s}$ is the symplectic graph associated with $G$.
Proof 11
In the commuting graph $\Gamma(G)$, which has vertex set $G$, there are exactly $p$ dominant vertices, corresponding to the $p$ elements of the centre $Z$. Let $Z$ also denote this subset of vertices of $\Gamma(G)$.
The closed neighborhood of a vertex $v\in G$ is defined as
$\bar{N}(v)=\{u\in V\mid(u,v)\in E\}\cup\{v\}.$
[…]
Claim 7.4
[…]
To see this claim, let $h\in G\setminus Z$. Consider the subgroup $H$ of $G$ generated by $Z\cup\{h\}$, where the centre $Z=\langle z\rangle$. Then $|H|=p^{2}$ and $H=\{z^{i}h^{j}\mid 0\leq i,j\leq p-1\}$. Furthermore,
$H\setminus Z=\{z^{i}h^{j}\mid 0\leq i\leq p-1,1\leq j\leq p-1\}.$
Now, as $\gcd(j,p)=1$ for nonzero $j\leq p-1$, it follows that any $x\in G\setminus Z$ commutes with $z^{i}h^{j}\in H\setminus Z$ if and only if $x$ commutes with $h$. Hence, every element of $H\setminus Z$ is a closed twin of $h$ which implies that $H\setminus Z\subseteq[h]$.
Let $g\notin G\setminus H$. We will show that $g\notin[h]$ by exhibiting an element of $G\setminus Z$ that commutes with $h$ but not with $g$. That will show that $H\setminus Z=[h]$, which will complete the proof of the claim.
As $g\notin H$, we have $Zg\notin H/Z$. Now, identifying $G/Z$ with the symplectic space, the coset $Zh$ maps to a vector $\bar{x}_{1}$ of $V(2r,p)$ and $H/Z$ maps to the $1$-dimensional subspace of $V(2r,p)$ spanned by $\bar{x}_{1}$. We can extend this vector to a symplectic basis $\bar{x}_{i},\bar{y}_{i},1\leq i\leq r$ of $V(2r,p)$ with the relations $1=\langle x_{i},y_{i}\rangle=-\langle y_{i},x_{i}\rangle$ along with the other relations (as in the proof of Lemma 7.2).
As $Zg\not\in H/Z$, its image $u\in V(2r,p)$ in the symplectic space is of the form
$u=\sum_{i=1}^{r}(\alpha_{i}\bar{x}_{i}+\beta_{i}\bar{y}_{i}),$
where either $\alpha_{i}\neq 0$ for some $i\neq 1$ or some $\beta_{j}\neq 0,1\leq j\leq r$. In the former case, $\langle u,\bar{x}_{i}\rangle\neq 0$ but $\langle\bar{x}_{1},\bar{x}_{i}\rangle=0$, and in the latter case $\langle\bar{x}_{1},\bar{y}_{j}\rangle=0$ but $\langle u,\bar{y}_{j}\rangle\neq 0$.
It follows that there is an element $x\in G\setminus Z$ that commutes with $h$ but not with $g$. Hence, $g\notin[h]$.
Thus, the equivalence classes defined by the closed-twins equivalence identifies all $p^{2}-p$ size vertex subsets of $G\setminus Z$ that corresponds to $H\setminus Z$ for each subgroup $H$ of order $p^{2}$ such that $Z\leq H\leq G$. The number of such equivalence classes is
$\frac{p^{2r+1}-p}{p^{2}-p}=1+p+p^{2}+\cdot+p^{2r-1}.$
The preceding discussion immediately implies the following.
Claim 7.5
For any two order $p^{2}$ subgroups $H$ and $H^{\prime}$ such that $Z\leq H\leq G$ and $Z\leq H^{\prime}\leq G$, the bipartition $H\setminus Z$ and $H^{\prime}\setminus Z$ either forms a complete bipartite graph in $\Gamma(G)$ or there are no edges between the parts.
To obtain the corresponding symplectic graph $X_{s}=(V_{s},E_{s})$ we apply the following three steps to $\Gamma(G)$.
  • (a)
    Collapse each equivalence class in $G\setminus Z$ into a single vertex. This gives rise to a set $V_{c}$ of $\sum_{i=0}^{2r-1}p^{i}$ many vertices. The edges between vertices in $V_{c}$ are naturally inherited from $\Gamma(G)$ by Claim 7.5.
  • (b)
    Include a new vertex $v_{0}$ corresponding to the 00 element of $V(2n,p)$; this corresponds to $p$ dominant vertices of $X$.
  • (c)
    The process of collapsing the closed-twin equivalence classes identifies vectors $v\in V(2r,p)$ with all the $p-1$ non-zero scalar multiples $\alpha v,\alpha\in\mathbb{F}^{*}_{p}$. We restore the $p-1$ copies by replacing each $v\in V_{c}$ by $p-1$ copies. The edges between these vertices are naturally inherited.
By construction, $X_{s}$ is the symplectic graph of the symplectic form on $V(2r,p)$ defined by the commuting graph $\Gamma(G)$. $\qed$
7.1 Recognizing the symplectic graph
We now show that the symplectic graph of a non-degenerate symplectic form on $V(2r,p)$ can be recognized in polynomial time.
Theorem 7.6
Given a simple undirected graph $X=(V,E)$ on a vertex set $V$ of size $p^{2r}$, for an odd prime $p$, in time polynomial in the size of $X$ we can recognize if $X$ is the associated symplectic graph of an extraspecial group of order $p^{2r+1}$. Furthermore, the algorithm will also compute a bijection from $V$ to $V(2r,p)$ and determine an associated symplectic form $\langle$,$\rangle$ , ⟩.
Proof 12
The somewhat detailed proof is structured as a series of claims followed by a description of the algorithm and its analysis.
Let $X=(V,E)$ be a graph with $p^{2r}$ vertices. Our algorithm will be recursive, with correctness proof based on an inductive argument. We know that the symplectic space $V(2r,p)$ has a symplectic basis of $2r$ vectors of the form $\bar{x}_{i},\bar{y}_{i},1\leq i\leq n$ such that
  • (a)
    $\langle\bar{x}_{i},\bar{y}_{i}\rangle=1$ for all $i$.
  • (b)
    $\langle\bar{x}_{i},\bar{y}_{j}\rangle=0$ for all $i\neq j$.
  • (c)
    $\langle\bar{x}_{i},\bar{x}_{j}\rangle=0$ for all $i,j$.
  • (d)
    $\langle\bar{y}_{i},\bar{y}_{j}\rangle=0$ for all $i,j$.
We first note that given a symplectic form $\langle$,$\rangle$ , ⟩ on $V(2r,p)$, we can construct a symplectic basis for $V(2r,p)$ with the following greedy procedure.
  • (a)
    Pick $\bar{x}_{1}\neq 0$ in $V(2r,p)$ arbitrarily.
  • (b)
    Since $\langle$,$\rangle$ , ⟩ is non-degenerate, we can find a vector $\bar{u}\in V(2r,p)$ such that $\langle\bar{x}_{1},\bar{u}\rangle=\alpha\neq 0$. Define $\bar{y}_{1}=\alpha^{-1}\bar{u}$ such that $\langle\bar{x}_{1},\bar{y}_{1}\rangle=1$.
  • (c)
    Then notice that the subspace $\{\bar{x}_{1},\bar{y}_{1}\}^{\perp}$ is a $22r-$ 2 dimensional symplectic space $V^{\prime}$ (w.r.t. the same symplectic form). We can continue with the basis construction, recursively applying the above two steps to $V^{\prime}$.
Now, suppose $X=(V,E)$ is the graph given as input on $p^{2r}$ vertices. If $X$ is indeed the symplectic graph of $(V,2r,p)$ for some symplectic form $\langle$,$\rangle$ , ⟩, we will describe a polynomial-time recursive procedure that recovers a bijection from $V$ to $V(2r,p)$ along with a symplectic form $\langle$,$\rangle$ , ⟩ as witness for $X$ being the associated symplectic graph. We will then argue that if $X$ is not a symplectic graph then this procedure detects that.
Essentially, the problem is to efficiently simulate the above two-step basis construction procedure given only the graph $X=(V,E)$ as input. As the symplectic space is non-degenerate, the zero vector 00 is the only vertex in $X$ adjacent to all others and is easily identified. Let $x_{1}$ be any other vertex in $V$ and we choose a non-zero vector $\bar{x}_{1}\in V(2r,p)$ as its image. Next, we will choose $y_{1}$ as any vertex not adjacent to $x_{1}$ in the graph $X$ and choose a non-zero vector $\bar{y}_{1}\in V(2r,p)$ as its image. Let
$V^{\prime}=\{v\in V\mid(v,x_{1})\in E\text{ and }(v,y_{1})\in E\}.$
That is, $V^{\prime}$ is the common neighborhood of $x_{1}$ and $y_{1}$. We can easily identify $V^{\prime}$ in $X$. Let $X^{\prime}$ be the subgraph of $X$ induced by the vertex subset $V^{\prime}$. We have the following simple observation.
Claim 7.7
If $X$ is the symplectic graph of a $2r$-dimensional symplectic space over $\mathbb{F}_{p}$ then the graph $X^{\prime}$ is the symplectic graph of a symplectic space of dimension $22r-$ 2 over $\mathbb{F}_{p}$.
The algorithm will now recursively check that $X^{\prime}$ is indeed the symplectic graph of a $22r-$ 2-dimensional symplectic space $V(2r-2,p)$ over $\mathbb{F}_{p}$. If so, it will compute a labeling of the vertices of $V^{\prime}$ by linear combinations $\sum_{i=2}^{r}(\alpha_{i}\bar{x}_{i}+\beta_{i}\bar{y}_{i}),\alpha_{i},\beta_{i}\in\mathbb{F}_{p}$ that is consistent with the orthogonality relation of a symplectic form $\langle$,$\rangle$ , ⟩. Here the basis vectors computed by the algorithm are $\bar{x}_{i},\bar{y}_{i},2\leq i\leq r$. At the end of the recursive call, the algorithm computes a bijection from $V^{\prime}$ to $V(2r-2,p)$ that labels each vertex in $V^{\prime}$ by a distinct linear combination of the basis vectors.
Now, the remaining task for the algorithm is to find a labeling of the vertices in $V\setminus V^{\prime}$ consistent with a symplectic form on $V(2r,p)$, assuming $X$ is a symplectic graph.
For each vertex $v$ we will denote its vector label by $\bar{v}$. For a labeled subset $S$ of vertices of $V$ we will use $\bar{S}\subseteq V(2r,p)$ to denote the set of vectors labeling vertices in $S$.
Claim 7.8
A vertex $v\in V\setminus V^{\prime}$ can be labeled by a non-zero vector $\alpha\bar{x}_{1}+\beta\bar{y}_{1}$ in $V(2r,p)$, $\alpha,\beta\in\mathbb{F}_{p}$, if and only if $(v,u)\in E$ for all $u\in V^{\prime}$.
Proof of Claim. Consider the symplectic graph of $V(2r,p)$. Let $V(2r-2,p)$ denote the subspace spanned by $\bar{x}_{i},\bar{y}_{i},2\leq i\leq n$. Clearly, every vector of the form $\alpha\bar{x}_{1}+\beta\bar{y}_{1},\alpha,\beta\in\mathbb{F}_{p}$ is orthogonal to each vector in $V(2r-2,p)$. Conversely, consider a vector $\alpha\bar{x}_{1}+\beta\bar{y}_{1}+v\in V(2r,p)$, where $v\in V(2r-2,p)$ is non-zero. Since $V(2r-2,p)$ is non-degenerate, there is a $u\in V(2r-2,p)$ such that $\langle v,u\rangle\neq 0$ which implies $\langle\alpha\bar{x}_{1}+\beta\bar{y}_{1}+v,u\rangle=\langle v,u\rangle\neq 0$.
Thus, the algorithm can find a subset $V_{12}$ of precisely $p^{2}-1$ many such vertices in $V$ that are adjacent to all of $V^{\prime}$, of which we have already labeled two vertices as $\bar{x}_{1}$ and $\bar{y}_{1}$. We analyze the subgraph $X_{12}$ of $X$ induced by $V_{12}$.
Claim 7.9
The subgraph $X_{12}$ induced by $V_{12}$ consists of a disjoint union of $p+1$ many $(p-1)$-size cliques, $C_{1},C_{2},\ldots,C_{p+1}$. Furthermore, the labeling procedure must label the vertices of each $C_{i}$ from the set $\{\alpha\bar{u}\mid\alpha\in\mathbb{F}^{*}_{p}\}$ of $p-1$ vectors, for some non-zero vector $\bar{u}\in\mathop{\mathrm{Span}}(\bar{x}_{1},\bar{y}_{1})$.
Proof of Claim. It is clear from the definition of $V_{12}$ that the vertices in it must be labeled by nonzero vectors in $\mathop{\mathrm{Span}}(\bar{x}_{1},\bar{y}_{1})$. The proof follows easily by observing that a non-zero vector $\alpha\bar{x}_{1}+\beta\bar{y}_{1}$ is orthogonal only to itself and its scalar multiples and to no other vectors in $\mathop{\mathrm{Span}}(\bar{x}_{1},\bar{y}_{1})$.
By the above claim, we can consistently label each vertex $u\in V_{12}$ by a vector $\bar{u}\in\mathop{\mathrm{Span}}(\bar{x}_{1},\bar{y}_{1})$ assuming $X$ is a symplectic graph.
Claim 7.10
If $X=(V,E)$ is the symplectic graph of the symplectic space $V(2r,p)$ and $V^{\prime}$ corresponds to the symplectic subspace $V(2r-2,p)$ spanned by $\{\bar{x}_{i},\bar{y}_{i}\mid 2\leq i\leq r\}$ then for any vertex $u\in V_{12}$ the vertex subset
$V[u]=\{v\in V\setminus V^{\prime}\mid(v,u)\in E\}$
has to be labeled by the vectors
$\{\alpha\bar{u}+\bar{v}\mid\alpha\neq 0,\bar{v}\in\mathop{\mathrm{Span}}(\cup_{i=2}^{r}\{\bar{x}_{i},\bar{y}_{i}\})\}$
of $V(2r,p)$.
Proof of Claim. Suppose $v\in V\setminus V^{\prime}$ is labeled by the vector $\bar{a}+\bar{b}$, where $\bar{a}\in\mathop{\mathrm{Span}}(\bar{x}_{1},\bar{y}_{1})$ and $\bar{b}\in V(2r-2,p)$. As $v\in V[u]$ we have
$\langle\bar{a}+\bar{b},\bar{u}\rangle=0.$
As $\langle\bar{b},\bar{u}\rangle=0$ it implies that $\langle\bar{a},\bar{u}\rangle=0$ which means $\bar{a}=\alpha\bar{u}$ for some non-zero scalar $\alpha$. Conversely, any vertex labeled by $\alpha\bar{u}+\bar{b}$ for $\bar{b}\in V(2r-2,p)$ has to lie in $V[u]$ if the labeling is consistent with the symplectic graph $X$.
It follows that $|V[u]|=(p-1)\cdot p^{2r-2}$. Furthermore, if vertices $u$ and $v$ are in the same $(p-1)$-clique in the graph $X_{12}$ then $V[u]=V[v]$. As $(p+1)(p-1)\cdot p^{2r-2}=p^{2r}-p^{2r-2}=|V\setminus V^{\prime}|$, it follows that the different $V[u]$ are all pairwise disjoint and form a partition of $V\setminus V^{\prime}$. Consequently, we can identify the $p+1$ many equisized vertex subsets $V[u]$ which form a partition of $V\setminus V^{\prime}$.
Labelling vertices in $V[u]$
By Claim 7.10, the vertices of $V[u]$ have to be labeled by vectors from $\{\alpha\bar{u}+\bar{v}\mid v\in V^{\prime},\alpha\neq 0\}$. We will examine the adjacencies in the subgraph induced by $V[u]$ and also the adjacencies across the different $V[u]$ and $V[u^{\prime}]$ to determine a consistent labeling. Let $X[u]$ denote the subgraph of $X$ induced by $V[u]$.
For distinct vertices $v_{1},v_{2}\in V^{\prime}$ notice that
$\langle\bar{v}_{1},\bar{v}_{2}\rangle=\langle\alpha\bar{u}+\bar{v}_{1},\beta\bar{u}+\bar{v}_{2}\rangle.$ (7)
Thus, in any consistent labeling of $V[u]$ suppose $S_{u,v_{i}}\subset V[u]$ are the $p-1$ vertices labeled by $\{\alpha\bar{u}+\bar{v}_{i}\mid\alpha\neq 0\}$ for $i\in\{1,2\}$. Then Equation 7 implies that in the subgraph $X[u]$ the bipartite graph induced by the subsets $S_{u,v_{1}}$ and $S_{u,v_{2}}$ is either complete or empty.
Furthermore, $\langle\alpha\bar{u}+\bar{v},\beta\bar{u}+\bar{v}\rangle=0$ for all non-zero $\alpha,\beta$. Hence each $S_{u,v},v\in V^{\prime}$ induces a $(p-1)$-clique.
How do we identify each $S_{u,v}$? Suppose $v_{1},v_{2}\in V^{\prime}$ are distinct vertices. In $V^{\prime}$ (which is already labeled by $V(2r-2,p)$) we can find a vertex $v_{3}$ such that $\langle v_{1},v_{3}\rangle\neq 0$ and $\langle v_{2},v_{3}\rangle=0$. Hence each $S_{u,v}$ can be identified because
  • (a)
    All $p-1$ vertices in $S_{u,v}$ are adjacent to each other. They all have identical neighborhoods in $X[u]$ by the above argument.
  • (b)
    Every vertex in $V[u]\setminus S_{u,v}$ has a different neighborhood in $X[u]$ than the vertices in $S_{u,v}$.
Putting it together, in summary, we have shown the following.
Claim 7.11
In polynomial time we can uniquely identify all the subsets $S_{u,v},v\in V^{\prime}$, each of size $p-1$, forming a partition of $V[u]$, for each $u\in V_{12}$.
Thus, we have identified the subsets $S_{u,v}\subset V[u]$ for $v\in V^{\prime}$. We now need to label the elements of $S_{u,v}$ by vectors $\{\alpha\bar{u}+\bar{v}\mid\alpha\in\mathbb{F}^{*}_{p}\}$.
Claim 7.12
Assuming $X$ is a symplectic graph, for each $v\in V^{\prime}$ and $u\in V_{12}$, we can label the $p-1$ vertices in $S_{u,v}$ by the $p-1$ vectors $\alpha\bar{u}+\bar{v},\alpha\in\mathbb{F}^{*}_{p}$ arbitrarily to get a labeling for $V\setminus V^{\prime}$ that is consistent with the labeling of $V^{\prime}$.
Proof of Claim. We first observe that adjacencies within $V[u]$ and between $V[u]$ and $V^{\prime}$ are consistent with such a labeling by the arguments preceding the claim.
We now examine the adjacencies across two disjoint $V[u]$ and $V[u^{\prime}]$. Consider $S_{u,v}\subset V[u]$ and $S_{u^{\prime},v^{\prime}}\subset V[u^{\prime}]$. There is an edge between these subsets if and only if for some non-zero scalars $\alpha$ and $\beta$ we have $\langle\alpha\bar{u}+\bar{v},\beta\bar{u^{\prime}}+\bar{v^{\prime}}\rangle=0$. Now, $\langle\bar{u},\bar{u^{\prime}}\rangle\neq 0$. Hence we must have
$\alpha\beta=\frac{-\langle\bar{v},\bar{v^{\prime}}\rangle}{\langle\bar{u},\bar{u^{\prime}}\rangle}.$
Since $\alpha$ and $\beta$ are non-zero scalars, it follows that $\langle\bar{v},\bar{v^{\prime}}\rangle\neq 0$. It follows that $v$ and $v^{\prime}$ being non-adjacent in the graph induced by $V^{\prime}$ is a necessary condition for $S_{v}$ and $S_{v^{\prime}}$ to have any edges between them. Furthermore, the above equation implies that for the vertex, say $v_{\alpha}\in S_{u,v}$ labeled $\alpha\bar{u}+\bar{v}$ there is a unique $\beta\neq 0$ such that $v_{\alpha}$ is adjacent to the vertex, say $v^{\prime}_{\beta}$ in $S_{u^{\prime},v^{\prime}}$ labeled $\beta\bar{u^{\prime}}+\bar{v^{\prime}}$. Hence, either there are no edges between $S_{v}$ and $S_{v^{\prime}}$ (if $v$ and $v^{\prime}$ are adjacent) or there is a perfect matching between $S_{v}$ and $S_{v^{\prime}}$( if $v$ and $v^{\prime}$ are non-adjacent).
Since these are all the possible edges in $V\setminus V^{\prime}$ and it is clearly consistent with any labeling of $S_{u,v}$ by the $p-1$ vectors $\alpha\bar{u}+\bar{v},\alpha\neq 0$ for all $u\in V_{12}$ and $v\in V^{\prime}$, this completes the proof of the claim.
Putting it together, clearly if $X$ is the symplectic graph of some non-degenerate symplectic form on $V(2r,p)$ the above algorithm verifies that by constructing a symplectic form consistent with the symplectic graph. The algorithm is described as a recursive procedure below. Essentially, given a graph $X=(V,E)$ with $|V|=p^{2r}$ vertices as input it will try to compute a labeling of $V$ by the $p^{2r}$ vectors in $V(2r,p)=\{\sum_{i=1}^{r}(\alpha_{i}\bar{x}_{i}+\beta_{i}\bar{y}_{i})\mid\alpha_{i},\beta_{i}\in\mathbb{F}_{p}\}$, such that the edges of $X$ are consistent with some symplectic form on $V(2r,p)$. If $X$ is a symplectic graph it will succeed. Otherwise, the procedure will fail and detect that $X$ is not symplectic.
Procedure Label$(X,2r,p)$
  1. The procedure takes as input a simple graph $X=(V,E)$ with $p^{2r}$ vertices for an odd prime $p$ and computes a labeling consistent with some symplectic form on $V(2r,p)$ if $X$ is a symplectic graph. Otherwise, it returns “fail”.
  2. (a)
    Identify the unique dominant vertex $v_{0}\in V$ and label it by 00 (the zero vector).
  3. (b)
    Pick any pair of nonadjacent vertices $x_{1},y_{1}\in V\setminus\{v_{0}\}$ and label them by the basis vectors $\bar{x}_{1}$ and $\bar{y}_{1}$ respectively.
  4. (c)
    Let $V^{\prime}$ be the set of all vertices adjacent to both $x_{1}$ and $y_{1}$. Check $|V^{\prime}|=p^{2r-2}$. Let $X^{\prime}$ be the subgraph induced by $V^{\prime}$. Recursively call Label$(X^{\prime},2r-2,p)$ and obtain a labeling of $V^{\prime}$ by vectors in $V(2r-2,p)$ defined by the basis $\bar{x}_{i},\bar{y}_{i},2\leq i\leq p$ and consistent with a symplectic form on $V(2r-2,p)$ (Claim 7.7).
  5. (d)
    Determine the $p^{2}-1$ vertices $V_{12}$ (Claim 7.8) consisting of vertices in $V\setminus V^{\prime}$ that are adjacent to all vertices in $V^{\prime}$. Determine the partition of $V_{12}$ into $p+1$ many $(p-1)$-cliques in the induced subgraph $G_{12}$ (Claim 7.9).
  6. (e)
    From the adjacencies between $V\setminus V^{\prime}$ and $V^{\prime}$ determine the partition of $V\setminus V^{\prime}$ into $p+1$ subsets $V[u]$ (Claim 7.10).
  7. (f)
    From the subgraphs $G[u]$ induced by the respective $V[u]$, determine the partition of $V[u]$ into $p-1$-subsets $S_{u,v},v\in V^{\prime}$ (Claim 7.11) .
  8. (g)
    Label the vertices in $S_{u,v},v\in V^{\prime}$ and $u\in V_{12}$ by the vectors $\alpha\bar{u}+\bar{v},\alpha\neq 0$ (Claim 7.12).
  9. (h)
    If any step above does not satisfy the stated property then output fail. Else output pass along with the complete labeling of $V$ by elements of $V(2r,p)$ as the witness that $X$ is a symplectic graph.
Running time analysis  Let $T(n)$ denote the running time for constructing a symplectic form $\langle$,$\rangle$ , ⟩ for $V(2r,p)$ consistent with $X$ as the symplectic graph, where $n=p^{2r}$ is the number of vertices in $X$. The inductive construction and the rest of the computation implies the recurrence $T(n)=O(|V|^{2})+T(n-1)$, which gives an overall cubic bound $T(n)=O(|V|^{3})=O(p^{6r+3})$ on the running time.
Putting it together, we have a polynomial-time algorithm that checks if $X$ is the symplectic graph for the symplectic space by finding out a labeling of vertices by the vectors along with a consistent symplectic form. $\qed$
In summary, we have the following.
Theorem 7.13
Given a graph $X=(V,E)$ with $p^{2r+1}$ vertices we can determine in polynomial time if it is the commuting graph of an extraspecial group of order $p^{2r+1}$ and, if so, label vertices by unique group elements satisfying the commuting relation.
Proof 13
By Lemma 7.3 we can obtain the graph $X_{s}=(V_{s},E_{s})$ from $X$ in polynomial time. By Theorem 7.6 we can check if $X_{s}$ is the symplectic graph of a symplectic form $\langle$,$\rangle$ , ⟩ on $V(2n,p)$ and also find the symplectic form.
From the proof of Lemma 7.3, we have the following observations.
  • (a)
    The vertex of degree $|V_{s}|-1$ in $X_{s}$ is the 00 element of $V(2n,p)$, and corresponds to $Z$ in $G$.
  • (b)
    For each vertex $x\in V_{s}$ there is a closed-twins equivalence class of size $p-1$ containing $x$. There is a corresponding subset of vertices $H_{x}\setminus Z$ in $X$ of size $p^{2}-p$, where $H_{x}$, in turn, corresponds to a subgroup of $G$ of order $p^{2}$ that contains $Z$. We have a labeling of the vertex $x$ by a linear combination $\sum_{i=1}^{n}(\alpha_{i}\bar{x}_{i}+\beta_{i}\bar{y}_{i})$, and the remaining $p-2$ vertices in the equivalence class are labeled by non-zero scalar multiples of this linear combination.
Now, using the description of the extraspecial group $G_{1}$ we can label the vertices of the clique $H_{x}\setminus Z$ with the group elements
$z^{j}\cdot\left(\prod_{i=1}^{n}x_{i}^{\alpha_{i}}.y_{i}^{\beta_{i}}\right)^{k},0\leq j\leq p-1,1\leq k\leq p-1.$
This will give us the labeling of $X$ with the elements of the extraspecial group $G_{1}$ of order $p^{2r+1}$, consistent with the commuting graph $X$.
Observe that if $X_{s}$ is not a symplectic graph then Procedure Label described in Theorem 7.6 will detect that and return fail. This completes the proof. $\qed$
8 A quasipolynomial time algorithm in the general case
Assuming CFSG (the classification of finite simple groups [41]), in this section we present a $2^{O(\log^{3}n)}$ time algorithm for checking if a given $n$-vertex graph is the commuting graph of some $n$-element group and, if so, labeling the vertices of the graph by the group elements consistent with all the commuting pairs. For CFSG in this section we will refer to the (concrete) descriptions of the finite simple groups in Wilson’s text [79].
  1. (a)
    The main idea is the notion of description $\mathop{\mathrm{desc}}(G)$ for any finite group $G$. It is essentially an encoding of the group $G$ as a binary string of length at most $c\cdot\log^{3}|G|$, for some fixed constant $c>0$, from which $G$ can be recovered efficiently. More precisely, we will describe an algorithm $\mathcal{D}$ that takes $\mathop{\mathrm{desc}}(G)$ as input and computes the multiplication table of $G$ as output in time polynomial in $|G|$.
  2. (b)
    This algorithm $\mathcal{D}$ can be used to generate the multiplication tables of all groups $G$ of order $n$ as follows. We enumerate all possible binary strings $w$ of length at most $c\cdot\log^{3}n$. For each $w$ we run the algorithm $\mathcal{D}$ on it. If $w=\mathop{\mathrm{desc}}(G)$ for some $G$ then $\mathcal{D}$ will output its multiplication table. We compute $\Gamma(G)$ from $G$ and check if $X$ is isomorphic to $\Gamma(G)$ in time $2^{O(\log^{3}n)}$ using Babai’s graph isomorphism algorithm [4].
Thus, the main task of this section is to define $\mathop{\mathrm{desc}}(G)$ and describe the algorithm $\mathcal{D}$. We will first define $\mathop{\mathrm{desc}}(G)$ for finite simple groups, along with how the algorithm $\mathcal{D}$ will compute the multiplication table for $G$ given $\mathop{\mathrm{desc}}(G)$ as input. We will then extend the notion of $\mathop{\mathrm{desc}}(G)$ to all groups by using a composition series for $G$ combined with $\mathop{\mathrm{desc}}(N)$ for each composition factor $N$ of $G$.
Remark 5
We emphasize here that our notion of “short” description is entirely different from the well-studied notion of generator-relator presentations of groups (see e.g. [5]). Indeed, $\mathop{\mathrm{desc}}(G)$ and its construction will be significantly simpler. It suffices for the purpose of recovering the multiplication table of the group $G$ in time polynomial in $|G|$.
8.1 Short descriptions for finite simple groups assuming CFSG
We first briefly summarize the classification theorem of finite simple groups (CFSG) [41].
  • (a)
    Cyclic groups of prime order.
  • (b)
    Alternating groups $A_{n},n\geq 5$.
  • (c)
    Finite simple groups of Lie type.
  • (d)
    Sporadic finite simple groups.
As there are only finite many sporadic finite simple groups, for the algorithm $\mathcal{D}$ they can be trivially dealt with. The cyclic groups of prime order are also easy to deal with for the algorithm $\mathcal{D}$ as there is a unique such subgroup of order $p$ for each prime $p$.
Lemma 8.1
If $G$ is cyclic of prime order $p$ then $\mathop{\mathrm{desc}}(G)$ can be defined as the binary encoding $\mathop{\mathrm{bin}}(p)$ of $p$ which is of length $\log p$ bits. We can include $1$ as the generator with the additive group $\mathbb{Z}_{p}$ as the group it encodes.
Proof 14
The algorithm $\mathcal{D}$ will take the binary encoding $\mathop{\mathrm{bin}}(p)$ of $p$, verify that $p$ is prime using the polynomial-time primality test, and then write out the group table for the additive group of integers modulo $p$. $\qed$
For several of the remaining cases, the following simple lemma gives a framework for defining $\mathop{\mathrm{desc}}(G)$ and computing $G$’s multiplication table from it. We use the well-known fact that any finite simple group can be generated with two generators (see [57]).
Lemma 8.2
Let $H$ be a finite group and $N\lhd K\leq H$ be subgroups such that
  • (a)
    $\mathop{\mathrm{desc}}(H)$ is of length $m$.
  • (b)
    The multiplication table for $H$ can be computed from $\mathop{\mathrm{desc}}(H)$ in time polynomial in $|H|$.
  • (c)
    We can test membership of elements in $N$ in time polynomial in $|H|$.
Suppose $G=K/N$ is simple. Then $\mathop{\mathrm{desc}}(G)$ has size $O(m+\log|H|)$ and in time polynomial in $|H|$ we can compute the multiplication table for $G$.
Proof 15
As $G=K/N$ is simple it can be generated by two elements $Ng_{1},Ng_{2}\in K/N$ where $g_{1},g_{2}\in K$. Let $\mathop{\mathrm{desc}}(G)$ be the indices of these two elements in $H$ along with $\mathop{\mathrm{desc}}(H)$. By indices here we mean the positions of these elements in the multiplication table for $H$. Clearly, $\mathop{\mathrm{desc}}(G)$ is of size $O(\log|H|)$. The claimed algorithm $\mathcal{D}$ will first use $\mathop{\mathrm{desc}}(H)$ to compute the multiplication table for $H$. It can enumerate $H$ and use the indices to identify $g_{1}$ and $g_{2}$. Using the multiplication table for $H$ it can generate the subgroup $K=\langle g_{1},g_{2}\rangle$, listing down all its elements in time polynomial in $|H|$. Next, the algorithm can identify the normal subgroup $N$ (as given in the hypothesis) in time polynomial in $|H|$ and obtain the multiplication table for $G=K/N$$\qed$
We will apply the above lemma to the remaining cases of finite simple groups.
Lemma 8.3
If $G=A_{n}$ for some $n\geq 5$ then we can define $\mathop{\mathrm{desc}}(G)$ of size $O(\log|G|)$.
Proof 16
We can define $\mathop{\mathrm{desc}}(A_{n})$ as the encoding of two generators of $A_{n}$, by giving the index of the two generating permutations, say $g_{1},g_{2}$ in $S_{n}$ in the natural order of enumerating $S_{n}$. Hence, $\mathop{\mathrm{desc}}(G)$ will be of size $2\log n!=O(\log|A_{n}|)$. We can now apply Lemma 8.2 with $H=S_{n}$, $K=A_{n}$ and $N=\{1\}$ to complete the proof.$\qed$
We now consider the finite simple groups of Lie type. It is convenient to deal with them in two categories: classical and exceptional. We will follow the (concrete) constructions of these groups as explained in Wilson’s text [79].
8.1.1 Classical Lie type
Lemma 8.4
For each finite simple group $G$ that is of classical Lie type, there is a description $\mathop{\mathrm{desc}}(G)$ of $O(\log|G|)$ size such that the multiplication table for $G$ can be computed from $\mathop{\mathrm{desc}}(G)$ in time polynomial in $|G|$.
Proof 17
These finite simple groups are categorized below, following Wilson’s text [79, Chapter 3].
  1. (a)
    $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$ for either $n>2$ or $q>3$. These are the projective special linear group arising from the matrix group $\mathop{\mathrm{GL}}\nolimits_{n}(q)$ (the finite group of $n\times n$ invertible matrices over the finite field $\mathbb{F}_{q}$).
  2. (b)
    $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$ which is the projective symplectic group, obtained as quotient of the symplectic group $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$, and where $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ itself is a subgroup of $\mathop{\mathrm{GL}}\nolimits_{2m}(q)$.
  3. (c)
    $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ which is the projective unitary group, obtained as a quotient of the special unitary group $\mathop{\mathrm{SU}}\nolimits_{n}(q)$. The general unitary group $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ itself is a subgroup of the matrix group $\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})$.
  4. (d)
    Finally, the orthogonal groups which are also obtained essentially as quotients of matrix groups, but there are different subcases depending on both $n$ and $q$.
Consider $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$. The special linear group $\mathop{\mathrm{SL}}\nolimits_{n}(q)=\{A\in\mathop{\mathrm{GL}}\nolimits_{n}(q)\mid\det(A)=1\}$ is a subgroup of $\mathop{\mathrm{GL}}\nolimits_{n}(q)$ of index $q-1$. $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$ is obtained as the quotient $\mathop{\mathrm{SL}}\nolimits_{n}(q)/Z$ where $Z$ consists of all “scalar matrices” $\lambda I_{n}$ of determinant $1$. Thus, $|\mathop{\mathrm{PSL}}\nolimits_{n}(q)|\geq|\mathop{\mathrm{GL}}\nolimits_{n}(q)|/q^{2}$. We can apply Lemma 8.2, letting $H=\mathop{\mathrm{GL}}\nolimits_{n}(q)$, $K=\mathop{\mathrm{SL}}\nolimits_{n}(q)$ and $N=Z$ in it. Notice that $\mathop{\mathrm{GL}}\nolimits_{n}(q)$ has a description of size $O(\log n+\log q)$. Given $n$ and $q$ we need to enumerate $\mathop{\mathrm{GL}}\nolimits_{n}(q)$ in time polynomial in $|\mathop{\mathrm{GL}}\nolimits_{n}(q)|$. The algorithm will enumerate all $n\times n$ matrices $M$ over $\mathbb{F}_{q}$, one by one, run a polynomial time algorithm for checking if $M$ is invertible (say, using Gaussian elimination with $O(n^{3})$ many field operations) and if so include it in $\mathop{\mathrm{GL}}\nolimits_{n}(q)$. Since a large fraction of matrices are invertible (for $n$ more than some constant), this procedure will run in time polynomial in $|GL_{n}(q)|$. Thus, by Lemma 8.2 we can obtain the desired description for $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$.
Next, consider $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$. Recall that the symplectic group $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ is the subgroup of $\mathop{\mathrm{GL}}\nolimits_{2m}(q)$ consisting of all isometries of a non-singular alternating bilinear form on the vector space $\mathbb{F}_{q}^{2m}$. It turns out that $|\mathop{\mathrm{Sp}}\nolimits_{2m}(q)|>q^{m^{2}}$. As $|\mathop{\mathrm{GL}}\nolimits_{2m}(q)|=q^{4m^{2}}$, clearly
$|\mathop{\mathrm{GL}}\nolimits_{n}(q)|\leq|\mathop{\mathrm{Sp}}\nolimits_{2m}(q)|^{4}.$
Thus, the description of the group $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ is easy to construct. Given $2m$ and $q$ in binary, in time polynomial in $m$ and $q$ we can define a non-singular alternating bilinear form $\langle$,$\rangle$ , ⟩ on $\mathbb{F}_{q}^{2m}$. Then, as done for the previous case, we can enumerate the elements of $\mathop{\mathrm{GL}}\nolimits_{2m}(q)$, one by one, and put each into $\mathop{\mathrm{Sp}}\nolimits_{2m}(q)$ if it is an isometry for $\langle$,$\rangle$ , ⟩. Now, the simple group $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$ by quotienting out the subgroup of scalar matrices (which has size at most $2$). Clearly, Lemma 8.2 can be applied to obtain the desired description for $\mathop{\mathrm{PSp}}\nolimits_{2m}(q)$ along with the algorithm for obtaining its multiplication table.
For the next case $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$, we start with the group $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ which is the subgroup of $\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})$ that preserves the sesquilinear form $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}^{q}$. From there on the process of obtaining $\mathop{\mathrm{SU}}\nolimits_{n}(q)$ and then $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ is exactly like in the case of $\mathop{\mathrm{PSL}}\nolimits_{n}(q)$. In order to obtain the desired description for $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ it suffices to generate $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ in time polynomial in $|\mathop{\mathrm{PSU}}\nolimits_{n}(q)|$ given $n$ and $q$. Clearly, $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ can be generated in time polynomial in $|\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})|$ by looking all these matrices. However, $|\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})|<|\mathop{\mathrm{GU}}\nolimits_{n}(q)|^{3}$. Thus, the brute-force algorithm for generating all of $\mathop{\mathrm{GL}}\nolimits_{n}(q^{2})$ by looking through all $n\times n$ invertible matrices over $\mathbb{F}_{q}$ and picking those which preserve the sesquilinear form will again work. Finally, the group $\mathop{\mathrm{PSU}}\nolimits_{n}(q)$ is only smaller than $\mathop{\mathrm{GU}}\nolimits_{n}(q)$ by a small power of $q$.
The fourth category are the simple groups arising from the orthogonal groups. There are different cases to consider here.
When $q$ is odd, there are exactly two non-singular $n$-variate symmetric bilinear forms over $\mathbb{F}_{q}$. If $f$ is one of them then $\mathop{\mathrm{GO}}\nolimits_{n}(f)$ is the subgroup of $GL_{n}(q)$ that leaves $f$ invariant. If $n=2m+1$ is odd, it turns out that there is only one orthogonal group which we may denote by $\mathop{\mathrm{GO}}\nolimits_{2m+1}(q)$. We can then construct $\mathop{\mathrm{SO}}\nolimits_{2m+1}(q)$ and $\mathop{\mathrm{PSO}}\nolimits_{2m+1}(q)$ analogous to the earlier cases. But there is a further subgroup of index $2$ in $\mathop{\mathrm{PSO}}\nolimits_{2m+1}(q)$ denoted $\mathop{\mathrm{P}\Omega}\nolimits_{2m+1}(q)$. As before, we can apply Lemma 8.2 to obtain the desired description of $\mathop{\mathrm{P}\Omega}\nolimits_{2m+1}(q)$ from which the group’s multiplication table can be obtained in polynomial time.
For $n=2m$, there are two orthogonal groups $\mathop{\mathrm{GO}}\nolimits^{+}_{2m}(q)$ and $\mathop{\mathrm{GO}}\nolimits^{-}_{2m}(q)$. As in the previous case we can obtain the groups $\mathop{\mathrm{P}\Omega}\nolimits^{\epsilon}_{2m}(q)$ for $\epsilon\in\{+,-1\}$. For odd $q$ these are always simple and Lemma 8.2 is again easily applicable.
The characteristic $2$ case of orthogonal groups requires different techniques. However, Lemma 8.2 is still applicable in defining $\mathop{\mathrm{desc}}(G)$ for the simple groups $G$ arising in this case. We omit the details.$\qed$
8.1.2 Exceptional Lie type
We will closely follow Wilson’s exposition [79, Chapter 4]. These finite simple groups come in ten families:
  • (a)
    The five families of Chevalley groups $G_{2}(q)$, $F_{4}(q)$, and $E_{n}(q),n\in\{6,7,8\}$.
  • (b)
    The two families of Steinberg–Tits–Herzig groups ${}^{3}{D}_{4}(q)$ and ${}^{2}{E}_{6}(q)$.
  • (c)
    The Suzuki and Ree groups $Sz(2^{2n+1})$, $R_{1}(3^{2n+1})$ and $R_{2}(2^{2n+1})$ (also known as ${}^{2}B_{2}(q)$, ${}^{2}G_{2}(q)$ and ${}^{2}F_{4}(q)$ for the appropriate $q$).
Unlike the classical case, these families have only one parameter which is the field size $q$. It turns out that the following lemma (analogous to Lemma 8.2) captures how these simple groups arise and suggest a suitable $\mathop{\mathrm{desc}}(G)$ for each $G$ along with the corresponding algorithm for recovering $G$ from $\mathop{\mathrm{desc}}(G)$.
Lemma 8.5
Let $G$ be finite simple group from any of the ten families listed above. Then we can define a description $\mathop{\mathrm{desc}}(G)$ of size $O(\log q)$ such that $G$’s multiplication table can be obtained from it in time polynomial in $|G|$.
Proof 18
We will briefly describe how the groups are constructed for each of these ten families. The easy descriptions and algorithms for generating their multiplication tables will suggest themselves.
  1. (a)
    $\mathop{\mathrm{Sz}}\nolimits(2^{2n+1})$:   Let $q=2^{2n+1}$. The Suzuki group $Sz(q)$ (see [79, Chapter 4.2]) is a subgroup of the symplectic group $\mathop{\mathrm{Sp}}\nolimits_{4}(2^{2n+1})$. It arises as the centralizer of a particular automorphism of order $2$ in $\mathop{\mathrm{Sp}}\nolimits_{4}(2^{2n+1})$. As $|\mathop{\mathrm{Sz}}\nolimits(q)|=(q^{2}+1)q^{2}(q-1)$ $|\mathop{\mathrm{GL}}\nolimits_{4}(q)|\leq|\mathop{\mathrm{Sz}}\nolimits(q)|^{4}$. Thus, we can cycle through all elements of $\mathop{\mathrm{Sp}}\nolimits_{4}(q)$ looking for the elements of this centralizer. That automorphism of order $2$ in $\mathop{\mathrm{Sp}}\nolimits_{4}(2^{2n+1})$ is also easily described (in size independent of $q$).
  2. (b)
    $G_{2}(q)$:  It turns out that $G_{2}(q)$ (see [79, Chapter 4.3]) arises as the automorphism group of the octonion algebra over $\mathbb{F}_{q}$. The entire octonion algebra can be listed in time $O(q^{8})$. As the automorphisms permute fix the identity in the octonion algebra, $G_{2}(q)$ is essentially a subgroup of $\mathop{\mathrm{GO}}\nolimits_{7}(q)$. We can enumerate $\mathop{\mathrm{GO}}\nolimits_{7}(q)$ and pick out the automorphisms which will give us $G_{2}(q)$. As $|G_{2}(q)|$ is $q^{6}(q^{6}-1)(q^{2}-1)$, this will be polynomial time in the group size. We can define the $\mathop{\mathrm{desc}}(G_{2}(q))$ of size $O(\log q)$. It is essentially just the encoding of $q$.
  3. (c)
    $R_{1}(3^{2n+1})={}^{2}G_{2}(3^{2n+1})$: This can be handled quite analogously to $G_{2}(q)$ [79, Section 4.5].
  4. (d)
    ${}^{3}{D}_{4}(q)$:  This group (see [79, Chapter 4.6]) arises as centralizer of a certain order $3$ automorphism of $\mathop{\mathrm{P}\Omega}\nolimits^{+}_{8}(q^{3})$. Because of the special nature of this automorphism (product of an $\mathbb{F}_{q}$ automorphism and the triality automorphism), it has a short description of size $O(\log q)$. As $\mathop{\mathrm{P}\Omega}\nolimits^{+}_{8}(q^{3})$ can be enumerated in time polynomial in $q$, we can pick out the centralizer elements in time polynomial in $q$.
  5. (e)
    $F_{4}(q)$:  when $q$ is not a power of $2$ or $3$, $F_{4}(q)$ arises as automorphisms of the finite (non-associative) Albert algebra which is $27$-dimensional [79, Chapter 4.8]. Thus, $F_{4}(q)$ can be picked out from $GL_{27}(q)$ by enumerating it and checking each element for being an automorphism. The finite Albert algebra can be defined in $O(\log q)$ size by defining the multiplication of its basis elements.
  6. (f)
    $R_{2}(2^{2n+1})={}^{2}F_{4}(2^{2n+1})$:  This case is quite similar to the previous one [79, Section 4.9].
  7. (g)
    $E_{6}(q)$:  This group can be defined as automorphisms of a certain trilinear form over the Albert algebra [79, Chapter 4.10]. After quotienting out the scalar subgroup from this automorphism group we will obtain $E_{6}(q)$. The same algorithm of enumerating the bigger group and selecting the automorphisms and then quotienting out the scalars will work.
  8. (h)
    $E_{7}(q)$ and $E_{8}(q)$:  Both these arise as automorphisms of a finite Lie algebra [79, Chapter 4.12]. For $E_{8}(q)$ we have to consider a $248$ dimensional Lie algebra. Thus we can look for $E_{8}(q)$ inside $GL_{248}(q)$ by enumerating all its elements and picking out the automorphisms. All we need is that from a $O(\log q)$ size description we should be able to obtain the Lie algebra. For that we only need to express the multiplication of basis elements as an $\mathbb{F}_{q}$ linear combination of the basis (which is of size $248$). $E_{7}(q)$ is similarly dealt with as automorphisms of a smaller dimensional Lie algebra.$\qed$
The above lemmas immediately imply the following.
Theorem 8.6
There is an algorithm that, given as input $\mathop{\mathrm{desc}}(G)$ for a finite simple group $G$, outputs its multiplication table in time polynomial in $|G|$.
8.2 Short descriptions for finite groups
We first show that every finite group $G$ has a description $\mathop{\mathrm{desc}}(G)$ of bit-length $O(\log^{3}|G|)$. Then we describe the algorithm that given as input a bit string $y$, if $y=\mathop{\mathrm{desc}}(G)$ for some finite group $G$ then the algorithm runs in time $2^{O(\log^{3}|G|)}$ and outputs the multiplication table of a group $G^{\prime}$ isomorphic to $G$.
Building on the construction of $\mathop{\mathrm{desc}}(G)$ for finite simple groups $G$, described in Section 8.1, we obtain $\mathop{\mathrm{desc}}(G)$ for all finite groups with a straightforward construction that uses a composition series for $G$. It is convenient to use Erdős-Rényi cube generating sequences [33].
Definition 4 (Erdős-Rényi cube generating sequences)
Let $G$ be a finite group. A cube generating sequence (abbreviated as c.g.s.) for $G$ is a sequence $(g_{1},g_{2},\ldots,g_{k})$ of elements from $G$ such that
$G=\{g_{1}^{\epsilon_{1}}g_{2}^{\epsilon_{2}}\cdots g_{k}^{\epsilon_{k}}\mid\epsilon_{i}\in\{0,1\}\hbox{ for each }i\}.$
Thus, given the cube generating sequence $(g_{1},g_{2},\ldots,g_{k})$, each element of $G$ can be defined by the bit vector $(\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{k})$. We note that this bit vector need not be unique for a group element.
Theorem 8.7
[33] Every finite group $G$ has a cube generating sequence of size at most $2\log|G|$.
Let $X=(g_{1},g_{2},\ldots,g_{k})$ be cube generating sequence for a group $G$. Its length $k$ is denoted $|X|$. Abusing notation, by $g\in X$ we will mean that $g=g_{i}$ for some $i$.
We note that cube generating sequences can be computed in polynomial time.
Lemma 8.8
There is a polynomial-time algorithm that takes a group $G$ by its multiplication table as input and outputs a cube-generating sequence of length at most $2\log|G|$ for $G$.
Proof 19
The simple algorithm is directly based on a “cube doubling” trick due to Babai and Szemerédi [6]. Suppose we have picked the sequence $(g_{1},g_{2},\ldots,g_{i})$ of group elements from $G$. Let $C_{i}$ be the cube it generates. That is,
$C_{i}=\{g_{1}^{\epsilon_{1}}g_{2}^{\epsilon_{2}}\cdots g_{i}^{\epsilon_{i}}\mid\epsilon_{j}\in\{0,1\}\hbox{ for }1\leq j\leq i\}.$
For any $g\in G$ let $C_{i}g=\{xg\mid x\in C_{i}\}$. As $|C_{i}|\leq 2^{i}$ and each element in $C_{i}$ can be computed with $i-1$ group multiplications, the set $C_{i}$ can be enumerated with at most $(i-1)2^{i}$ many group multiplications. We can enumerate $C_{i}g$ with $2^{i}$ additional group multiplications.
Now, if $C_{i}g\cap C_{i}\neq\emptyset$ for $g\in G$ then $g$ is in the cube generated by the sequence $(g_{i}^{-1},g_{i-1}^{-1},\ldots,g_{1}^{-1},g_{1},\ldots,g_{i})$. Thus, if for all $g\in G$ $C_{i}g\cap C_{i}\neq\emptyset$ then we have a cube generating sequence of length $2i$. Otherwise, there is a $g_{i+1}\in G$ such that
$C_{i}g_{i+1}\cap C_{i}=\emptyset,$
and the cube $C_{i+1}$ generated by $(g_{1},g_{2},\ldots,g_{i+1})$ is twice the cardinality of $C_{i}$ by a simple counting argument. Clearly, we can find $g_{i+1}$ in time polynomial in $|G|$ by searching through $G$.
As $|C_{i+1}|=2\cdot|C_{i}|$, it follows that this procedure will eventually find a cube generating sequence $C_{k},k\leq 2\log|G|$ with at most polynomial in $|G|$ group multiplications and hence in time polynomial in $|G|$. $\qed$
We now explain the construction of $\mathop{\mathrm{desc}}(G)$ for a finite group $G$, where $G$ is given by its multiplication table. Compute a composition series for $G$ (which can be done in time polynomial in $|G|$ by computing normal closures of elements):
$1=N_{0}\lhd N_{1}\lhd\cdots\lhd N_{r}=G.$
For each quotient group $H_{i}=N_{i}/N_{i-1}$ for $1\leq i\leq r$, we can compute a cube generating sequence $Y_{i}=(N_{i-1}g_{i,1},N_{i-1}g_{i,2},\ldots,N_{i-1}g_{i,k_{i}})$ of length $k_{i}\leq 2\log|H_{i}|$ in polynomial time, by Lemma 8.8. For each $i$, let $X_{i}=(g_{i,1},g_{i,2},\ldots,g_{i,k_{i}})$ be the corresponding sequence of coset representatives.
Proposition 8.9
The sequence $X=(X_{1},X_{2},\cdots,X_{i})$ obtained by concatenating the sequences $X_{j},1\leq j\leq i$, defined above, in that order is a cube generating sequence for the subgroup $N_{i}$ of length at most $\sum_{j=1}^{i}2\log|H_{j}|=2\log|N_{i}|$, for $1\leq i\leq r$. In particular, $X=(X_{1},X_{2},\cdots,X_{r})$ is a cube generating sequence for $G$ of length at most $2\log|G|$.
Proof 20
Suppose $H\lhd G$ and $Y=\{y_{1},y_{2},\ldots,y_{\ell}\}$ is a cube generating sequence for $H$ and $Z=\{Hz_{1},Hz_{2},\ldots,Hz_{m}\}$ is a cube generating sequence for $G/H$. We note that any element $Hg\in G/H$ can be expressed as
$Hg=Hz_{1}^{\epsilon_{1}}z_{2}^{\epsilon_{2}}\cdots z_{m}^{\epsilon_{m}},\epsilon_{i}\in\{0,1\},$
which implies that any element of $G$ can be expressed as
$y_{1}^{\delta_{1}}y_{2}^{\delta_{2}}\cdots y_{\ell}^{\delta_{\ell}}z_{1}^{\epsilon_{1}}z_{2}^{\epsilon_{2}}\cdots z_{m}^{\epsilon_{m}},\,\epsilon_{i},\delta_{j}\in\{0,1\}.$
Now, building on this, an easy inductive argument yields the proposition.$\qed$
For $1\leq i\leq r$, let $Y_{i}=(X_{1},X_{2},\ldots,X_{i})$. By the above proposition, $Y_{i}$ is a cube generating sequence for the group $N_{i}$ of length at most $2\log|N_{i}|$. As $N_{i-1}\lhd N_{i}$, for each $x\in X_{i}$ we have $xN_{i-1}=N_{i-1}x$, $1<i\leq r$. In particular, for $y\in Y_{i}$ there is a unique element $w(x,y)\in N_{i-1}$ such that we have
$xy=w(x,y)x,$ (8)
where $w(x,y)$ is a word over the cube generating sequence $Y_{i-1}$. Thus, $w(x,y)$ is a word of length at most $2\log|N_{i-1}|$, expressed using the c.g.s. $Y_{i-1}$. Let $R_{i}$ consist of the set of all such equations defined by Equation (8) for $y\in Y_{i-1}$ and $x\in X_{i}$. Notice that $R_{i}$ has $|Y_{i-1}|\cdot|X_{i}|$ equations in it and each word $w(x,y)$ over $Y_{i-1}$ is defined by the bit vector of exponents in the cube product over $Y_{i-1}$. Thus, the size of $R_{i}$ is
$\mathop{\mathrm{size}}(R_{i})=O(|Y_{i-1}|\cdot|X_{i}|\cdot|Y_{i-1}|)=O((\log|N_{i-1}|)^{2}\cdot\log|H_{i}|).$
We are now ready to define $\mathop{\mathrm{desc}}(G)$ for the finite group $G$. It is defined as an efficient [^2] We can encode an $r$-tuple of $3$-tuples in binary by using some standard encoding which is polynomial time computable and its inverse is also polynomial time computable. binary encoding of the $r$-tuple defined below
$\mathop{\mathrm{desc}}(G)~{}=~{}[(\mathop{\mathrm{desc}}(H_{1}),X_{1},R_{1}),(\mathop{\mathrm{desc}}(H_{2}),X_{2},R_{2}),\ldots,(\mathop{\mathrm{desc}}(H_{r}),X_{r},R_{r})],$
where $R_{1}$ is the empty set. The size of $\mathop{\mathrm{desc}}(G)$ is clearly
$O(\sum_{i=1}^{r}\log~{}|H_{i}|)+O(\sum_{i=1}^{r}\mathop{\mathrm{size}}(X_{i}))+O(\sum_{i=1}^{r}\mathop{\mathrm{size}}(R_{i})),$
where $\mathop{\mathrm{size}}(X_{i})$ and $\mathop{\mathrm{size}}(R_{i})$ denote the bit length of their encodings described above. Notice that $\sum_{i=1}^{r}\log~{}|H_{i}|=\log|G|$. It follows that $\sum_{i=1}^{r}\mathop{\mathrm{size}}(X_{i})\leq O(\log^{2}|G|)$, because $|X_{i}|\leq 2\log|H_{i}|\leq 2\log|G|$ and each element of $X_{i}$ requires at most $\log|H_{i}|$ bits to encode it. We now bound $\sum_{i=1}^{r}\mathop{\mathrm{size}}(R_{i})$.
$\sum_{i=1}^{r}\mathop{\mathrm{size}}(R_{i})$ $=$ $O(\sum_{i=1}^{r}(\log^{2}|N_{i-1}|\cdot\log|H_{i}|)$
$\leq$ $\sum_{i=1}^{r}O(\log^{2}|G|)\cdot\log~{}|H_{i}|$
$=$ $O(\log^{2}|G|)\cdot\sum_{i=1}^{r}\log~{}|H_{i}|$
$=$ $O(\log^{3}|G|).$
The above discussion has the following immediate consequence.
Lemma 8.10
Given as input the multiplication table of a finite group $G$, there is an algorithm that computes in time polynomial in $|G|$ its description $\mathop{\mathrm{desc}}(G)$.
However, we are interested in the converse statement, about generating $G$ from its description $\mathop{\mathrm{desc}}(G)$ in time polynomial in $|G|$. The next theorem, which shows this, will allow us to enumerate the multiplication tables of all groups of order $n$ upto isomorphism in time $2^{O(\log^{3}n)}$.
Theorem 8.11
There is an algorithm that given as input $\mathop{\mathrm{desc}}(G)$, for a finite group $G$, outputs its multiplication table in time polynomial in $|G|$. As a consequence, there is an algorithm that outputs the list of multiplication tables of all distinct groups of order $n$ in time $2^{O(\log^{3}n)}$.
Proof 21
As argued above, every finite group $G$ has a description $\mathop{\mathrm{desc}}(G)=[(\mathop{\mathrm{desc}}(H_{1}),X_{1},R_{1}),(\mathop{\mathrm{desc}}(H_{2}),X_{2},R_{2}),\ldots,(\mathop{\mathrm{desc}}(H_{r}),X_{r},R_{r})]$ where the $r$-tuple is encoded in a binary string of length bounded by $c\cdot\log^{3}|G|$ for some constant $c>0$ (which is independent of $G$).
Let $w$ be a binary string. If it encodes $\mathop{\mathrm{desc}}(G)$ for some group $G$ such that $|G|=n$, then
  • (a)
    $|w|\leq c\log^{3}n$ and $w$ decodes to an $r$-tuple of $3$-tuples
    $[(\mathop{\mathrm{desc}}(H_{1}),X_{1},R_{1}),(\mathop{\mathrm{desc}}(H_{2}),X_{2},R_{2}),\ldots,(\mathop{\mathrm{desc}}(H_{r}),X_{r},R_{r})].$
  • (b)
    Moreover, for some composition series $1=N_{0}\lhd N_{1}\lhd\cdots\lhd N_{r}=G$ of $G$, $\mathop{\mathrm{desc}}(N_{i})$ is the binary encoding of the $i$-tuple
    $\mathop{\mathrm{desc}}(N_{i})=[(\mathop{\mathrm{desc}}(H_{1}),X_{1},R_{1}),(\mathop{\mathrm{desc}}(H_{2}),X_{2},R_{2}),\ldots,(\mathop{\mathrm{desc}}(H_{i}),X_{i},R_{i})],$
    for $1\leq i\leq r$, where $H_{i}=N_{i}/N_{i-1}$.
The algorithm will compute the multiplication table for $N_{i}$, inductively for increasing values of $i,1\leq i\leq r$. For $i=1$, it can be obtained from $\mathop{\mathrm{desc}}(H_{1})$ by Theorem 8.6.
Given the multiplication table for $N_{i-1}$ and the triple $(\mathop{\mathrm{desc}}(H_{i}),X_{i},R_{i})$ we first obtain $H_{i}$ by Theorem 8.6, and the actual elements of the c.g.s. $X_{i}=(x_{1},x_{2},\ldots,x_{\ell})$ for $\ell\leq 2\log|H_{i}|$. The c.g.s. for $N_{i-1}$ is $Y_{i}=(X_{1},X_{2},\ldots,X_{i-1})=(y_{1},y_{2},\ldots,y_{k})$ for some $k\leq 2\log|N_{i-1}|$. Applying Equation (8) we have as part of $R_{i}$
$x_{\mu}y_{\nu}=w(x_{\mu},y_{\nu})x_{\mu},$ (9)
for each $\mu\in[\ell]$ and each $\nu\in[k]$. Here, note that $w(x_{\mu},y_{\nu})$ is expressed as a product $y_{1}^{e_{1}}y_{2}^{e_{2}}\cdots y_{k}^{e_{k}}$ for $(e_{1},e_{2},\ldots,e_{k})\in\{0,1\}^{k}$.
We now explain how to obtain the multiplication table for the subgroup $N_{i}$ from the above data in time polynomial in $|N_{i}|$. More precisely, the data comprises of the multiplication table for $N_{i-1}$, the multiplication table for the quotient group $H_{i}=N_{i}/N_{i-1}$ and $R_{i}$ (Equation (9) for all $y_{\nu}$ and $x_{\mu}$).
We first note that $N_{i}=\cup_{x}N_{i-1}x$, where $x$ is in the cube
$C=\{\prod_{i=1}^{\ell}x_{j}^{\varepsilon_{j}}\mid\varepsilon_{j}\in\{0,1\}\textrm{ and }x_{j}\in X_{i}\}.$
Thus, $N_{i}=N_{i-1}C$. So, any element in $N_{i}$ is of the form $ab,a\in N_{i-1},b\in C$. We need to express the product of any two elements $a_{1}b_{1},a_{2}b_{2}\in N_{i}$, where $a_{1},a_{2}\in N_{i-1}$ and $b_{1},b_{2}\in C$ in the form $a_{3}b_{3},a_{3}\in N_{i-1},b_{3}\in C$.
Claim 8.12
For any $x_{\mu}\in X_{i}$ and $a\in N_{i-1}$ we can find $a^{\prime}\in N_{i-1}$ such that $x_{\mu}a=a^{\prime}x_{\mu}$ in time polynomial in $|N_{i-1}|$.
Proof of Claim 1
We first express $a$ using the c.g.s. $Y_{i}$ for $N_{i-1}$ to write it as
$a=y_{1}^{\beta_{1}}y_{2}^{\beta_{2}}\cdots y_{\ell}^{\beta_{\ell}}.$
Then we can use Equation (9) $\ell$ times to write $x_{\mu}a=x_{\mu}\prod_{j=1}^{\ell}y_{j}^{\beta_{j}}$ as the product $x_{\mu}a=(\prod_{j=1}^{\ell}w(x_{\mu},y_{j})^{\beta_{j}})x_{\mu}$. We can compute $a^{\prime}=\prod_{j=1}^{\ell}w(x_{\mu},y_{j})^{\beta_{j}}$ using the multiplication table for $N_{i-1}$ as each $w(x_{\mu},y_{j})$ is in $N_{i-1}$.
Using the above claim repeatedly we can express products of the form $xa$ for $x\in C,a\in N_{i-1}$ as $a^{\prime}x$ for some $a^{\prime}\in N_{i-1}$. Finally, we can express the product of any two elements $a_{1}b_{1},a_{2}b_{2}\in N_{i}$, where $a_{1},a_{2}\in N_{i-1}$ and $b_{1},b_{2}\in C$ in the form $a_{3}b_{3},a_{3}\in N_{i-1},b_{3}\in C$. The algorithm runs in time polynomial in $|N_{i-1}|$ and $|X_{i}|$ and hence in time polynomial in $|N_{i}|$.
Applied for increasing values of $i$, finally, for $i=r$ we obtain the table for $N_{r}=G$. The overall algorithm clearly runs in polynomial in $|G|$ time.
The algorithm for enumerating all order $n$ group multiplication tables is easy to obtain from the above. We run through all strings $w$ of length at most $c\log^{3}n$, where $n=|G|$, and apply the above algorithm to generate the multiplication for $G$ if $w$ encodes $\mathop{\mathrm{desc}}(G)$.
Notice that this enumeration may contain duplicates because the same group can have different composition series and hence different descriptions. We can eliminate the duplicates by running Miller’s group isomorphism test [59]. Each such test takes $n^{\log n}$. Since there are at most $2^{\log^{3}n}$ groups being listed the running time to detect multiplicities takes time at most $2^{\log^{3}n}\cdot n^{\log n}\cdot 2^{\log^{3}n}=2^{O(\log^{3}n)}$$\qed$
Remark 6
Some notes concerning the preceding results:
  1. (a)
    Computing a composition series for $G$ can be done in time polynomial in $|G|$ by computing normal closures of elements:
    $1=N_{0}\rhd N_{1}\rhd\cdots\rhd N_{r}=G.$
    More precisely, for each $g\in G$ we compute its normal closure by starting with the cyclic subgroup $N=\langle g\rangle$ and repeatedly replacing it with the (larger) subgroup generated by the set $\{xyx^{-1}\mid x\in G,y\in N\}$ until $N$ does not grow any further. By construction, the final $N$ is a normal subgroup of $G$ and its computation takes time polynomial in $|G|$. If $G$ is not simple, then for some $g\in G$ this construction will yield a non-trivial normal subgroup $N$. We can now recursively find composition series for $N$ and $G/N$ (we have multiplication tables for both) and put them together to obtain a composition series for $G$ in time polynomial in $|G|$.
  2. (b)
    Let $H_{i}=N_{i}/N_{i-1}$ for $1\leq i\leq r$ be the simple quotient groups. Let $n_{i}=|H_{i}|$ for $1\leq i\leq r$ denote their respective orders.
  3. (c)
    For each group $N_{i}$ we compute a cube generating sequence $X_{i}$ of size at most $k_{i}\leq 2\log|N_{i}|$. This computation is done in time polynomial in $|N_{i}|$, by Lemma 8.8. As part of the description of $G$, for each $1\leq i\leq r$ we obtain a set of relations $R_{i}$ as defined in Equation (8).
Our algorithm described in Theorem 8.11 is broadly based on a result of McIver and Neumann [58] that bounds the number of $n$ element groups by $2^{O(\log^{3}n)}$. There is related work by Babai et al. [5] which shows that all simple groups of $n$ elements (except one family called the Ree groups) have generator-relator presentations of size $O(\log^{2}n)$.
Theorem 8.13
There is a $2^{O(\log^{3}n)}$ time algorithm that recognizes if a given $n$-vertex graph is the commuting graph of a group of order $n$.
Proof 22
Let $X$ be an $n$-vertex graph. The algorithm cycles through all multiplication tables of distinct groups of order $n$, listed by the $2^{O(\log^{3}n)}$ time algorithm described in Theorem 8.11. For each group $G$ that is listed we can compute its commuting graph $\Gamma(G)$ in polynomial time. We can test if $\Gamma(G)$ is isomorphic to the input graph $X$ using Babai’s $2^{O(\log^{3}n)}$ time algorithm [4].
Thus, the overall computation takes $2^{O(\log^{3}n)}$ time.$\qed$
Remark 7
This algorithm is not very interesting because it does not reveal any new relationship between the group and its commuting graph. Can we do better?
9 Isoclinism of groups and commuting graphs
Motivated by the polynomial-time algorithm for recognizing commuting graphs of extraspecial groups, in this section we briefly discuss isoclinism of groups and its connected to commuting graphs.
The lexicographic product of graphs $\Gamma_{1}$ and $\Gamma_{2}$ is the graph on the vertex set $V(\Gamma_{1})\times V(\Gamma_{2})$, in which $(v_{1},w_{1})$ and $(v_{2},w_{2})$ are joined if and only if either $v_{1}$ and $v_{2}$ are joined in $\Gamma_{1}$, or $v_{1}=v_{2}$ and $w_{1}$ and $w_{2}$ are joined in $\Gamma_{2}$. If $\Gamma_{2}$ is a complete graph, then the lexicographic product is obtained from $\Gamma_{1}$ by blowing up each vertex to a complete graph on $|V\Gamma_{2}|$ vertices.
Commutation in a group $G$ is a map from $G\times G$ to the derived group $G^{\prime}$; multiplying the arguments by elements of $Z(G)$ does not change the function value, so commutation induces a map from $G/Z(G)\times G/Z(G)$ to $G^{\prime}$. We say that groups $G$ and $H$ are isoclinic if there are isomorphisms $\theta:G/Z(G)\to H/Z(H)$ and $\phi:G^{\prime}\to H^{\prime}$ which intertwine the commutation map; that is, for all $g_{1},g_{2}\in G$,
$[\theta(g_{1}Z(G)),\theta(g_{2}Z(G))]=\phi([g_{1}Z(G),g_{2}Z(G)]).$
Now the commuting graph of a group $G$ is the lexicographic product of a graph on $G/Z(G)$ with $g_{1}Z(G)$ and $g_{2}Z(G)$ joined if $[g_{1},g_{2}]=1$ (this graph can also be described as the induced subgraph of the commuting graph of $G$ on a transversal for $Z(G)$ in $G$) with a complete graph on $Z(G)$; the graph on $G/Z(G)$ is determined by the isoclinism type. Hence, as observed by Pezzott and Nakaoka [62], we have:
Theorem 9.1
Isoclinic groups of the same order have isomorphic commuting graphs.
Remark 8
It follows from Lemma 7.2 that the extraspecial groups $G_{1}$ and $G_{2}$, defined in Section 7, are isoclinic. Indeed, the recognition algorithm of Section 7 implicitly uses this property.
Is there a converse to Theorem 9.1? That is, if two groups have isomorphic commuting graphs, must they be isoclinic? This holds for various classes of groups, such as abelian groups, non-abelian simple groups, and extraspecial groups (as we have seen).
However, it is not true in general; we recycle an example taken from [23] to show this. Let $G$ be the group of order $64$ which is SmallGroup(64,182) in the SmallGroups library in GAP [37]. A Schur cover of $G$ is a group $H$ of maximal order subject to having a subgroup $Z\leq H^{\prime}\cap Z(H)$ such that $H/Z\cong G$. The Schur multiplier of $G$ has order $2$, so any Schur cover of $G$ has order $128$. Moreover, the Bogomolov multiplier (see [23] for context and other references) of $G$ is equal to the Schur multiplier, which implies that any Schur cover $H$ is commutation-preserving: that is, two elements $a,b\in H$ commute if and only if their projections $Za,Zb\in G$ commute. This implies that the commuting graph of a Schur cover is obtained from the commuting graph of $G$ by replacing each vertex with a clique of size $2$, with all edges between cliques corresponding to adjacent vertices. This procedure also describes the commuting graph of $G\times C_{2}$. On the other hand, it is easy to verify computationally that the derived groups of $H$ and $G\times C_{2}$ are not isomorphic, so these groups cannot be isoclinic.
We note that the group $G$ has nilpotency class $3$, as do all of its Schur covers (these are SmallGroup(128,$i$) for $i\in\{789,790,791,815,816,817\}$ in the GAP library). So the following question is still open:
Conjecture
Is it true that a nilpotent group of class $2$ is determined up to isoclinism by its commuting graph?
This is interesting because, as observed, extraspecial groups are nilpotent of class two $2$ and are determined up to isoclinism by their commuting graphs.
Part II Forbidden subgraphs
10 Preliminaries
A number of important graph classes can be defined either structurally or in terms of forbidden induced subgraphs. Here are structural descriptions.
  1. (a)
    $\Gamma$ is called a cograph if it can be built from $1$-vertex graphs by the operations of disjoint union and complementation.
  2. (b)
    $\Gamma$ is called a chordal graph provided that $\Gamma$ contains no induced cycles of length greater than $3$; in other words, any cycle of length at least $4$ contains a chord.
The characterization by forbidden subgraphs is a composite of results from [36, 27, 72, 26].
Proposition 10.1
  1. (a)
    A graph is a cograph if and only if it forbids $P_{4}$.
  2. (b)
    A graph is chordal if and only if it forbids $\{C_{n}:n\geq 4\}$.
Since a cycle of length $5$ or more contains a $4$-vertex path, we have the following observation.
Observation 3
If a graph is $\{P_{4},C_{4}\}$-free, then it is chordal.
We conclude this section with a general remark. In the commuting graph $\Gamma(G)$ of a group $G$, elements of the centre $Z(G)$ are dominating vertices; this is the reason why, to investigate questions about connectedness, they are excluded. However, in our case, it doesn’t matter:
Proposition 10.2
  1. (a)
    The class of groups whose commuting graph is $\mathcal{F}$-free is subgroup-closed.
  2. (b)
    If no graph in $\mathcal{F}$ has a dominating vertex, then the class of groups with $\mathcal{F}$-free commuting graph is independent of whether vertices in $Z(G)$ are included or excluded.
Proof 23
(a) This is immediate from the fact that, if $H$ is a subgroup of $G$, then $\Gamma(H)$ is an induced subgraph of $\Gamma(G)$. (We remark that the same result holds for various other graphs such as the power graph, enhanced power graph, nilpotency and solvability graphs. But it does not hold for yet others such as the generating graph or the deep commuting graph.)
(b) If $X$ is an induced subgraph of $\Gamma(G)$, then vertices of $X\cap Z(G)$ would be dominating vertices in $X$. $\qed$
In the cases we are interested in (chordal graphs and cographs), the excluded subgraphs are paths and cycles with more than three vertices. None of these have dominating vertices, so we can take the vertex set of the commuting graph to be all of $G$, but to decide if it has any of these properties we can delete vertices in $Z(G)$.
Also to prove main results of this paper, we need the following assertion which generalizes a result obtained by Wielandt [78] in 1939.
Proposition 10.3
Let $G$ be a finite group. Assume that one of the following statements holds:
  1. (a)
    $Z(G)=1$;
  2. (b)
    $G=G^{\prime}$.
Then $C_{\mathop{\mathrm{Aut}}(G)}(\mathop{\mathrm{Inn}}(G))=1$.
Proof 24
Let $\phi$ be an automorphism of $G$ which commutes with each inner automorphism of $G$. Then for each $g,h\in G$, we have
$\phi(h)^{g}=\phi(h^{g})=\phi(h)^{\phi(g)},$
i. e. $[g,\phi]=g^{-1}\phi(g)\in Z(G)$ and $\phi(g)=g\cdot z_{g}$ for some $z_{g}\in Z(G)$ (depending of $g$).
If $Z(G)=1$, then $z_{g}=1$ for each $g$, therefore $\phi$ is trivial and $C_{\mathop{\mathrm{Aut}}(G)}(\mathop{\mathrm{Inn}}(G))=1$, as required.
Now assume that $G$ is perfect, i.e. $G^{\prime}=G$. Note that $\phi([x,y])=[\phi(x),\phi(y)]=[x,y]$ for each $x,y\in G$. Since $G$ is generated by commutators, we have $z_{g}=1$ for each $g$. Thus, again $\phi$ is trivial and so, $C_{\mathop{\mathrm{Aut}}(G)}(\mathop{\mathrm{Inn}}(G))=1$. $\qed$
11 Structure of groups whose commuting graphs are cographs or chordal graphs
In this section, we will investigate the group-theoretical structure of a group whose commuting graph is a cograph or a chordal graph. In particular, in some sense, the class of such groups generalizes the class of finite solvable groups.
First, we investigate special cases of groups like direct and central products and Frobenius groups whose commuting graph is a cograph or a chordal graph, and then prove the general structure theorem.
11.1 Direct and central products
Observation 4
Let $A$ be an abelian group. Then $\Gamma(A\times G)$ is a cograph (resp. a chordal graph) if and only if $\Gamma(G)$ is a cograph (resp. a chordal graph).
Lemma 11.1
Let $H$ and $K$ be non-abelian groups. Then $\Gamma(H\times K)$ has an induced subgraph isomorphic to $P_{4}$ and an induced subgraph isomorphic to $C_{4}$.
Proof 25
Now let $a$ and $b$ be non-commuting elements of $G$, and $u$ and $v$ non-commuting elements of $H$. Then
$(a,u)\sim(e_{H},u)\sim(b,e_{K})\sim(e_{H},v)$
is an induced path with four vertices, and
$(a,e_{K})\sim(e_{H},u)\sim(b,e_{K})\sim(e_{H},v)\sim(a,e_{K})$
is an induced cycle of length $4$, as desired. $\qed$
By Lemma 11.1 and Observation 4, the following result holds.
Proposition 11.2
Let $H$ and $K$ be groups. Then
  1. (a)
    $\Gamma(H\times K)$ is a cograph if and only if one of $H$ and $K$ is abelian and the commuting graph of the other is $P_{4}$-free;
  2. (b)
    $\Gamma(H\times K)$ is chordal if and only if one of $H$ and $K$ is abelian and the other is $C_{n}$-free for all $n\geq 4$.
Recall that a finite nilpotent group is the direct product of its Sylow subgroups. Thus, Proposition 11.2 implies the following result.
Corollary 11.3
Let $G$ be a non-abelian nilpotent group. Then $\Gamma(G)$ is a cograph (resp. a chordal graph) if and only if there exists exactly one non-abelian Sylow subgroup in $G$ whose commuting graph is $P_{4}$-free (resp. $C_{n}$-free for all $n\geq 4$), while the other Sylow subgroups are abelian.
We saw in the first part the definition of the strong product of graphs and the fact that the commuting graph of $G\times H$ is $\Gamma(G)\boxtimes\Gamma(H)$ (Proposition 5.1).
We can extend this to central products as follows. Let $\mathfrak{N}_{2}$ be the class of all finite nilpotent groups of nilpotency class at most $2$ (thus, the class consists of finite abelian groups and finite non-abelian groups of nilpotency class $2$).
Lemma 11.4
Let $H$ and $K$ be two groups not in the class $\mathfrak{N}_{2}$. Then $\Gamma(H\circ K)$ is neither a cograph nor a chordal graph.
Proof 26
By assumption, we can choose $a,b\in H$ such that the commutator $[a,b]$ is not in the centre of $H$, and similarly $u,v\in K$. Then
$(a,u)\sim(a,e_{K})\sim(e_{H},v)\sim(b,v)$
carries a subgroup $P_{4}$ of $\Gamma(H\times K)$ such that the projection to $H\circ K$ is an isomorphism, and also
$(a,e_{K})\sim(e_{H},u)\sim(b,e_{K})\sim(e_{H},v)\sim(a,e_{K})$
is a subgraph isomorphic to $C_{4}$ which projects isomorphically to $H\circ K$. $\qed$
11.2 Frobenius groups
In this sections we investigate when the commuting graph of a Frobenius group is a cograph or a chordal graph. We use the structure of the commuting graphs of Frobenius groups established in Theorem 6.2. We prove the following assertion.
Proposition 11.5
Let $G$ be a Frobenius group with Frobenius complement $H$ and Frobenius kernel $K$ and $\mathcal{F}$ be a collection of graphs such that no graph in $\mathcal{F}$ has a dominating vertex. Then $\Gamma(G)$ is $\mathcal{F}$-free if and only if $\Gamma(K)$ is $\mathcal{F}$-free and $\Gamma(H)$ is $\mathcal{F}$-free.
Proof 27
This is clear from our remarks about dominating vertices and the fact (observed in Theorem 6.2) that $\Gamma(G\setminus\{1\})$ is the disjoint union of $\Gamma(H\setminus\{1\})$ and $|H|$ copies of $\Gamma(K\setminus\{1\})$.
Corollary 11.6
Let $G$ be a Frobenius group with Frobenius complement $H$ and Frobenius kernel $K$. Then $\Gamma(G)$ is a cograph (resp. chordal graph) if and only if $\Gamma(K)$ is a cograph (resp. a chordal graph) and $\Gamma(H)$ is a cograph (resp. a chordal graph).
Question 4
For which Frobenius groups is the commuting graph a cograph or a chordal graph?
Remark 9
Since $K$ is nilpotent, Corollary 11.3 gives a criterion for $\Gamma(K)$ to be a cograph (resp. a chordal graph).
11.3 A general structure theorem
The main result of this section is the following theorem.
Theorem 11.7
Let $G$ be a finite group such that $\Gamma(G)$ is a cograph (resp. a chordal graph), and let $F(G)$ be the Fitting subgroup of $G$.
  1. (a)
    One of the following statements holds:
    1. (i)
      $C_{G}(F(G))=Z(F(G))$;
    2. (ii)
      $F(G)$ is of nilpotent class at most $2$ and $G/F(G)$ is almost simple; in particular, $G$ has a unique non-abelian composition factor.
  2. (b)
    If $\mathop{\mathrm{Soc}}(G)$ is non-abelian, then $G$ has a normal subgroup $N=A\times S$, where $A$ is abelian, $S$ is simple and $\Gamma(S)$ is a cograph (resp. a chordal graph), such that $G/N$ is isomorphic to a subgroup of $\mathop{\mathrm{Out}}(S)$ and, therefore, $G/N$ is solvable.
Proof 28
Let $F^{*}(G)$ be the generalized Fitting subgroup of $G$. Let $H_{1}$ and $H_{2}$ be distinct elements from $\mathop{\mathrm{Comp}}(G)$. By [3, 31.5], $H_{1}$ and $H_{2}$ commute, therefore $\langle H_{1},H_{2}\rangle=H_{1}\circ H_{2}$ is a subgroup of $G$. By Lemma 11.4, $\Gamma(G)$ is neither a cograph nor a chordal graph, a contradiction to Proposition 10.2. Thus, $|\mathop{\mathrm{Comp}}(G)|\leq 1$.
If $|\mathop{\mathrm{Comp}}(G)|=0$, then $F^{*}(G)=F(G)$, and because we have $C_{G}(F^{*}(G))\leq F^{*}(G)$ by [3, 31.13], statement (a)(i) holds.
Assume that $|\mathop{\mathrm{Comp}}(G)|=1$. Then $G$ has a quasisimple characteristic subgroup $N=E(G)$. Consider $C=C_{G}(N)$, which is a normal subgroup of $G$. Let $H=\langle N,C\rangle=N\circ C$. By Proposition 10.2 and Lemma 11.4, $C$ is nilpotent of nilpotency class at most $2$, therefore $C\leq F(G)$. On the other hand, $F(G)\leq C$ by (1). Thus, $C=F(G)$ and, therefore $G/F(G)$ is a subgroup of $\mathop{\mathrm{Aut}}(N)$ which has a simple normal subgroup $H/C\cong N/Z(N)$.
Now we show that $\mathop{\mathrm{Aut}}(N)$ is an almost simple group and $\mathop{\mathrm{Soc}}(\mathop{\mathrm{Aut}}(N))\cong N/Z(N)$. Indeed, $N/Z(N)\cong\mathop{\mathrm{Inn}}(N)$ is a normal subgroup of $\mathop{\mathrm{Aut}}(N)$ and $N/Z(N)$ is simple. By Proposition 10.3, we have
$\mathop{\mathrm{Aut}}(N)\cong\mathop{\mathrm{Aut}}(N)/C_{\mathop{\mathrm{Aut}}(N)}(\mathop{\mathrm{Inn}}(N))$
is a subgroup of $\mathop{\mathrm{Aut}}(\mathop{\mathrm{Inn}}(N))\cong\mathop{\mathrm{Aut}}(N/Z(N))$. So, we have that $\mathop{\mathrm{Aut}}(N)$ is a subgroup of an almost simple subgroup $\mathop{\mathrm{Aut}}(N/Z(N))$ which contains a normal subgroup $\mathop{\mathrm{Inn}}(N)\cong N/Z(N)$. Thus, $\mathop{\mathrm{Aut}}(N)$ is an almost simple group with socle $N/Z(N)$, as required.
So, $G/F(G)$ is isomorphic to a subgroup of $\mathop{\mathrm{Aut}}(N)$ which is almost simple with socle isomorphic to $\mathop{\mathrm{Inn}}(N)$ and contains a subgroup $H/F(G)\cong\mathop{\mathrm{Inn}}(N)$, therefore $G/F(G)$ is almost simple. Thus, statement (a)(ii) holds.
Now we prove statement (b). Assume that $L=\mathop{\mathrm{Soc}}(G)$ is non-abelian. Then $L=A\times S$, where $A$ is abelian and $S$ is the direct product of non-abelian simple groups, moreover, both $A$ and $S$ are normal subgroups of $G$. By Propositions 10.2 and 11.2, $S$ is simple and $\Gamma(S)$ is a cograph (resp. a chordal graph). If $C=C_{G}(S)$, then $C\unlhd G$ and $C\cap S=1$ since $S$ is simple. Thus, $N=SC=S\times C$ is a normal subgroup of $G$. By Proposition 10.2, $\Gamma(N)$ is also a cograph (resp. a chordal graph). Using Proposition 11.2, we conclude that $C$ is abelian. Moreover, $G/C$ is isomorphic to a subgroup of $\mathop{\mathrm{Aut}}(S)$ which is almost simple, therefore $G/N$ is isomorphic to a subgroup of $\mathop{\mathrm{Out}}(S)$. Thus, by the Schreier Conjecture (Theorem 2.1), $G/N$ is solvable, and therefore statement (b) holds. $\qed$
Remark 10
If $\Gamma(G)$ is a cograph, then in statement $(b)$ of Theorem 11.7 the group $G/N$ is cyclic. This result follows from the classification of finite simple groups whose commuting graphs are cographs, see Theorem 12.15 below.
The following problem is of interest.
Problem 1
Classify finite simple groups whose commuting graphs are chordal.
12 Simple groups whose commuting graph is a cograph or a chordal graph
In this section, we investigate when the commuting graph of a simple sporadic or alternating group is a cograph or a chordal graph, and then obtain a complete classification of finite simple groups whose commuting graph is a cograph.
12.1 CA-groups and AC-groups
A group $G$ is a CA-group if the centralizer of every non-identity element of $G$ is abelian. Another name is CT-groups, since they are the groups for which the commutation relation on non-identity elements is transitive. (If $x\sim y\sim z$ in the CA-group $G$, then $x,z\in C_{G}(y)$, so $x$ and $z$ commute. The converse is similar.)
If $G$ is a CA-group, then the induced subgraph $\Gamma(G)$ on non-identity elements is a disjoint union of complete graphs. As a result, we see:
Proposition 12.1
Let $G$ be a CA-group. Then $\Gamma(G)$ is a cograph, chordal, and perfect.
CA-groups played an important historical role in developing techniques for attacking problems on finite groups; they are now well-understood. From results of Weisner [77], Brauer, Suzuki and Wall [13] and Suzuki [73], we obtain:
Proposition 12.2
A finite CA-group $G$ satisfies one of the following:
  1. (a)
    $G$ is abelian;
  2. (b)
    $G$ is a Frobenius group with abelian Frobenius kernel and cyclic Frobenius complement; or
  3. (c)
    $G\cong\mathop{\mathrm{PSL}}\nolimits(2,q)$ for $q$ a power of $2$.
More generally, $G$ is said to be an AC-group if $C_{G}(x)$ is abelian for every element $x\in G\setminus Z(G)$. This is a weaker condition, though clearly a group with trivial centre is an AC-group if and only if it is a CA-group.
AC-groups investigated by Schmidt [65] and Rocke [64]. (Schmidt investigated groups in which two maximal centralizers intersect just in $Z(G)$; it is not hard to see that, for finite groups, this is equivalent to the AC property.) Abdollahi et al. [2] investigated the connection of this property with the non-commuting graph, the complement of the commuting graph. The classification of the AC-groups is not as complete as that of CA-groups, but the cited authors have proved a number of results about these groups.
From our point of view, AC-groups are as good as CA-groups:
Proposition 12.3
Let $G$ be an AC-group. Then $\Gamma(G)$ is a cograph and a chordal graph.
Proof 29
By Proposition 10.2, it suffices to consider the induced subgraph on the set $G\setminus Z(G)$; and this graph is a disjoint union of complete graphs. $\qed$
Here is a simple general result which we will use.
Proposition 12.4
Let $G$ be a finite group with the property that every non-identity element of $G/Z(G)$ has cyclic centralizer. Then $G$ is an AC-group.
Proof 30
Take $a\in G\setminus Z(G)$. Then $Z(G)\leq C_{G}(a)$, and $C_{G}(a)/Z(G)\leq C_{G}(aZ(G))$ (in $G/Z(G)$); so $C_{G}(a)/Z(G)$ is cyclic, whence $C_{G}(a)$ is abelian. $\qed$
12.2 Generalized dihedral groups and generalized quaternion groups
In this subsection we show that these groups are AC-groups, so that their commuting graphs are cographs as well as chordal graphs.
Proposition 12.5
Let $D(A)$ be the generalized dihedral group as presented in Equation (4). Then $D(A)$ is an AC-group. Hence $\Gamma(D(A))$ is a cograph as well as a chordal graph. In particular, the commuting graph of a dihedral group is a cograph as well as a chordal graph.
Proof 31
The centre of $D(A)$ is the subgroup $T$ of $A$ consisting of elements of orders $1$ or $2$. If we remove it, then $A\setminus T$ is a clique, and the elements outside $A$ fall into cliques of the form $\langle t,T\rangle\setminus T$ for $t\notin A$; there are no edges between these cliques. So the induced subgraph is a disjoint union of complete graphs. Now Proposition 10.2 gives the result.$\qed$
Proposition 12.6
Let $Q_{4m}$ be the generalized quaternion group as presented in Equation (5). Then $Q_{4m}$ is an AC-group. Hence $\Gamma(Q_{4m})$ is a cograph as well as a chordal graph.
Proof 32
The proof is similar to the preceding theorem. We find that, with $G=Q_{4m}$, we have $Z(G)=\{e,x^{m}\}$ (Observation 2); for $11\leq i\leq 2m-$ 1 and $i\neq m$, we find that $C_{G}(x^{i})=\langle x\rangle$, while for $10\leq i\leq 2m-$ 1 we have $C_{G}(x^{i}y)=\langle x^{i}y\rangle=\{e,x^{i}y,x^{m},x^{m+i}y\}$. So again the induced subgraph on $G\setminus Z(G)$ consists of a complete graph of size $22m-$ 2 together with $m$ disjoint edges. $\qed$
12.3 Symmetric groups and alternating groups
Proposition 12.7
$\Gamma(S_{n})$ is a cograph if and only if $n\leq 3$.
Proof 33
Proposition 10.2 shows that it is enough to show that $\Gamma(S_{3})$ is a cograph but $\Gamma(S_{4})$ is not (since $S_{n-1}\leq S_{n}$ for any $n$). By Proposition 12.5, we have that $\Gamma(S_{3})$ is a cograph as $S_{3}\cong D_{6}$. As $S_{n-1}\leq S_{n}$, Proposition 10.2 shows that if $n\leq 3$, then $\Gamma(S_{n})$ is a cograph. In $\Gamma(S_{4})$,
$(12)\sim(12)(34)\sim(13)(24)\sim(13)$
is an induced path isomorphic to $P_{4}$$\qed$
Proposition 12.8
$\Gamma(S_{n})$ is a chordal graph if and only if $n\leq 4$.
Proof 34
It is easy to see that
$(12)\sim(34)\sim(15)\sim(24)\sim(35)\sim(12)$
is an induced cycle isomorphic to $C_{5}$ in $\Gamma(S_{5})$. As in Proposition 12.7, it suffices to prove that $\Gamma(S_{4})$ is a chordal graph. Note that in $\Gamma(S_{4})$, for any non-central element, the cycle type of this element must be one of
$\{(ij),(ij)(kl),(ijk),(ijkl)\},$ (10)
where $\{i,j,k,l\}=\{1,2,3,4\}$. Moreover, it is easy to see that
$C_{S_{4}}((ij))=\{e,(ij),(kl),(ij)(kl)\},C_{S_{4}}((ij))=\{e,(ij),(kl),(ij)(kl)\}$
$C_{S_{4}}((ijkl))=\langle(ijkl)\rangle$ (11)
and
$C_{S_{4}}((ij)(kl))=\{e,(ij),(kl),(ij)(kl),(ik)(jl),(il)(jk),(ikjl),(iljk)\}.$
Suppose, for the sake of contradiction, that $\Gamma(S_{4})$ contains an induced subgraph isomorphic to $C_{n}$ with $n\geq 4$, say $\Delta$. Let $a,b,c,d$ be $4$ distinct vertices of $\Delta$. Then $C_{S_{4}}(x)$ is non-abelian for any $x\in\{a,b,c,d\}$. Note that every centralizer in (11) is abelian. It follows from (10) that every element in $\{a,b,c,d\}$ has cycle type $(ij)(kl)$, which is impossible because $S_{4}$ has precisely three elements that are the product of two disjoint cycles. $\qed$
Proposition 12.9
The following are equivalent:
$(a)$ $\Gamma(A_{n})$ is a cograph;
$(b)$ $\Gamma(A_{n})$ is a chordal graph;
$(c)$ $n\leq 5$.
Proof 35
The group $A_{5}$ is a CA-group (it is easily checked from [28] that all its non-abelian subgroups have trivial centre), so the result for $n=5$ follows from Proposition 12.2. Moreover, $A_{n}\leq A_{5}$ for $n\leq 5$, so the result follows for these groups as well.
On the other hand, in $\Gamma(A_{6})$, it is easy to verify that
$(12)(56)\sim(12)(34)\sim(13)(24)\sim(13)(56)$
is an induced path isomorphic to $P_{4}$. Also, it is not hard to verify that
$(13)(45)\sim(14)(35)\sim(14)(26)\sim(12)(46)\sim(12)(35)\sim(13)(25)\sim(13)(46)$
$\sim(14)(36)\sim(14)(25)\sim(12)(45)\sim(12)(36)\sim(13)(26)\sim(13)(45)$
is an induced cycle isomorphic to $C_{12}$, so $\Gamma(A_{6})$ is not chordal. Now $A_{6}\leq A_{n}$ for $n>6$, so $\Gamma(A_{n})$ is neither a cograph nor chordal for $n\geq 6$, by Proposition 10.2. $\qed$
Remark 11
The graph $\Gamma(A_{6})$ has no induced cycles of length $n$, where $123<n<$ 12. However, $\Gamma(A_{7})$ has an induced cycle isomorphic to $C_{6}$, such as $(123)\sim(456)\sim(127)\sim(345)\sim(126)\sim(457)\sim(123)$. Also, $\Gamma(A_{8})$ has an induced cycle isomorphic to $C_{4}$, such as $(12)(34)\sim(567)\sim(348)\sim(12)(56)\sim(12)(34)$.
12.4 Sporadic groups
In this subsection, we resolve the question for the sporadic simple groups.
Proposition 12.10
If $G$ is a sporadic simple group, then $\Gamma(G)$ is neither a cograph nor a chordal graph.
Proof 36
There are $26$ sporadic simple groups. Observe that $S_{5}$ is a subgroup of the Mathieu group $M_{11}$ by using information in the $\mathbb{ATLAS}$ of finite groups [28]. Also, note that the fact that the Mathieu group $M_{11}$ is a subgroup of all the other sporadic simple groups except $M_{22}$, $J_{1}$, $J_{2}$, $J_{3}$, $He$, $Ru$ and $Th$. Thus, by Proposition 10.2 and Propositions 12.7 and 12.8, the commuting graph of a sporadic simple group containing $M_{11}$ is neither a cograph nor a chordal graph.
Now by [28] again, it is easy to observe that $A_{7}\leq M_{22}$, $D_{6}\times D_{10}\leq J_{1}$, $A_{4}\times A_{5}\leq J_{2}$, $A_{6}\leq J_{3}$, $S_{6}\leq He$, $A_{8}\leq Ru$, $S_{5}\leq Th$. Thus, it follows from Proposition 10.2, Lemma 11.1 and Proposition 12.9 that the commuting graphs of these $7$ sporadic simple groups are neither cographs nor chordal graphs, as desired. $\qed$
12.5 Simple groups of Lie type whose commuting graph is a cograph
Now we turn to simple groups of Lie type. In this analysis, we only consider cographs, not chordal graphs.
Proposition 12.11
Let $G=\mathop{\mathrm{PSL}}\nolimits(2,q)$, with $q\geq 4$. Then $\Gamma(G)$ is a cograph if and only if either $q$ is a power of $2$ or $q=5$.
Note that $\mathop{\mathrm{PSL}}\nolimits(2,5)\cong\mathop{\mathrm{PSL}}\nolimits(2,4)$.
Proof 37
The subgroup structure of $\mathop{\mathrm{PSL}}\nolimits(2,q)$ was determined by Dickson [31], and can be found, for example, in [14].
Suppose first that $q\geq 4$ is a power of $2$. Then every non-trivial element of $\mathop{\mathrm{PSL}}\nolimits(2,q)$ has order $2$ or a prime divisor of $q-1$ and $q+1$. Also, the centralizers of involutions of $\mathop{\mathrm{PSL}}\nolimits(2,q)$ are elementary abelian groups with order $q$, and the centralizers of other non-trivial elements of $\mathop{\mathrm{PSL}}\nolimits(2,q)$ are cyclic of order $q-1$ or $q+1$. It follows that in this case $\mathop{\mathrm{PSL}}\nolimits(2,q)$ is a CA-group, and so by Proposition 12.2, $\Gamma(\mathop{\mathrm{PSL}}\nolimits(2,q))$ is a cograph.
Suppose next that $q$ is odd at least $5$. Let $\epsilon\in\{+1,-1\}$ be chosen so that $q+\epsilon 1$ is divisible by $4$. Note that $\mathrm{PSL}(2,q)$ contains a Klein group such that the centralizers of its involutions are Dihedral groups $D_{2(q+\epsilon 1)}$ which contain disjoint cyclic groups of order $(q+\epsilon)/2$. If $q>5$, then $(q+\varepsilon)/2>2$. So two involutions together with generators of cyclic groups of order $(q+\epsilon)/2$ in their centralizers induce a path of length $4$. (The list of subgroups of $G$ shows that these generators do not commute.) So $\Gamma(\mathop{\mathrm{PSL}}\nolimits(2,q))$ is not a cograph. $\qed$
Remark 12
It is well-known that $\mathop{\mathrm{PSL}}\nolimits(2,2)\cong S_{3}$ and $\mathop{\mathrm{PSL}}\nolimits(2,3)\cong A_{4}$, therefore their commuting graphs are cographs by Propositions 12.7 and 12.9, respectively.
Proposition 12.12
Let $G$ be either $\mathop{\mathrm{PSL}}\nolimits(3,q)$ or a group of Lie type of BN-rank at least $3$. Then $\Gamma(G)$ is not a cograph.
Proof 38
We begin by showing that, if $G=\mathop{\mathrm{SL}}\nolimits(3,q)$, then $\Gamma(G)$ is not a cograph.
Suppose first that $q\notin\{2,4\}$. Choose a non-zero element $a\in\mathop{\mathrm{GF}}\nolimits(q)$ such that $a^{3}\neq 1$, and let $b=a^{-2}$, so that $b\neq a$.
Consider the four elements $g_{1},\ldots,g_{4}$ given by
$\left(\begin{array}[]{ccc}1&1&0\\ 0&1&0\\ 0&0&1\end{array}\right),\,\left(\begin{array}[]{ccc}a&0&0\\ 0&a&0\\ 0&0&b\end{array}\right),\,\left(\begin{array}[]{ccc}b&0&0\\ 0&a&0\\ 0&0&a\end{array}\right),\,\left(\begin{array}[]{ccc}1&0&0\\ 0&1&1\\ 0&0&1\end{array}\right).$
It is clear that all of them belong to $G$, and that $g_{1}\sim g_{2}\sim g_{3}\sim g_{4}$ in $\Gamma(G)$. A short calculation gives the commutators of the other three pairs; none of them is trivial. So we have an induced path in $\Gamma(G)$.
But the argument gives more. In fact none of the commutators is a scalar matrix, so if $Z$ is any subgroup of the centre of $G$ (the group of scalar matrices), then we also have an induced path in $\Gamma(G/Z)$, so that $\Gamma(G/Z)$ is not a cograph. In particular, $\Gamma(\mathop{\mathrm{PSL}}\nolimits(3,q))$ is not a cograph.
Now we can cover the remaining two cases. Since $\mathop{\mathrm{PSL}}\nolimits(3,2)\cong\mathop{\mathrm{PSL}}\nolimits(2,7)$, the commuting graph of $\mathop{\mathrm{PSL}}\nolimits(3,2)$ is not a cograph. Also, $\mathop{\mathrm{PSL}}\nolimits(3,2)$ is a subgroup of both $\mathop{\mathrm{SL}}\nolimits(3,4)$ and $\mathop{\mathrm{PSL}}\nolimits(3,4)$, so their commuting graphs are not cographs either.
For the remainder of the proof, we use the Levi decomposition of parabolic subgroups of groups with BN-pairs (see [25, Section 8.5]). The Borel subgroup $H$ of such a group is generated by the root subgroups corresponding to the positive roots with respect to a fundamental system $\Pi$ of roots. Parabolic subgroups $P_{\Phi}$ are those which contain $H$; they are generated by $H$ together with the root subgroups corresponding to the negatives of a subset $\Phi$ of the negatives of the fundamental group. According to the Levi decomposition, $P_{\Phi}$ is the semidirect product of its unipotent radical by the Levi factor $L_{\Phi}$, a group with BN-pair corresponding to the roots $\Phi$. In the covering group, $L_{\Phi}$ may have a centre, which may or may not be killed when we factor out the centre of $G$. (For example, $\mathop{\mathrm{SL}}\nolimits(n,q)$ has a parabolic subgroup corresponding to the first two roots, generated by all upper unitriangular matrices together with the lower unitriangular matrices with non-zero off-diagonal elements in positions $1,2,3$ only. These generate $\mathop{\mathrm{SL}}\nolimits(3,q)$; its image in $\mathop{\mathrm{PSL}}\nolimits(n,q)$ is either $\mathop{\mathrm{SL}}\nolimits(3,q)$ or $\mathop{\mathrm{PSL}}\nolimits(3,q)$ depending on the congruence of $n$ and $q$ mod $3$.)
Now the Coxeter–Dynkin diagram of any simple group $G$ of Lie type with BN-pair of rank at least $3$ contains a single edge, corresponding to a root system of type $A_{2}$; so $G$ has a parabolic subgroup whose Levi factor is a central quotient of $\mathop{\mathrm{SL}}\nolimits(3,q)$, as in the example above. From Proposition 10.2, we conclude that $\Gamma(G)$ is not a cograph. $\qed$
Proposition 12.13
Let $G$ be a central quotient of $\mathop{\mathrm{SU}}\nolimits(3,q)$, with $q>2$. Then $\Gamma(G)$ is not a cograph.
Proof 39
Suppose that $q>2$. We take the Hermitian form $H(x_{1},x_{2},x_{3})=x_{1}^{q+1}+x_{2}^{q+1}+x_{3}^{q+1}$. A diagonal matrix with diagonal entries $(a,b,c)$ preserves this form if and only if
$a^{q+1}=b^{q+1}=c^{q+1}=1,$
and belongs to $\mathop{\mathrm{SU}}\nolimits(3,q)$ if additionally $abc=1$. So the set of such matrices forms a subgroup of $\mathop{\mathrm{SU}}\nolimits(3,q)$ isomorphic to $\mathbb{Z}_{q+1}^{2}$, and its quotient in $\mathop{\mathrm{PSU}}\nolimits(3,q)$ is $\mathbb{Z}_{q+1}\times\mathbb{Z}_{(q+1)/\gcd(q+1,3)}$. We observe that $(-1)^{q+1}=1$ (this is clear if $q$ is even and holds if $q$ is odd since then $(-1)^{2}=1$).
Let $a\in GF(q^{2})$ satisfy $a^{q+1}=1$, $a^{3}\neq 1$, and let $b=a^{-2}$, so that $a\neq b$. Now consider the four matrices $g_{1}$, …, $g_{4}$ given by
$\left(\begin{array}[]{ccc}0&1&0\\ 1&0&0\\ 0&0&-1\end{array}\right),\,\left(\begin{array}[]{ccc}a&0&0\\ 0&a&0\\ 0&0&b\end{array}\right),\,\left(\begin{array}[]{ccc}b&0&0\\ 0&a&0\\ 0&0&a\end{array}\right),\,\left(\begin{array}[]{ccc}-1&0&0\\ 0&0&1\\ 0&1&0\end{array}\right).$
It is readily checked that all four matrices belong to $\mathop{\mathrm{SU}}\nolimits(3,q)$, and that $g_{2}$ commutes with $g_{1}$ and $g_{3}$ while $g_{3}$ commutes with $g_{4}$. The other pairs fail to commute; indeed in each case $g_{i}g_{j}$ is not a scalar multiple of $g_{j}g_{i}$. So the four matrices induce a $P_{4}$ subgraph of the commuting graph, which remains as a $P_{4}$ subgraph when the scalar matrices in $\mathop{\mathrm{SU}}\nolimits(3,q)$ are factored out. $\qed$
Among the groups of Lie type, this leaves only those of ranks $1$ or $2$, and we have already handled $\mathop{\mathrm{PSL}}\nolimits(2,q)$, $\mathop{\mathrm{PSL}}\nolimits(3,q)$, and $\mathop{\mathrm{PSU}}\nolimits(3,q)$ for $q>2$. The remaining groups are:
  1. (a)
    $\mathrm{PSp}(4,q)$;
  2. (b)
    $\mathrm{PSU}(n,q)$ for $n\in\{4,5\}$;
  3. (c)
    $G_{2}(q)$, ${}^{2}B_{2}(q)\cong\mathop{\mathrm{Sz}}\nolimits(q)$, ${}^{2}G_{2}(q)$, ${}^{2}F_{4}(q)$, and ${}^{3}D_{4}(q)$.
We have omitted the orthogonal groups here. For these satisfy
  1. (a)
    $\mathop{\mathrm{P}\Omega}\nolimits(3,q)\cong\mathop{\mathrm{PSL}}\nolimits(2,q)$;
  2. (b)
    $\mathop{\mathrm{P}\Omega}\nolimits^{+}(4,q)\cong\mathop{\mathrm{PSL}}\nolimits(2,q)\times\mathop{\mathrm{PSL}}\nolimits(2,q)$ and $\mathop{\mathrm{P}\Omega}\nolimits^{-}(4,q)\cong\mathop{\mathrm{PSL}}\nolimits(2,q^{2})$;
  3. (c)
    $\mathop{\mathrm{P}\Omega}\nolimits(5,q)\cong\mathop{\mathrm{PSp}}\nolimits(4,q)$;
  4. (d)
    $\mathop{\mathrm{P}\Omega}\nolimits^{+}(6,q)\cong\mathop{\mathrm{PSL}}\nolimits(4,q)$;
  5. (e)
    $\mathop{\mathrm{P}\Omega}\nolimits^{-}(6,q)\cong\mathop{\mathrm{PSU}}\nolimits(4,q)$
(see the $\mathbb{ATLAS}$ of Finite Groups [28]). So (a), (b) and (d) are covered by our earlier results, and (c) and (e) will be treated as symplectic or unitary groups.
Also $G_{2}(q)$ contains a subgroup $\mathop{\mathrm{PSL}}\nolimits(3,q)$, and ${}^{3}D_{4}(q)$ contains $G_{2}(q)$ [29, 50, 51], so their commuting graphs are not cographs, by Proposition 10.2 and Proposition 12.12.
We now finish off the classical groups.
The group $\mathop{\mathrm{PSp}}\nolimits(4,q)$ contains the central product of two copies of $\mathop{\mathrm{SL}}\nolimits(2,q)$. (To see this, note that the $4$-dimensional vector space can be written as the orthogonal direct sum of two $2$-dimensional non-degenerate subspaces; so $\mathop{\mathrm{Sp}}\nolimits(4,q)$ contains $\mathop{\mathrm{SL}}\nolimits(2,q)\times\mathop{\mathrm{SL}}\nolimits(2,q)$. If $q$ is odd, take the quotient by {±$\{\pm I\}$ to obtain a central product.) So its commuting graph is not a cograph, by Lemma 11.4.
For unitary groups, the $\mathop{\mathrm{PSU}}\nolimits(3,2)$ is not simple; so for $q=2$ it remains to deal with $\mathop{\mathrm{PSU}}\nolimits(4,2)$ and $\mathop{\mathrm{PSU}}\nolimits(5,2)$. The first is isomorphic to $\mathop{\mathrm{PSp}}\nolimits(4,3)$ [28], which we have dealt with. The second contains $\mathop{\mathrm{PSL}}\nolimits(2,11)$, see [28]. For $q>2$, by Proposition 12.13 central quotients of $\mathop{\mathrm{SU}}\nolimits(3,q)$ have commuting graphs which are not cographs. For $n=5$, $\mathop{\mathrm{PSU}}\nolimits(n,q)$ contains such a central quotient, obtained as above by writing the $5$-dimensional unitary space as the sum of non-degenerate subspaces of dimensions $2$ and $3$. For $n=4$, the space is the sum of two non-degenerate $2$-dimensional subspaces, so $\mathop{\mathrm{PSU}}\nolimits(4,q)$ contains a central product of two copies of $\mathop{\mathrm{SU}}\nolimits(2,q)$. So the commuting graphs of these groups are not cographs either.
This concludes dealing with the classical groups.
We show next that the commuting graphs of the Suzuki groups are cographs.
Proposition 12.14
For each $q$, the commuting graph of $\mathop{\mathrm{Sz}}\nolimits(q)$ is a cograph.
Proof 40
It is well-known that $\mathop{\mathrm{Sz}}\nolimits(2)\cong 5:4$ is a Frobenius groups with cyclic kernel and cyclic complement. Therefore, by Proposition 11.5, $\Gamma(\mathop{\mathrm{Sz}}\nolimits(2))$ is a cograph. Thus, we can assume that $q>2$.
The information used in the next two paragraphs is taken from Suzuki’s original paper [74]; it can also be found in [79, Section 4.2].
It suffices to show that the graph obtained by removing the identity is a cograph. Now $\mathop{\mathrm{Sz}}\nolimits(q)$ has maximal cyclic subgroups of orders $q-1$ and $q\pm\sqrt{2q}+1$, and every element of odd order lies in one of these; so the non-identity elements of odd order carry a disjoint union of complete graphs, with no edges to the remaining elements. The remaining elements have order $2$ or $4$, and since $\mathop{\mathrm{Sz}}\nolimits(q)$ is a CIT-group (one in which the centralizer of any involution is a $2$-group), the remaining vertices lie in a disjoint union of copies of the commuting graph on $P\setminus\{e\}$, where $P$ is a Sylow $2$-subgroup. Now $|P|=q^{2}$ and $|Z(P)|=q$; elements of $Z(P)$ are joined to every element of $P$ so we may remove them. The remaining elements of $P$ have order $4$. If $g$ is one of these, then $|C_{P}(g)|=2q$ and $C_{P}(g)$ consists of two cosets of $Z(P)$, so it is abelian; thus the elements of order $4$ fall into $q-1$ sets of size $q$, each set a complete graph with no edges between different sets. Thus the graph is indeed a cograph. $\qed$
Next, the small Ree group ${}^{2}G_{2}(q)$, with $q$ an odd power of $3$, is not simple for $q=3$; for larger $q$, it contains $\mathop{\mathrm{PSL}}\nolimits(2,q)$ (the centralizer of an involution in ${}^{2}G_{2}(q)$ is $\mathbb{Z}_{2}\times\mathop{\mathrm{PSL}}\nolimits(2,q)$ [51]).
The Tits group ${}^{2}F_{4}(2)^{\prime}$ contains a subgroup isomorphic to $\mathop{\mathrm{PSL}}\nolimits(2,25)$ (this can be found in the $\mathbb{ATLAS}$ [28]); so its commuting graph is not a cograph. The group ${}^{2}F_{4}(2)^{\prime}$ is a subgroup of the large Ree group ${}^{2}F_{4}(2^{d})$ for all odd $d$ [54].
To summarize our results, combined with the earlier results on alternating and sporadic groups, and noting that $\mathop{\mathrm{Alt}}\nolimits(5)\cong\mathop{\mathrm{PSL}}\nolimits(2,4)$:
Theorem 12.15
Let $G$ be a non-abelian finite simple group. Then the commuting graph of $G$ is a cograph if and only if $G$ is isomorphic to $\mathop{\mathrm{PSL}}\nolimits(2,q)$, with $q$ a power of $2$ and $q>2$, or to $\mathop{\mathrm{Sz}}\nolimits(q)$, with $q$ an odd power of $2$ and $q>2$.
13 Appendix
In this section, we consider small groups whose commuting graphs are cographs or chordal graphs.
Lemma 13.1
If $|G/Z(G)|=pqr$ with primes $p,q,r$ (not necessary distinct), then $G$ is an AC-group.
Proof 41
Let $a\in G\setminus Z(G)$. It suffices to prove that $C_{G}(a)$ is abelian. Note that $Z(G)\subseteq Z(C_{G}(a))\subseteq C_{G}(a)<G$. Since $|G/Z(G)|=pqr$, we have that $|C_{G}(a)/Z(G)|<pqr$ and $|C_{G}(a)/Z(G)|$ is a divisor of $pqr$. By the argument of Proposition 12.4, it suffices to consider the case where $C_{G}(a)/Z(G)$ is non-cyclic and has order $pq$.
Then $Z(G)a$ has order $p$ or $q$. Thus, $C_{G}(a)/Z(G)$ has an element, say $Z(G)b$, such that $C_{G}(a)/Z(G)=\langle Z(G)b,Z(G)a\rangle$. Note that $[b,a]=1$. Hence, $C_{G}(a)=\langle b,a,Z(G)\rangle$ is an abelian group, as desired. $\qed$
A similar but easier proof shows the following.
Corollary 13.2
If $|G/Z(G)|=pq$ or $p$ where $p,q$ are primes, then $G$ is an AC-group. In particular, If $|G|=pq$ or $pqr$ with primes $p,q,r$, then $G$ is either an abelian group or an AC-group.
Note that these primes $p,q,r$ in Lemma 13.1 and Corollary 13.2 are not necessarily distinct. Now combining Corollary 13.2, Lemma 13.1 and Proposition 12.3, we have the following result.
Lemma 13.3
$\Gamma(G)$ is a cograph as well as a chordal graph if one of the following holds:
  1. (a)
    $G$ is abelian;
  2. (b)
    $|G|=pqr$ or $pq$, where $p,q,r$ are primes (not necessarily distinct);
  3. (c)
    $|G/Z(G)|=p$, $pq$ or $pqr$, where $p,q,r$ are primes (not necessary distinct).
Proposition 13.4
If $G$ is a group of minimal order such that $\Gamma(G)$ is not a cograph, then $|G|=24$. In particular, for $|G|\leq 24$, $\Gamma(G)$ is a cograph if and only if $G\ncong S_{4}$.
Proof 42
Proposition 12.7 implies that $\Gamma(S_{4})$ is not a cograph. Thus, it suffices to show that if $|G|\leq 24$, then $\Gamma(G)$ is a cograph if and only if $G\ncong S_{4}$. Now suppose that $|G|\leq 24$. It follows from Lemma 13.3 that if $|G|\notin\{16,24\}$, then $\Gamma(G)$ is a cograph. If $|G|=16$, since a $2$-group has non-trivial centre, we have that $|G/Z(G)|=2^{k}$ with $k\leq 3$, and so Lemma 13.3(c) implies that $\Gamma(G)$ is a cograph. Finally, we suppose that $|G|=24$. Note that the fact that if $G\ncong S_{4}$, then $G$ has non-trivial centre (cf. Burnside [18]). So, by Lemma 13.3(c), we have that if $G\ncong S_{4}$, then $\Gamma(G)$ is a cograph, as desired. $\qed$
In this paper, we identify a finite group by the unique identifier of this finite group in SmallGroups library [10] in the GAP system [37]. To be specific, the $m$-th group of order $n$ is identified as SmallGroup$[n,m]$, where $[n,m]$ is called the GAP ID of the group.
By Lemma 11.1, $\Gamma(S_{3}\times S_{3})$ is not a cograph. One can use the GAP system [37] to classify all finite groups $G$ satisfying $|G|\leq 36$ and $\Gamma(G)$ is not a cograph; these groups are presented in Table 1. In particular, by Lemma 11.1, $\Gamma(S_{3}\times S_{3})$ is not a cograph.
Table 1: All groups $G$ such that $\Gamma(G)$ is not a cograph and $|G|\leq 36$
GAP ID Structure
$[24,12]$ $S_{4}$
$[32,6]$ $(\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2})\rtimes\mathbb{Z}_{4}$
$[32,7]$ $(\mathbb{Z}_{8}\rtimes\mathbb{Z}_{2})\rtimes\mathbb{Z}_{2}$
$[32,8]$ $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})_{\cdot}(\mathbb{Z}_{4}\times\mathbb{Z}_{2})$
$[32,43]$ $\mathbb{Z}_{8}\rtimes(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$
$[32,44]$ $(\mathbb{Z}_{2}\times Q_{8})\rtimes\mathbb{Z}_{2}$
$[32,49]$ $(\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2})\rtimes(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$
$[32,50]$ $(\mathbb{Z}_{2}\times Q_{8})\rtimes\mathbb{Z}_{2}$
$[36,10]$ $S_{3}\times S_{3}$
For the groups of order $32$, there are precisely two extra-special $2$-groups $D_{8}\circ D_{8}$ and $D_{8}\circ Q_{8}$ as follows:
$D_{8}\circ D_{8}$ $=$ $\langle x,y,z,w|x^{2}=y^{2}=z^{2}=w^{2}=[x,y]^{2}=[z,w]^{2}=1,$ (13)
$[x,z]=[z,y]=[y,w]=[w,x]=1\rangle$
where $[x,y]=[z,w]$ is an involution belonging to $Z(D_{8}\circ D_{8})$;
$D_{8}\circ Q_{8}$ $=$ $\langle x,y,z,w|x^{2}=y^{2}=z^{4}=w^{4}=[x,y]^{2}=[z,w]^{2}=1,$ (15)
$[x,z]=[z,y]=[y,w]=[w,x]=1\rangle$
where $z^{2}=w^{2}=[x,y]=[z,w]$ is an involution belonging to $Z(D_{8}\circ Q_{8})$.
Proposition 13.5
If $G$ is a group of minimal order such that $\Gamma(G)$ is not a chordal graph, then $|G|=32$. In particular, if $G\cong D_{8}\circ D_{8}$ or $D_{8}\circ Q_{8}$, then $\Gamma(G)$ is not chordal.
Proof 43
Suppose that $G\cong D_{8}\circ D_{8}$ or $D_{8}\circ Q_{8}$. Then by (13) and (15), it is easy to see that the induced subgraph of $\Gamma(G)$ by the set $\{x,y,z,w\}$ is isomorphic to $C_{4}$, which implies that $\Gamma(G)$ is not chordal.
Now, it suffices to show that if $|G|<32$, then $\Gamma(G)$ is a chordal graph. Suppose that $|G|\leq 31$. Note that a $p$-group has non-trivial centre where $p$ is a prime. From Lemma 13.3, it follows that if $|G|\neq 24$, then $\Gamma(G)$ is a chordal graph. Now consider $|G|=24$. If $G\cong S_{4}$, then by Proposition 12.8, $\Gamma(G)$ is chordal. If $G\ncong S_{4}$, then by Lemma 13.3 and the proof of Proposition 13.4, $\Gamma(G)$ is chordal, as desired. $\qed$
We have taken these computations further and determined, using GAP with the package GRAPE [70], the groups of order up to $128$ whose commuting graphs are not cographs. The numbers of groups are given in Table 2.
Table 2: Small groups whose commuting graph is not a cograph
$\begin{array}[]{|r|r|r|r|r|r|}\hline\cr\hbox{Order}&\hbox{Groups}&\hbox{Order}&\hbox{Groups}&\hbox{Order}&\hbox{Groups}\\ \hline\cr 24&1&32&7&36&1\\ 48&10&54&2&60&2\\ 64&115&72&11&80&12\\ 84&1&96&112&100&2\\ 108&10&112&8&120&15\\ 126&2&128&1539&&\\ \hline\cr\end{array}$
Acknowledgements
The authors of this paper contributed equally to this work. The authors are ordered with respect to alphabet ordering in English.
This work was initiated during the online series of Research Seminars on “Groups and Graphs” in March–August 2021, run by Ambat Vijayakumar and Aparna Lakshmanan, Cochin Univ. of Science and Technology.
The second author acknowledges the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Groups, representations and applications: new perspectives (supported by EPSRC grant no. EP/R014604/1), where he held a Simons Fellowship.
The third author was supported by National Natural Science Foundation of China (Grant Nos. 11801441 and 12326333) and Shaanxi Fundamental Science Research Project for Mathematics and Physics (Grant No. 22JSQ024).
The fourth author is thankful to Prof. Andrey V. Vasil’ev for his helpful comments which improved this text. Also the fourth author started thinking on this project during her visit to the Steklov Institute of Mathematics, and she is very thankful to colleagues from this institution for hospitality and moral support.
Acknowledgement.  We are grateful to the referee for thoughtful comments and suggestions that have helped improve our presentation.